### v1

# Sheffer polynomials and the s-ordering of exponential boson operators

Abstract
The $s$-ordered form of any product of single-mode boson creation and annihilation operators, containing only a single annihilator, is computed explicitly. The $s$-ordering concept originated in quantum optics, and subsumes normal, symmetric (Weyl), and anti-normal ordering for any two operators satisfying a canonical commutation relation. Because the $s$-ordering map can be viewed as producing a function of a complex variable, its inverse is a quantization map that takes such “classical” functions to quantum operators. The explicit $s$-ordered expressions are derived with the aid of a parametric family of Sheffer polynomial sequences (or equivalently an exponential Riordan array of polynomial coefficients), called the Hsu–Shiue family. To yield orderings interpolating between normal and anti-normal, this family is extended in an intricate way.
keywords:
Exponential boson operator , Heisenberg–Weyl algebra , operator ordering , s-ordering , normal ordering , Sheffer sequence , Riordan array
MSC:
[2020] 81S30 , 81S05 , 05A40
1 Introduction
The Heisenberg–Weyl algebra $\mathcal{W}$, generated by ladder operators $a,{a^{\dagger}}$ with canonical commutation relation $a,{a^{\dagger}}]=1$, appears in the modeling of a quantum harmonic oscillator, a single quantized radiation mode of the electromagnetic field, or in general, any quantized, single-mode bosonic field.
Even in a single-mode theory, when calculating quantum-mechanical expectation values of observables, it is important to be able to express the corresponding operators in $\mathcal{W}$, and operators in general (such as the density operator representing the current state, and also non-hermitian operators), in various ordered ways. An example is the exponential operator ${\rm e}^{\lambda({a^{\dagger}}a)}$, an exponentiated version of the number operator ${a^{\dagger}}a$, which when $\lambda<0$ appears as the density operator for a thermal state.
In the then new field of quantum optics, Cahill and Glauber [1] derived as identities in $\mathcal{W}[[\lambda]]$, the space of formal power series in $\lambda$ with coefficients in $\mathcal{W}$, the representations
${\rm e}^{\lambda({a^{\dagger}}a)}=\left\{\begin{aligned} &\mathopen{:}\,\mathrel{\exp\left[\left({\rm e}^{\lambda}-1\right){a^{\dagger}}a\right]}\,\mathclose{:}_{N},\\ &\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}}{a^{\dagger}}a\right]}\,\mathclose{:}_{W},\\ &\mathopen{:}\,\mathrel{{\rm e}^{-\lambda}\exp\left[\left(1-{\rm e}^{\lambda}\right){a^{\dagger}}a\right]}\,\mathclose{:}_{A}.\end{aligned}\right$. (1)
[…]
They also introduced a concept of $s$-ordering $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{s}$, which generalizes the three more familiar types: it involves a free parameter $s\in\mathbb{C}$ (usually, $s\in[-1,1]$), and when s=1,0,+1s=-1,0,+1, $s$-ordering reduces to N-, W-, and A-ordering. They derived the identity
${\rm e}^{\lambda({a^{\dagger}}a)}=\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}{a^{\dagger}}a\right]}\,\mathclose{:}_{s},$ (2)
which subsumes the preceding three.
The $s$-ordering concept has proved useful in quantum mechanics on phase space, and specifically in the quasiprobability approach to quantum optics [2, 3] and other fields of quantum physics [4]. A fundamental result is Sudarshan’s optical equivalence theorem, which makes the quantum theory of light formally resemble a classical probabilistic theory [5]. Suppose a density operator $\hat{\rho}$ for a state has anti-normal ordered representation $\mathopen{:}\,\mathrel{F^{(A)}(\alpha,\alpha^{*})}\,\mathclose{:}_{A}$ (which means that $F^{(A)}$ is some function on the complex $\alpha$-plane, real due to hermiticity, and that anti-normal quantization of $F^{(A)}$ yields $\hat{\rho}$). Then, $F^{(A)}$ can be viewed as a quasiprobability density on $\mathbb{C}\ni\alpha$, from which quantum-mechanical expectations of normal ordered operators, with respect to the state, can be computed by integration in a formally classical way.
This $F^{(A)}$, which is obtained (by setting $s=\nobreak+1$) from a general $s$-parametrized representation of $\hat{\rho}$, which is denoted by $F^{(s)}$, may not be non-negative on the $\alpha$-plane; in which case it cannot be interpreted as a classical probability density, and the state is said to be non-classical. But non-classicality can be quantified: as $s$ is lowered from +1+1 toward 0 and 1-1, regions of negativity (if any) of $F^{(s)}$ and its marginals shrink at some rate and vanish [6].
Quantum optics models in which the time evolution of $s$-parametrized quasiprobability densities has been studied include the widely applicable Jaynes–Cummings model of a two-level atom (a qubit) interacting with a single light mode of a cavity [7], and the Drummond–Walls model of dispersive optical bistability [8]. Plots of evolving densities, at multiple alternative values of $s$, reveal much about model dynamics.
In this paper, we go beyond previous work on the foundations of $s$-ordering by finding closed-form, analytic expressions of $s$-ordered representations denoted by $F^{(s)}$ (i.e., $s$-dependent functions on the complex $\alpha$-plane), for certain single-mode exponential boson operators of the form ${\rm e}^{\lambda\omega}$, where $\omega\in\mathcal{W}$. That is, for each $s$ we find an $F^{(s)}$ with $s$-quantization equal to $\omega$.
The case most amenable to an analytic treatment is when $\omega$ is a monomial or “boson string”: a product of $a$’s and $a^{\dagger}$}’s, i.e., a word over $\{a,{a^{\dagger}}\}$. The “single annihilator” word $\omega={a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ will be treated, where $e=L+\nobreak R-\nobreak 1\geqslant 0$ is the excess of $a^{\dagger}$}’s over $a$’s. The Cahill–Glauber result, eq. (2) above, is the $s$-representation of ${\rm e}^{\lambda\omega}$ when $(L,R)=(1,0)$ and $\omega={a^{\dagger}}a$.
The chief tool is the theory of Sheffer polynomials [9, 10], or equivalently the theory of exponential Riordan arrays (infinite triangular arrays of polynomial coefficients) [11, 12, 13]. A Sheffer polynomial sequence is a sequence of polynomials $s_{n}(t)$ with $s_{0}(t)\equiv 1$ and $\deg s_{n}=n$, $n\geqslant 0$, which can be constructed in any of several ways, including the use of differential ladder operators that increment and decrement $n$. By extending previous work of Lang [14, 15] and Błasiak and collaborators [16, 17], we first derive normal and anti-normal representations that subsume the first and third identities in (1), namely
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,-e,1,R)}\,\mathclose{:}_{N}\\ &=\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl{(}1-e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}\end{aligned}$ (3)
and
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}(a{a^{\dagger}},\lambda{a^{\dagger}}^{e};\,e,-1,-L)}\,\mathclose{:}_{A}\\ &=\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl{(}1+e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A},\end{aligned}$ (4)
[…]
Our main result (Theorem 5.3) is the explicit $s$-ordering of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$, valid for arbitrary $s\in\mathbb{C}$. It subsumes eqs. (3) and (4), which are its $s=\nobreak-1$ and $s=\nobreak+1$ special cases, and includes a new and explicit Weyl ordering formula ($s=\nobreak 0$); for the latter, see Theorem 5.5. The general $s$-ordered representation is expressed in terms of what will be called a two-point Hsu–Shiue EGF, interpolating in an intricate sense between the EGF’s in (3) and (4).
Many examples, including a full treatment of the case $e=1$, are given. However, if $e\geqslant 2$ the two-point EGF is obtained by solving a high-degree polynomial equation (if $e=2$, a quartic), and may not be expressible in terms of radicals.
Any ordering of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ gives an ordering of $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}$ by expanding in $\lambda$. The normal or anti-normal ordering of the $n$’th power of a boson string can be accomplished with the aid of combinatorics [18], or algorithmically [19]. In Proposition 6.2 we give a Weyl-ordered representation of $({a^{\dagger}}a{a^{\dagger}})^{n}$, derived in a different way that exploits a known Riordan array.
The paper is structured as follows. In Sec. 2, operator orderings including $s$-ordering are defined, in part using quantum optics language. Sheffer polynomial sequences and exponential Riordan arrays are introduced in Sec. 3, and the parametric Hsu–Shiue family is defined. Theorem 4.4 subsumes eqs. (3) and (4) above, and serves as a lemma in the derivation by interpolation of the main result, in Sec. 5. Examples are given in Sec. 6.
2 Definitions and $s$-ordering
2.1 Definitions
The single-mode Heisenberg–Weyl algebra $\mathcal{W}$ is an associative algebra over the scalar field $\mathbb{C}$ generated by noncommuting indeterminates $a,{a^{\dagger}}$. Any element of $\mathcal{W}$ can be written as a polynomial in $a,{a^{\dagger}}$, i.e., a linear combination over $\mathbb{C}$ of finite products of $a,{a^{\dagger}}$. $\mathcal{W}$ has a unit (multiplicative identity), written $1$, which is the empty product multiplied by the scalar $1$.
The representation as polynomials in $a,{a^{\dagger}}$ is not unique, as the product on $\mathcal{W}$ satisfies $a,{a^{\dagger}}]=1$. $\mathcal{W}$ can be identified with $\mathbb{C}\langle a,{a^{\dagger}}\rangle/\allowbreak(a{a^{\dagger}}-\nobreak{a^{\dagger}}a-\nobreak 1)$, the quotient of the free algebra $\mathbb{C}\langle a,{a^{\dagger}}\rangle$ by the two-sided ideal generated by $a{a^{\dagger}}-\nobreak\nobreak{a^{\dagger}}a-\nobreak 1$. Thus each element of $\mathcal{W}$ is an equivalence class of polynomials in $a,{a^{\dagger}}$, though a distinguished representative can be singled out by imposing an ordering criterion.
One can also view $\mathcal{W}$ as $\mathbb{C}[{a^{\dagger}}][a]$, the ring of polynomials in $a$ with coefficients that are polynomials in $a^{\dagger}$}, equipped with a product operation that comes from $a,{a^{\dagger}}]=1$. In effect, this imposes normal ordering. Alternatively, one can view $\mathcal{W}$ as the isomorphic algebra $\mathbb{C}[a][{a^{\dagger}}]$, which imposes anti-normal ordering. The isomorphism $\mathbb{C}[{a^{\dagger}}][a]\leftrightarrow\mathbb{C}[a][{a^{\dagger}}]$ is also based on $a,{a^{\dagger}}]=1$: any polynomial can be written in either normal or anti-normal ordered form.
So, any $G\in\mathcal{W}$ can be viewed as a finite sum over $\mathbb{C}$ of terms of the type ${a^{\dagger}}^{n}a^{m}$, or equivalently one of terms of the type $a^{m}{a^{\dagger}}^{n}$. To avoid the problem of convergence when dealing with ${\rm e}^{G}$, which is formally an infinite sum $\sum_{n=0}^{\infty}G^{n}/n!$ and is not an element of $\mathcal{W}$, one can embed ${\rm e}^{G}$ in the one-parameter family ${\rm e}^{\lambda G}\in\mathcal{W}[[\lambda]]$, a power series in an indeterminate $\lambda$ with coefficients in $\mathcal{W}$.
It is useful to define an extension algebra $\overline{\mathcal{W}}{(a)}$ of $\mathcal{W}$, each element of which is a sum of ${a^{\dagger}}^{n}a^{m}$ terms, finite or infinite, but with the exponent of $a$ taking only a finite number of values in all. (The product on $\overline{\mathcal{W}}{(a)}$ comes from $a,{a^{\dagger}}]=1$.) One can equally well say that each element is a sum of $a^{m}{a^{\dagger}}^{n}$ terms, finite or infinite, with the same condition imposed. Similarly, one defines the extension algebra $\overline{\mathcal{W}}{({a^{\dagger}})}$ of $\mathcal{W}$, where it is the exponent of $a^{\dagger}$} that can take only a finite number of values. $\overline{\mathcal{W}}{(a)}$, $\overline{\mathcal{W}}{({a^{\dagger}})}$ could be denoted by $\mathbb{C}[[{a^{\dagger}}]][a]$, $\mathbb{C}[[a]][{a^{\dagger}}]$. The power-series algebras $\mathbb{C}[[{a^{\dagger}}]][a][[\lambda]]$, $\mathbb{C}[[a]][{a^{\dagger}}][[\lambda]]$ will also appear.
2.2 Orderings
The following defines $s$-ordering as an interpolation between normal and anti-normal ordering [1, 2, 3]. It was introduced by Cahill–Glauber [1] in quantum optics, where $a,{a^{\dagger}}$ are ladder (annihilation and creation) operators for a quantized radiation field. But the treatment here is purely algebraic and does not single out a representation of $\mathcal{W}$, such as one on a Fock space. So it makes no mention of coherent states or number (i.e., Fock) states, and though the Glauber displacement operator appears as an algebraic object, it will not act on anything.
Let $G\in\mathcal{W}$, i.e., let $G(a,{a^{\dagger}})$ be a polynomial in the noncommuting $a,{a^{\dagger}}$ that represents $G$ (and is not unique). One can define two c-number or classical representations (sometimes called transforms) associated to $G$, denoted by $F^{(N)},F^{(A)}\in\mathcal{W}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$, the ring of polynomials in commuting indeterminates $\alpha,\alpha^{*}$, which is $\mathbb{C}\langle\alpha,\alpha^{*}\rangle/\allowbreak(\alpha\alpha^{*}-\nobreak\alpha^{*}\alpha)$. They are defined by requiring that
$G(a,{a^{\dagger}})$ $=\mathopen{:}\,\mathrel{F^{(N)}(\alpha,\alpha^{*})}\,\mathclose{:}_{N},$ (5a)
$G(a,{a^{\dagger}})$ $=\mathopen{:}\,\mathrel{F^{(A)}(\alpha,\alpha^{*})}\,\mathclose{:}_{A},$ (5b)
as equalities in $\mathcal{W}$. In this, the normal and anti-normal double dot operations $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{N}$ and $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{A}$, each linear over $\mathbb{C}$ in its single argument, map $\mathcal{W}_{c}$ to $W$ by quantizing any monomial ${\alpha^{*}}^{n}\alpha^{m}$ in $\mathcal{W}_{c}$ to the elements ${a^{\dagger}}^{n}a^{m}$ and $a^{m}{a^{\dagger}}^{n}$ of $\mathcal{W}$, respectively. Thus the inverse images $F^{(N)},F^{(A)}\in\mathcal{W}_{c}$ are standardized, ordered representations of $G\in\mathcal{W}$.
[…]
In the same way, one defines a standardized Weyl-ordered representation $F^{(W)}\in\mathcal{W}_{c}$ of any $G\in\mathcal{W}$ by requiring that $G(a,{a^{\dagger}})$ equal $\mathopen{:}\,\mathrel{F^{(W)}(\alpha,\alpha^{*})}\,\mathclose{:}_{W}$. Here, the Weyl quantization $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{W}$ maps any monomial ${\alpha^{*}}^{n}\alpha^{m}$ to the average of the $\binom{n+m}{n}=\binom{n+m}{m}$ distinct orderings of a product of $n$ $a^{\dagger}$}’s and $m$ $a$’s. (An equivalent definition is that in $\mathcal{W}$, $\mathopen{:}\,\mathrel{(\mu\alpha+\nobreak\nu\alpha^{*})^{p}}\,\mathclose{:}_{W}$ must equal $(\mu a+\nobreak\nu{a^{\dagger}})^{p}$.) For instance,
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{2}\alpha}\,\mathclose{:}_{W}=({a^{\dagger}}^{2}a+{a^{\dagger}}a{a^{\dagger}}+a{a^{\dagger}}^{2})/3,$ (6)
where the right-hand side, an element of $\mathcal{W}$, can be rewritten with the aid of $a,{a^{\dagger}}]=1$ in various ways, including ${a^{\dagger}}^{2}a+\nobreak{a^{\dagger}}$, ${a^{\dagger}}a{a^{\dagger}}$, $a{a^{\dagger}}^{2}-\nobreak{a^{\dagger}}$. So
$G(a,{a^{\dagger}})={a^{\dagger}}a{a^{\dagger}}\quad\Longrightarrow\quad\left\{\begin{aligned} F^{(N)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha+\alpha^{*},\\ F^{(W)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha,\\ F^{(A)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha-\alpha^{*},\end{aligned}\right$. (7)
and if $G(a,{a^{\dagger}})={a^{\dagger}}^{2}a$, then $\alpha^{*}$ would be subtracted from each $F(\alpha,\alpha^{*})$, etc.
It is useful to define a formal power series $D(\beta,\beta^{*})\in\mathcal{W}[[\beta,\beta^{*}]]$ by
$D(\beta,\beta^{*})$ $=\exp(\beta{a^{\dagger}}-\beta^{*}a)=\sum_{p=0}^{\infty}\frac{1}{p!}(\beta{a^{\dagger}}-\beta^{*}a)^{p}$ (8)
$=\sum_{n,m=0}^{\infty}\frac{\beta^{n}(-\beta^{*})^{m}}{n!\,m!}\,\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{W},$
so that
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{W}=\frac{\partial^{n+m}}{\partial\beta^{n}\partial(-\beta^{*})^{m}}D(\beta,\beta^{*})\Bigm{|}_{(\beta,\beta^{*})=(0,0)}.$ (9)
In quantum optics, $D(\beta,\beta^{*})$ appears as the displacement operator.
In such equations as the preceding, a notational convention common in quantum optics is adopted: paired quantities such as $\alpha,\alpha^{*}$ and $\beta,\beta^{*}$, which in some contexts are treated as a complex number and its conjugate, are treated in others as being independent. (Not least, in the exponentiation of $\mathcal{W}$ to the complexified Heisenberg–Weyl group [20].) For greater rigor, one could instead introduce commuting quantities $\alpha_{r},\alpha_{i}$ and $\beta_{r},\beta_{i}$, the real and imaginary parts of $\alpha$ and $\beta$, to serve as real coordinates on the complex $\alpha$- and $\beta$-planes [2].
Define an $s$-parametrized deformation of the exponential operator $D(\beta,\beta^{*})$, namely $D(\beta,\beta^{*};s)={\rm e}^{s\beta(-\beta^{*})/2})D(\beta,\beta^{*})$, where $s\in\mathbb{C}$, and for each $n,m$, define the $s$-quantization $\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}$ of ${\alpha^{*}}^{n}\alpha^{m}$ by
$D(\beta,\beta^{*};s)=\sum_{n,m=0}^{\infty}\frac{\beta^{n}(-{\beta^{*}})^{m}}{n!\,m!}\,\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s},$ (10)
so that
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}=\frac{\partial^{n+m}}{\partial\beta^{n}\partial(-\beta^{*})^{m}}D(\beta,\beta^{*};s)\Bigm{|}_{(\beta,\beta^{*})=(0,0)}.$ (11)
Equivalently, one requires that
$D(\beta,\beta^{*})=\exp(\beta{a^{\dagger}}-\beta^{*}a)={\rm e}^{s|\beta|^{2}}\mathopen{:}\,\mathrel{{\rm e}^{\beta\alpha^{*}-\beta^{*}\alpha}}\,\mathclose{:}_{s},$ (12)
as an equality in $\mathcal{W}[[\beta,\beta^{*}]]$.
As an example of (11), one finds with the aid of $a,{a^{\dagger}}]=1$ that
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{2}\alpha}\,\mathclose{:}_{s}={a^{\dagger}}a{a^{\dagger}}+s\,{a^{\dagger}}.$ (13)
The $s$-quantization operator $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{s}$ extends by linearity from monomials in the (commuting) indeterminates $\alpha,\alpha^{*}$, to polynomials in same, i.e., to elements of ${\mathcal{W}}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$, and to elements of $\mathcal{W}_{c}[[\lambda]]=\mathbb{C}[\alpha,\alpha^{*}][[\lambda]]$, as well.
Thus, one can define a unique $s$-ordered representation or $s$-transform of any $G=G(a,{a^{\dagger}})$ in $\mathcal{W}$, denoted by $F^{(s)}(\alpha,\alpha^{*})$, by requiring that
$G(a,{a^{\dagger}})=\mathopen{:}\,\mathrel{F^{(s)}(\alpha,\alpha^{*})}\,\mathclose{:}_{s}.$ (14)
So, for example,
$G(a,{a^{\dagger}})={a^{\dagger}}a{a^{\dagger}}\quad\Longrightarrow\quad F^{(s)}(\alpha,\alpha^{*})={\alpha^{*}}^{2}\alpha-s\,\alpha^{*}.$ (15)
When $s=\nobreak 0$, $F^{(s)}$ reduces to the Weyl-ordered representation $F^{(W)}$. It is also not difficult to show, using the Baker–Campbell–Hausdorff formula
${\rm e}^{A}{\rm e}^{B}={\rm e}^{A+B+[A,B]/2},$ (16)
which holds as an equality between power series if $[A,B]$ commutes with $A$ and $B$, that when $s=\nobreak-1$ and $s=\nobreak+1$, $F^{(s)}$ reduces respectively to the normal and anti-normal ordered representations, $F^{(N)}$ and $F^{(A)}$. The s=1,0,+1s=-1,0,+1 cases of (15) are reflected in eq. (7).
In modern quantum optics, the most commonly encountered c-number representations of any element of the Heisenberg–Weyl algebra are the classical functions $F^{(-1)}=F^{(N)}$, $F^{(0)}=F^{(W)}$, $F^{(+1)}=F^{(A)}$ on $\mathbb{C}$ that arise from $G=\hat{\rho}$, where $\hat{\rho}$ is a specified (single-mode) density operator that is an element of $\mathcal{W}$, or an extension of $\mathcal{W}$ that includes formal power series in $a,{a^{\dagger}}$ [2, 21, 22]. These are called the $Q$, $W$, and $P$ quasiprobability densities, and are attributed respectively to Husimi, Wigner, and Glauber–Sudarshan. But in older works on phase-space methods in quantum optics, operators other than density operators and observables were frequently “classicized,” and it was in that general context that $s$-ordering was introduced. (See [23, 24, 25], and more recently [26]. [^1] There are two conventions on the sign of $s$. In the original paper of Cahill–Glauber [1] and a few more recent ones [26, 27, 28], s=+1,1s=+1,-1 correspond to normal and anti-normal ordering; e.g., for a density operator, to $Q$ and $P$. But usually, they correspond to anti-normal and normal ordering (i.e., to $P$ and $Q$). See, e.g., [2, Chapter 12], [6, 29], and [3, Chapter 3]. The latter convention has been adopted here, so the sign of $s$ in eq. (2), taken from [1], has been changed. Note also that in some works (e.g., [21] and [24, Table V]), the meanings of the so-called reciprocal orderings $F^{(N)}$ and $F^{(A)}$, related by negation of $s$, are interchanged.)
Differentiating the exponential operator $D(\beta,\beta^{*};s)={\rm e}^{(s^{\prime}-s)|\beta|^{2}/2}D(\beta,\beta^{*};s^{\prime})$ with respect to $\beta,\beta^{*}$, in a procedure that can be traced to Wilcox [30], and using (11), yields the $s$-ordering conversion formula
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}=\sum_{k=0}^{\min(m,n)}k!\,\binom{n}{k}\binom{m}{k}\left(\frac{s-s^{\prime}}{2}\right)^{k}\mathopen{:}\,\mathrel{{\alpha^{*}}^{n-k}\alpha^{m-k}}\,\mathclose{:}_{s^{\prime}}.$ (17)
The simplest case of (17) is the contraction
$\mathopen{:}\,\mathrel{\alpha^{*}\alpha}\,\mathclose{:}_{s}-\mathopen{:}\,\mathrel{\alpha^{*}\alpha}\,\mathclose{:}_{s^{\prime}}=\frac{s-s^{\prime}}{2},$ (18)
from which the general formula (17) can be derived, much as Wick’s theorem is derived [28]. As another example, the $s=\nobreak+1$, $s^{\prime}=-1$ case of (17) is
$a^{m}{a^{\dagger}}^{n}=\sum_{k=0}^{\min(m,n)}k!\,\binom{n}{k}\binom{m}{k}{a^{\dagger}}^{n-k}a^{m-k},$ (19)
an identity in $\mathcal{W}$ that converts an anti-normal ordering ($s=\nobreak+1$) to a normal one ($s=\nobreak-1$). Thus one can convert between the coefficients of an anti-normal and a normal series, $\sum_{m,n}f_{m,n}^{(A)}a^{m}{a^{\dagger}}^{n}$ and $\sum_{m,n}f_{n,m}^{(N)}{a^{\dagger}}^{n}a^{m}$ [31].
Equation (17) can be written in terms of a two-variable Hermite polynomial [27, 28], as was recognized early (when $s\in\{0,\pm 1\}$) by Agarwal and Wolf [24, Table V]. It can also be written as
$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}$ $=\sum_{k=0}^{\min(m,n)}\frac{1}{k!}\left(\frac{s-s^{\prime}}{2}\right)^{k}\mathopen{:}\,\mathrel{\left(\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right)^{k}{{\alpha^{*}}^{n}\alpha^{m}}}\,\mathclose{:}_{s^{\prime}}$ (20)
$=\mathopen{:}\,\mathrel{\exp\left[\left(\frac{s-s^{\prime}}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]{{\alpha^{*}}^{n}\alpha^{m}}}\,\mathclose{:}_{s^{\prime}}.$
This extends by linearity to
$\mathopen{:}\,\mathrel{F}\,\mathclose{:}_{s}=\mathopen{:}\,\mathrel{\exp\left[\left(\frac{s-s^{\prime}}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]F}\,\mathclose{:}_{s^{\prime}}$ (21)
for any $F\in\mathcal{W}_{c}$, i.e., for any polynomial $F=F(\alpha,\alpha^{*})$.
Suppose now that $F^{(s)},F^{(s^{\prime})}\in{\mathcal{W}}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$ are the $s$-ordered and $s^{\prime}$-ordered “c-number” versions of some $G\in\mathcal{W}$, computed from $G$ by the requirement (14). By (20), they must be related to each other by the dual relation
$F^{(s)}(\alpha,\alpha^{*})=\exp\left[\left(\frac{s^{\prime}-s}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]F^{(s^{\prime})}(\alpha,\alpha^{*}),$ (22)
in which the exponentiated second-order differential operator has been inverted. In the formally infinite series in (22) obtained by expanding the exponential in a power series, only a finite number of terms will be nonzero. Equivalently, $F=F^{(s)}(\alpha,\alpha^{*})$ satisfies the diffusion or heat equation
$\frac{\partial F}{\partial(-s)}=\frac{1}{2}\,\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}F,$ (23)
an unexponentiated version of (22). (See [2, Chapter 12].)
Equation (22) can be extended to the case when $F^{(s)},F^{(s^{\prime})}$ are convergent power series in $\alpha,\alpha^{*}$ rather than polynomials, at least formally. The nicest case is when each of $F^{(s)},F^{(s^{\prime})}$, regarded as a Maclaurin expansion at $(\alpha,\alpha^{*})=(0,0)$, i.e., $(\alpha_{r},\alpha_{i})=(0,0)$, converges globally and defines an analytic function of the real variables $\alpha_{r},\alpha_{i}$, or even a complex analytic function of $\alpha=\alpha_{r}+\nobreak{\rm i}\alpha_{i}$. Because $\partial^{2}/\partial\alpha\partial\alpha^{*}$ equals $\Delta/4$, where $\Delta=\partial^{2}/\partial\alpha_{r}^{2}+\partial^{2}/\partial\alpha_{i}^{2}$ is the Laplacian on the complex $\alpha$-plane coordinatized by $\alpha_{r},\alpha_{i}$, one has
$F^{(s)}(\alpha,\alpha^{*})$ $={\rm e}^{(s^{\prime}-s)\Delta/8}F^{(s^{\prime})}(\alpha,\alpha^{*})$ (24)
$\varpropto\int F^{(s^{\prime})}(\beta,\beta^{*})\,{\rm e}^{-2|\beta-\alpha|^{2}/(s^{\prime}-s)}\,{\rm d}^{2}\beta.$
So in this case, when $s<s^{\prime}$ it may be possible to compute $F^{(s)}$ from $F^{(s^{\prime})}$ by a smoothing Weierstrass transform operation on the complex $\alpha$-plane: convolution with a Gaussian kernel of variance $(s^{\prime}-\nobreak s)/4$.
This operation is familiar from quantum optics [2, 3, 4, 20]: among the quasiprobability densities $F^{(s)}$ obtained when $G=G(a,{a^{\dagger}})$ is a density operator $\hat{\rho}$, the Wigner $W$ function ($s=\nobreak 0$) is a smoothed version of the Glauber–Sudarshan $P$ function ($s=\nobreak+1$), and the Husimi $Q$ function ($s=\nobreak-1$) is a smoothed version of the $W$ function. The smoothing of $P$ to $W$ and $Q$ has been extended from functions of $\alpha\in\mathbb{C}$ to tempered and other distributions on the complex $\alpha$-plane, and the variation with $s$ of the mathematical properties and physical interpretation of smoothed quasiprobability densities has been explored [2, 6, 20].
In the proof of the main result of this paper (Theorem 5.3), the unexponentiated form (23), and not the exponentiated forms (22),(24), will be employed to relate $F^{(s)}$, $F^{(s^{\prime})}\in\mathcal{W}_{c}[[\lambda]]$, to obviate consideration of series convergence.
3 Riordan, Sheffer, and Hsu–Shiue polynomials
The following definitions and facts are standard. For the theory of (exponential) Riordan arrays and polynomials, introduced under that name by Shapiro et al. to derive combinatorial identities, see [11, 12, 13]. For the equivalent theory of Sheffer polynomial sequences (originally called Sheffer of $A$-type zero), which are used in approximation theory and elsewhere, see [9, 10]. In their present form, both emerged from the work of G.-C. Rota and collaborators on the refashioning of the classical umbral calculus as a finite operator calculus [32, 33].
Hsu and Shiue [34] defined a parametric family of generalized Stirling numbers, which now have their own literature [35, Chapter 4]. For any choice of parameters, they can be viewed as the elements of a exponential Riordan array, the “row polynomials” of which are Sheffer polynomials that generalize the classical Touchard polynomials, which have Stirling numbers as coefficients.
3.1 Exponential Riordan arrays and polynomials
Let $d,h\in\mathbb{C}[[z]]$ be formal power series in $z$ of order 0 and $1$ respectively, meaning that their first nonzero terms are the $z^{0}$ and $z^{1}$ terms; so $1/d$ and the compositional inverse $\bar{h}$ both exist as formal power series. (For instance, $d,h$ may be Maclaurin expansions of functions $d(z),h(z)$ analytic at $z=0$.) One usually requires that $d$ be monic, $d=1\cdot z^{0}+\nobreak O(z^{1})$, and often that $h=1\cdot z^{1}+\nobreak O(z^{2})$. Then, a polynomial sequence $s_{n}(t)$, $n\geqslant 0$, where $\deg s_{n}=n$, with EGF $\mathcal{S}(t,z)=d(z){\rm e}^{th(z)}$ that is an element of $\mathbb{C}[t][[z]]$, is defined by
$\sum_{n=0}^{\infty}s_{n}(t)\frac{z^{n}}{n!}=\mathcal{S}(t,z),$ (25)
with $s_{0}(t)\equiv 1$. The sequence $s_{n}(t)$ is said to be exponential Riordan for the ordered pair $[d,h]$, and is denoted briefly by $\mathbf{R}[d,h]$. Such notations as $s_{n}(t)=\mathbf{R}[d,h]_{n}(t)$ and $\mathcal{S}(t,z)=\mathbf{R}[d,h](t,z)$ are also used.
Equivalently, one can define a triangular array of coefficients $s_{n,k}$, $0\leqslant k\leqslant n$, by
$\sum_{n=0}^{\infty}\sum_{k=0}^{n}s_{n,k}t^{k}\frac{z^{n}}{n!}=\mathcal{S}(t,z),$ (26)
so that
$s_{n,k}=\frac{n!}{k!}\,[z^{n}]\left[d(z)h(z)^{k}\right],$ (27)
where $z^{n}]$ extracts the coefficient of $z^{n}$ from the element of $\mathbb{C}[[z]]$ that follows. Then $s_{n}(t)=\sum_{k=0}^{n}s_{n,k}t^{k}$, so the $s_{n}(t)$ are the row polynomials of the array. The array $s_{n,k}$ can be viewed as the lower-triangular part of an infinite matrix, and is called an exponential Riordan array. The notation $s_{n,k}=\mathbf{R}[d,h]_{n,k}$ is also used, with $\mathbf{R}[d,h]$ alternatively signifying the matrix as a whole.
Row-finite lower-triangular matrices form a group under matrix multiplication, and exponential Riordan arrays turn out to form a subgroup, which will be denoted by $\mathfrak{R}$. The following is a fundamental fact.
Proposition 3.1 ([13]).
In the exponential Riordan group $\mathfrak{R}$ the product operation is $\mathbf{R}[d_{1},h_{1}]\,\mathbf{R}[d_{2},h_{2}]=\allowbreak\mathbf{R}[(d_{2}\circ\nobreak h_{1})d_{1},\allowbreak h_{2}\circ\nobreak h_{1}]$, and the inversion operation is $\mathbf{R}[d,h]^{-1}=\allowbreak\mathbf{R}[1/(d\circ\nobreak\bar{h}),\bar{h}]$, where $\bar{h}$ is the compositional inverse of $h$, i.e., $\bar{h}(h(z))=\nobreak z$. The multiplicative identity is $[1,z]$, which signifies the infinite identity matrix.
For matrix inverses, the notation
$\mathbf{R}[d,h]^{-1}\mathrel{=:}\mathbf{R}\overline{[d,h]}=\mathbf{R}[\bar{d},\bar{h}],$ (28)
where $\bar{d}=1/(d\circ\bar{h})$, can be used. Note that $\bar{d},\bar{h}$, like $d,h$, are formal power series of orders 0 and $1$.
The generating functions of the theory facilitate the computation of matrix–vector products.
Proposition 3.2 ([13]).
For any column vector $u=(u_{k})$ with EGF $u(t)=\sum_{k=0}^{\infty}u_{k}t^{k}/k!$, an element of $\mathbb{C}[[t]]$, and for any exponential Riordan array $s_{n,k}$ equaling $\mathbf{R}[d,h]_{n,k}$, define the column vector $v=(v_{n})$ by
$v_{n}=\sum_{k=0}^{n}s_{n,k}u_{k},$ (29)
or symbolically by $v=\mathbf{R}[d,h]u$. Then, its EGF $v(z)=\sum_{n=0}^{\infty}v_{n}z^{n}/n!$, an element of $\mathbb{C}[[z]]$, equals $d(z)u(h(z))=[d\cdot(u\circ h)](z)$, where $\cdot$ and $\circ$ are the binary operations of product and composition on formal power series.
3.2 Sheffer polynomials
The theories of Sheffer and exponential Riordan polynomials are equivalent, being related by the inversion operation [36, 37]. A polynomial sequence $s_{n}(t)$, $n\geqslant 0$, where $\deg s_{n}=n$ and usually $s_{0}(t)\equiv 1$, is said to be Sheffer for the ordered pair $[g,f]$ if and only if it is exponential Riordan for the pair $[d,h]$ equal to $\overline{[g,f]}$, the inverse of $[g,f]$ in the group $\mathfrak{R}$, which will be written as $\bar{g},\bar{f}]$. The sequence as a whole will be denoted briefly by $\mathbf{S}[g,f]$, or equivalently by $\mathbf{R}[d,h]=\mathbf{R}[\bar{g},\bar{f}]$.
It is convenient in the Sheffer theory to use the indeterminate $D$ in the series $g,f$, and $z$ in the series $\bar{g},\bar{f}$. That is, to reduce ambiguity one adopts the convention that $g,f\in\mathbb{C}[[D]]$ and $\bar{g},\bar{f}\in\mathbb{C}[[z]]$. This makes it possible to write both $z=f(D)$ and $D=\bar{f}(z)$, related by series reversion.
So a polynomial sequence $s_{n}$ is Sheffer for $[g,f]$, where $g,f\in\mathbb{C}[[D]]$ are of order 0 and $1$ respectively, and $\bar{g},\bar{f}\in\mathbb{C}[[z]]$ are likewise, if
$\sum_{n=0}^{\infty}s_{n}(t)\frac{z^{n}}{n!}=\bar{g}(z){\rm e}^{t\bar{f}(z)}=\frac{1}{g(\bar{f}(z))}{\rm e}^{t\bar{f}(z)},$ (30)
and one writes $s_{n}=\mathbf{S}[g,f]_{n}=\mathbf{R}[\bar{g},\bar{f}]_{n}$ and $s_{n,k}=\mathbf{S}[g,f]_{n,k}=\mathbf{R}[\bar{g},\bar{f}]_{n,k}$, etc.
Many well-known polynomial sequences can be viewed as Sheffer sequences [9, 38], and there is some overlap with the orthogonal polynomials of theoretical physics. An example is the sequence of Hermite polynomials $\mathop{{}\it He}\nolimits_{n}(t)$, which is $\mathbf{S}[{\rm e}^{D^{2}/2},D]$ or equivalently $\mathbf{R}[{\rm e}^{-z^{2}/2},z]$. The $\mathop{{}\it He}\nolimits_{n}(t)$ are orthogonal with respect to an appropriate measure (namely, ${\rm e}^{-t^{2}/2}\,{\rm d}t$), as they happen to satisfy a three-term recurrence relation. The squared Hermite polynomials $\mathop{{}\it He}\nolimits_{n}(\sqrt{t})]^{2}$ turn out to be $\mathbf{S}[\sqrt{1-\nobreak 2D}/\allowbreak(1-\nobreak D),\allowbreak D/(1-\nobreak D)]$ or $\mathbf{R}[1/\sqrt{1-\nobreak z^{2}},\allowbreak z/\allowbreak(1+\nobreak z)]$, but do not satisfy such a relation.
For a Sheffer sequence $\mathbf{S}[g,f]$, the formal power series $z=f(D)$ may be the Maclaurin expansion of an elementary function of $D$, without the compositional inverse $D=\bar{f}(z)$ appearing in the EGF (30) having the same property. An example is the sequence of Abel polynomials $A_{n}(t)=t(t-\nobreak n)^{n-1}$, which turns out to be ${\bf S}[1,D{\rm e}^{D}]$. (See [9].) But the compositional inverse of $z=f(D)=D{\rm e}^{D}$ is the Lambert $W$ function $D=\bar{f}(z)=\allowbreak W_{L}(z)=\allowbreak z-\nobreak z^{2}+\nobreak\frac{3}{2}z^{3}+\nobreak\cdots$, which is not an elementary function. It is more convenient to refer to the Abel sequence as ${\bf S}[1,D{\rm e}^{D}]$ than as ${\bf R}[1,W_{L}(z)]$. The same phenomenon will appear in Sec. 5.
The following reveals the deep connection between the Sheffer theory and the Heisenberg–Weyl algebra. It also clarifies why the seemingly arbitrary symbol ‘$D$’ is used here as a power-series variable.
Proposition 3.3 ([9, 17, 39]).
For any pair $g,f\in\mathbb{C}[[D]]$ of orders $0,1$ that defines a Sheffer sequence $s_{n}=\mathbf{S}[g,f]_{n}$ of polynomials in $t$, the Heisenberg–Weyl algebra generated by $P,M$ with $[P,M]=1$ has a representation on $\mathbb{C}[t]$, the space of polynomials in $t$, given by lowering and raising operators
$P=g(D),\qquad M=\left[t-\frac{g^{\prime}(D)}{g(D)}\right]\frac{1}{f^{\prime}(D)},$
the action of which on the Sheffer polynomials is
$Ps_{n}=ns_{n-1},\qquad Ms_{n}=s_{n+1}.$ (31)
Here, $D$ is interpreted as $D_{t}={\rm d}/{\rm d}t$, and though $P,M$ may be of infinite order in $D$, as differential operators they map $\mathbb{C}[t]$ into itself.
The operators $P,M$ are called monomiality operators, and under a suitable invertible linear transformation of $\mathbb{C}[t]$, the polynomials $s_{n}$ can be taken to the monomials $t^{n}$, and $P,M$ to $D_{t},t$. (See [39] and references therein.) Another differential statement involving $D=D_{t}$ is
$t^{0}]\,g(D_{t}){\rm e}^{\lambda f(D_{t})}s_{n}(t)=\lambda^{n},$
an equality in $\mathbb{C}[t][[\lambda]]$ which like (31) can be used as an alternative definition of the polynomials $s_{n}$ [9].
The exponentiated raising operator ${\rm e}^{\lambda M}$ is of interest. Each non-negative power $M^{n}$ is an element of $\mathbb{C}[t][[D_{t}]]$, and because $M^{n}s_{0}=s_{n}$ and $s_{0}(t)\equiv 1$, it follows from (30) that
$\left[{\rm e}^{\lambda M}1\right](t)=\bar{g}(\lambda){\rm e}^{t\bar{f}(\lambda)},$ (32)
an equality in $\mathbb{C}[t][[\lambda]]$. The right-hand side is the Sheffer EGF at $z=\lambda$.
For later reference, note that replacing $D,t$ by $a,{a^{\dagger}}$ converts $M$ to an element $\tilde{M}$ of the algebra $\mathbb{C}[[a]][{a^{\dagger}}]$, which is strictly larger than $\mathbb{C}[a][{a^{\dagger}}]$, the realization of $\mathcal{W}$ by anti-normal ordered polynomials in $a,{a^{\dagger}}$.
3.3 Hsu–Shiue arrays and polynomials
Hsu–Shiue arrays $\mathcal{HS}(A,B;r)$, where $A,B,r\in\mathbb{C}$, are a parametric family of exponential Riordan arrays, i.e., a family of elements of the group $\mathfrak{R}$, with associated row polynomials and EGF’s. They can be introduced as follows.
The classical Stirling numbers (of the second kind), denoted by $\left\{{n\atop k}\right\}$ or $S(n,k)$, count the partitions of an $n$-set into $k$ disjoint, nonempty subsets. They make up a triangular array indexed by $n,k$ with $0\leqslant k\leqslant n$, and satisfy
$x^{n}=\sum_{k=0}^{n}\left\{{n\atop k}\right\}x^{\underline{k}},$ (33)
where $x^{\underline{k}}=x(x-1)\dots(x-k+1)$ is the falling factorial with unit “stride.” So, they are coefficients of connection between bases of $\mathbb{C}[x]$. The sequence of row polynomials $\sum_{k=0}^{n}\left\{{n\atop k}\right\}t^{k}$, $n\geqslant 0$, which could be called Stirling polynomials (of the second kind), but are more often called Touchard, Bell, or exponential polynomials, is the exponential Riordan sequence ${\bf R}[1,\allowbreak{\rm e}^{z}-\nobreak 1]$ or equivalently the Sheffer sequence ${\bf S}[1,\allowbreak\log(1+\nobreak D)]$.
The generalized Stirling numbers of Hsu–Shiue [34] are a family $\mathop{{}\it HS}\nolimits_{n,k}=\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)$ that subsumes the classical Stirling numbers and many previous generalizations. For $A,B,r\in\mathbb{C}$, they are defined by
$(x+r)^{\underline{n},A}=\sum_{k=0}^{n}HS_{n,k}(A,B,r)\,x^{\underline{k},B},$ (34)
where $(x+r)^{\underline{n},A}$, $x^{\underline{k},B}$ are falling factorials with stride not necessarily equal to unity, e.g., $x^{\underline{k},B}=x(x-\nobreak B)\dots[x-\nobreak(k-\nobreak 1)B]$. The $\mathop{{}\it HS}\nolimits_{n,k}$ reduce to $\left\{{n\atop k}\right\}$ if $(A,B,r)=(0,1,0)$. The parametric row polynomials
$\mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)t^{k},\qquad n\geqslant 0,$ (35)
will be called Hsu–Shiue polynomials. (Note that $\mathop{{}\it HS}\nolimits_{0}(t)\equiv 1$ and $\deg\mathop{{}\it HS}\nolimits_{n}=n$ for all $n$.)
By examination, this $n$-indexed sequence of polynomials is exponential Riordan for the pair $d_{\textit{HS}}(\cdot;A,r),h_{\textit{HS}}(\cdot;A,B)]$, where if $A,B\neq 0$,
$d_{\textit{HS}}(z;A,r)\mathrel{:=}(1+Az)^{r/A},$ (36)
$h_{\textit{HS}}(z;A,B)\mathrel{:=}\frac{(1+Az)^{B/A}-1}{B},$
the right-hand sides being elements of $\mathbb{C}[[z]]$. (The cases $A=0$, $B=0$ are handled by taking limits.) One way of confirming this is to verify that eq. (34), viewed as an equality between a column vector indexed by $n$ and a matrix–vector product, is consistent with Proposition 3.2.
It is easily checked that
$\overline{[d_{\textit{HS}}(\cdot;A,r),h_{\textit{HS}}(\cdot;A,B)]}={[d_{\textit{HS}}(\cdot;B,-r),h_{\textit{HS}}(\cdot;B,A)]},$ (37)
which is a statement of duality: as infinite lower-triangular matrices and as elements of the exponential Riordan group $\mathfrak{R}$, Hsu–Shiue arrays $\mathcal{HS}(A,B,r)$ and $\mathcal{HS}(B,A,-r)$ are inverses of each other. (Cf. [34].) So, one can write
$\begin{aligned} \mathcal{HS}(A,B,r)&={\bf R}[d_{\textit{HS}}(z;A,r),h_{\textit{HS}}(z;A,B)]\\ &={\bf S}[d_{\textit{HS}}(D;B,-r),h_{\textit{HS}}(D;B,A)]\end{aligned}$ (38)
and
$\begin{aligned} \mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)&={\bf R}[d_{\textit{HS}}(z;A,r),h_{\textit{HS}}(z;A,B)]_{n}(t)\\ &={\bf S}[d_{\textit{HS}}(D;B,-r),h_{\textit{HS}}(D;B,A)]_{n}(t).\end{aligned}$ (39)
Regardless of whether one uses the exponential Riordan or the Sheffer notation to express the $\mathop{{}\it HS}\nolimits_{n}$, the following holds.
Proposition 3.4.
The parametric Hsu–Shiue polynomials $\mathop{{}\it HS}\nolimits_{n}(t)$ have the EGF
$\sum_{n=0}^{\infty}\mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)\frac{z^{n}}{n!}=(1+Az)^{r/A}\exp\left\{\frac{t}{B}\left[(1+Az)^{B/A}-1\right]\right\},$
with the cases $A=0$, $B=0$ handled by taking limits. The EGF can be viewed as an element of $\mathbb{C}[t][[z]]$, or more strongly as an analytic function on a neighborhood of $(t,z)=(0,0)$.
From this EGF, which will be denoted by $\mathcal{HS}(t,z)=\mathcal{HS}(t,z;A,B,r)$, the $\mathop{{}\it HS}\nolimits_{n}(t)$ can be derived by repeated differentiation. For each $k$ with $0\leqslant k\leqslant n$, the coefficient $\mathop{{}\it HS}\nolimits_{n,k}$ of the $k$’th power of $t$ in $\mathop{{}\it HS}\nolimits_{n}(t)$ will be a polynomial in $A,B,r$ with integer coefficients.
If $B\neq 0$, the $\mathop{{}\it HS}\nolimits_{n,k}$ can also be computed from
$\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)=\frac{1}{B^{k}k!}\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}(Bj+r)^{\underline{n},A},$ (40)
which comes from eq. (34) by finite-differencing to extract coefficients [40]. And for any $A,B,r$, the $\mathop{{}\it HS}\nolimits_{n,k}$ can be computed from the recurrence (42), below.
Proposition 3.5.
The following are equivalent characterizations of the EGF
$\mathcal{HS}(t,z;\,A,B,r)=\sum_{n=0}^{\infty}\mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)\frac{z^{n}}{n!}=\sum_{n=0}^{\infty}\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)\,t^{k}\frac{z^{n}}{n!},$
which is an element of $\mathbb{C}[t][[z]]$.
  1. (i)
    The explicit formula for $\mathcal{HS}(t,z;A,B,r)$ given in Proposition 3.4.
  2. (ii)
    The first-order partial differential equation (PDE)
    $\left[(-Az-1)\frac{\partial}{\partial z}+Bt\,\frac{\partial}{\partial t}+(r+t)\right]\mathcal{HS}=0,$ (41)
    on a neighborhood of $(t,z)=(0,0)$, with the initial condition $\mathcal{HS}(t,0)\equiv 1$.
  3. (iii)
    The triangular recurrence
    $\mathop{{}\it HS}\nolimits_{n+1,k+1}=\left[-An+B(k+1)+r\right]\mathop{{}\it HS}\nolimits_{n,k+1}+\mathop{{}\it HS}\nolimits_{n,k},$ (42)
    with the initial condition that the apex element $\mathop{{}\it HS}\nolimits_{0,0}$ equal unity, and the convention that $\mathop{{}\it HS}\nolimits_{n,k}=0$ if $k<0$ or $k>n$.
Proof.
That $\mathcal{HS}(t,z)$ as supplied by Proposition 3.4 satisfies the PDE follows by direct computation. The PDE (with initial condition) can be solved on a neighborhood of $(t,z)=(0,0)$ by the Lagrange–Charpit method of characteristics, yielding the same analytic formula for $\mathcal{HS}(t,z)$. And, it is easily checked that $\mathcal{HS}(t,z)$, a function analytic at $(0,0)$ with bivariate Maclaurin expansion in $\mathbb{C}[t][[z]]$, satisfies the PDE if and only if its expansion coefficients $\mathop{{}\it HS}\nolimits_{n,k}$ satisfy the recurrence (42). ∎
For some choices of $A,B,r$, explicit expressions for the polynomial coefficients $\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)$ and the polynomials $\mathop{{}\it HS}\nolimits_{n}(t;A,B,r)$ can be obtained from the summation formula (40) or the recurrence (42). For instance,
$\mathop{{}\it HS}\nolimits_{n,k}(B,B,r)=\binom{n}{k}r^{\underline{n-k},B},$ (43a)
$\mathop{{}\it HS}\nolimits_{n,k}(-B,B,r)=\binom{n}{k}(Bk+r)^{\overline{n-k},B},$ (43b)
with an obvious notation for the rising factorial. Notice that by (43b),
$\mathop{{}\it HS}\nolimits_{n,k}(-1,1,1)={\binom{n}{k}}\frac{n!}{k!},$ $\mathop{{}\it HS}\nolimits_{n,k}(1,-1,-1)=(-1)^{n-k}{\binom{n}{k}}\frac{n!}{k!},$ (44)
and
$\mathop{{}\it HS}\nolimits_{n}(t;\,-1,1,1)=n!\,L_{n}(-t),$ $\mathop{{}\it HS}\nolimits_{n}(t;\,1,-1,-1)=(-1)^{n}n!\,L_{n}(t),$ (45)
where $L_{n}$ is the $n$’th Laguerre polynomial. By examination, negating the parameters $A,B,r$ of any exponential Riordan array of Hsu–Shiue type multiplies $\mathop{{}\it HS}\nolimits_{n,k}$ by $(-1)^{n-k}$ and alters $\mathop{{}\it HS}\nolimits_{n}(t)$ to $(-1)^{n}\mathop{{}\it HS}\nolimits_{n}(-t)$.
One can see from (39) that the polynomial sequence $n!\,L_{n}(-t)$ is the Sheffer sequence $\mathbf{S}[1/(1+\nobreak D),\allowbreak D/\allowbreak(1+\nobreak D)]$, or equivalently $\mathbf{R}[1/(1-\nobreak z),z/\allowbreak(1-\nobreak z)]$; and $(-1)^{n}n!\,L_{n}(t)$ is $\mathbf{S}[1/(1-\nobreak D),\allowbreak D/\allowbreak(1-\nobreak D)]$ or $\mathbf{R}[1/(1+\nobreak z),z/\allowbreak(1+\nobreak z)]$. (Cf. [9].)
An unrelated fact will be needed below: besides $\mathop{{}\it HS}\nolimits_{n,k}(0,1,0)=\left\{{n\atop k}\right\}$, one can prove by induction that $\mathop{{}\it HS}\nolimits_{n,k}(0,1,1)=\left\{{{n+1}\atop{k+1}}\right\}$.
4 Exponentiated operators
The main result of this paper, on the $s$-ordered representation of the exponential of a single-mode “boson string” of the form ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$, will be deduced in Sec. 5 from a restricted result on a normal ordering; or equally well, an anti-normal one. The normal and anti-normal results, given in Theorem 4.4 below, appeared in the Introduction as eqs. (3),(4). They will follow from Theorem 4.2, a result of the author on the $n$’th power of ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ [41, Thm. 6.5], a short proof of which is supplied, to make this paper self-contained.
As will be sketched, they also follow from the results of Lang [14, 15] and Błasiak et al. [16, 17] on exponentiated vector fields and on applications of Sheffer sequences to ordering problems. That Fan et al. [42] obtained the $R=\nobreak 0$ case of eq. (3) by the formal IWOP (integration within ordered products) technique [43] must also be mentioned.
As motivation for introducing the coefficients of Hsu–Shiue and other Sheffer polynomials, consider the following identities. Firstly, the two in $\mathcal{W}[[\lambda]]$,
${\rm e}^{\lambda({a^{\dagger}}a)}$ $=\mathopen{:}\,\mathrel{\exp\left[\left({\rm e}^{\lambda}-1\right){a^{\dagger}}a\right]}\,\mathclose{:}_{N}$ (46a)
$=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda;\,0,1,0)}\,\mathclose{:}_{N}=\mathopen{:}\,\mathrel{\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,0,1,0)\frac{\lambda^{n}}{n!}}\,\mathclose{:}_{N}$ (46b)
and
${\rm e}^{\lambda(a{a^{\dagger}})}$ $=\mathopen{:}\,\mathrel{{\rm e}^{\lambda}\exp\left[\left({\rm e}^{\lambda}-1\right)a{a^{\dagger}}\right]}\,\mathclose{:}_{N}$ (46a)
$=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda;\,0,1,1)}\,\mathclose{:}_{N}=\mathopen{:}\,\mathrel{\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,0,1,1)\frac{\lambda^{n}}{n!}}\,\mathclose{:}_{N}.$ (46b)
Here, (46a) is the Cahill–Glauber normal ordering shown in (1), and (46a) is an immediate consequence. The rewritings in terms of $\mathcal{HS}(t,z)$ come from the $A\to 0$ limit of the EGF formula in Proposition 3.4. Secondly, the two in $\mathcal{W}$,
$({a^{\dagger}}a)^{n}$ $=\sum_{k=0}^{n}\left\{{{n}\atop{k}}\right\}{a^{\dagger}}^{k}a^{k}=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(0,1,0)\,{a^{\dagger}}^{k}a^{k}$ (47)
and
$(a{a^{\dagger}})^{n}$ $=\sum_{k=0}^{n}\left\{{{n+1}\atop{k+1}}\right\}{a^{\dagger}}^{k}a^{k}=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(0,1,1)\,{a^{\dagger}}^{k}a^{k}.$ (48)
These two are due to Katriel [44], who showed for the first time how Stirling numbers arise in boson algebra. Each can also be proved combinatorially, by applying Wick’s theorem and enumerating contractions [45, §9.3].
Clearly (46b),(46b) follow from (47),(48) by multiplying by $\lambda^{n}/n!$ and summing over $n$. What is needed for a generalization to the exponential operator ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ is a formula for $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}$. Duchamp et al. proved the following.
Proposition 4.1 ([46]).
If $\omega$ is any “single annihilator” word over the alphabet $\{a,{a^{\dagger}}\}$, i.e., $\omega={a^{\dagger}}^{L}a{a^{\dagger}}^{R}$, and $e=L+R-1\geqslant 0$, then for all $n\geqslant 0$,
$\omega^{n}=\left({a^{\dagger}}^{e}\right)^{n}\left[\sum_{k=0}^{n}\mathbf{R}[d,h]_{n,k}\,{a^{\dagger}}^{k}a^{k}\right],$ (49)
an identity in $\mathcal{W}$, where $\mathbf{R}[d,h]$ is some $\omega$-specific exponential Riordan array.
What is difficult is finding explicitly the $(L,R)$-dependent series $d,h\in\mathbb{C}[[z]]$ that define the Riordan array. Identities (46a),(46b) suggest that $\mathbf{R}[d,h]$ should be of Hsu–Shiue type. But the derivation of the identity (49) need not be combinatorial, or based on a generalization of Wick’s theorem.
Combinatorial interpretations of the reordered coefficients of such boson strings as ${a^{\dagger}}^{r_{n}}a^{s_{n}}\dots a^{\dagger}{}^{r_{1}}a^{s_{1}}\in\mathcal{W}$ are known [18]. Also, interpretations of the coefficients $\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)$ for general integer values of $A,B,r$ are known, such as ones involving the placing of balls in cells and their compartments, or the drawing of balls from an urn. But such combinatorial interpretations tend at present to be complicated. (See, e.g., [35, Chapter 4].) The following two parametric identities in $\mathcal{W}$ are extracted from [41, Theorem 6.5], and have a non-combinatorial proof. The first subsumes Katriel’s formulae, (47) and (48).
Theorem 4.2.
For all integer $L,R\geqslant 0$ with $e=L+R-1\geqslant 0$,
$\begin{aligned} \left({a^{\dagger}}^{L}a{a^{\dagger}}^{R}\right)^{n}&=\left({a^{\dagger}}^{e}\right)^{n}\left[\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(-e,1,R)\,{a^{\dagger}}^{k}a^{k}\right]\\ &=\mathopen{:}\,\mathrel{\left({a^{\dagger}}^{e}\right)^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-e,1,R)}\,\mathclose{:}_{N}\end{aligned}$ (50)
and
$\begin{aligned} \left({a^{\dagger}}^{L}a{a^{\dagger}}^{R}\right)^{n}&=\left[\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(e,-1,-L)\,a^{k}{a^{\dagger}}^{k}\right]\left({a^{\dagger}}^{e}\right)^{n}\\ &=\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}(a{a^{\dagger}};\,e,-1,-L)\left({a^{\dagger}}^{e}\right)^{n}}\,\mathclose{:}_{A},\end{aligned}$ (51)
where the $\mathop{{}\it HS}\nolimits_{n,k}$ coefficients can be computed from the summation formula (40) or the recurrence (42).
Proof.
Representing $a,{a^{\dagger}}$ by $D_{x},x$, acting on any convenient space of functions of $x$, yields a faithful representation of $\mathcal{W}$, so it suffices to obtain versions of (50),(51) with $a,{a^{\dagger}}$ replaced by $D_{x},x$.
Replacing $x$ by $xD_{x}$ in the definition (34) of the coefficients $\mathop{{}\it HS}\nolimits_{n,k}$ yields
$(xD_{x}+r)^{\underline{n},A}=\sum_{k=0}^{n}HS_{n,k}(A,B,r)\,(xD_{x})^{\underline{k},B}.$ (52)
This can be rewritten with the aid of Boole’s lemma $(xD_{x})^{\underline{n}}=x^{n}D_{x}^{n}$, which has a long history [35]. The more general forms of the lemma
$(xD_{x})^{\underline{n},A}$ $=(x^{A})^{n}(x^{1-A}D_{x})^{n},$ (53a)
$=(xD_{x}x^{A})^{n}(x^{-\alpha})^{n}$ (53b)
(cf. [47]) can be proved by induction, or operationally: the left and right-hand sides act identically on monomials in $x$. The yet more general forms
$(xD_{x}+r)^{\underline{n},A}$ $=x^{-r}\left[(x^{A})^{n}(x^{1-A}D_{x})^{n}\right]x^{r}=(x^{A})^{n}(x^{1-A-r}D_{x}x^{r})^{n},$ (54a)
$=x^{-r}\left[(xD_{x}x^{A})^{n}(x^{-A})^{n}\right]x^{r}=(x^{1-r}D_{x}x^{A+r})^{n}(x^{-A})^{n}$ (54b)
follow by the similarity transformation $(\cdot)\mapsto x^{-r}(\cdot)x^{r}$.
To obtain a version of (50) with $a,{a^{\dagger}}$ replaced by $D_{x},x$, rewrite the left-hand side of (52) with the aid of (54a), and the right-hand side with the aid of (53a); and also set (A,B,r)=(e,1,R)(A,B,r)=(-e,1,R). To obtain (51) similarly, use (54b) and (53b), and set $(A,B,r)=\allowbreak(e,-1,1-\nobreak L)$. ∎
Example 4.3.
If $L=R=1$, identities (50),(51) of the theorem reduce to
$\left({a^{\dagger}}a{a^{\dagger}}\right)^{n}=\left\{\begin{aligned} &\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-1,1,1)}\,\mathclose{:}_{N}=n!\;\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}L_{n}(-{a^{\dagger}}a)}\,\mathclose{:}_{N},\\ &\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}(a{a^{\dagger}};\,1,-1,-1)a^{n}}\,\mathclose{:}_{A}=(-1)^{n}n!\;\mathopen{:}\,\mathrel{L_{n}(a{a^{\dagger}}){a^{\dagger}}^{n}}\,\mathclose{:}_{A},\end{aligned}\right$. (55)
by (45), so that
$\left({a^{\dagger}}a{a^{\dagger}}\right)^{n}=\left\{\begin{aligned} &{a^{\dagger}}^{n}\cdot\sum_{k=0}^{n}\frac{n!^{2}}{k!^{2}(n-k)!}\,{a^{\dagger}}^{k}a^{k},\\ &\sum_{k=0}^{n}(-1)^{n-k}\frac{n!^{2}}{k!^{2}(n-k)!}\,a^{k}{a^{\dagger}}^{k}\cdot{a^{\dagger}}^{n}.\end{aligned}\right$. (56)
These are known Laguerre-polynomial representations [48].
Multiplying the identities in Theorem 4.2 by $\lambda^{n}/n!$ and summing over $n$ yields the following identities in $\mathcal{W}[[\lambda]]$, with the two explicit expressions coming from the EGF formula in Proposition 3.4. The first subsumes (46b) and (46b).
Theorem 4.4.
For all integer $L,R\geqslant 0$ with $e=L+R-1\geqslant 0$, $G={\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ has the ordered representations
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,-e,1,R)}\,\mathclose{:}_{N}\\ &=\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl{(}1-e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}\end{aligned}$ (57)
and
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}(a{a^{\dagger}},\lambda{a^{\dagger}}^{e};\,e,-1,-L)}\,\mathclose{:}_{A}\\ &=\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl{(}1+e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A},\end{aligned}$ (58)
with the case $e=0$ handled by taking the formal $e\to 0$ limit.
Remark 4.4.1.
[…]
Remark 4.4.2.
The two identities $G=\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{N}$ and $G=\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{A}$, with their explict right-hand sides coming from Proposition 3.4, imply each other. To convert either to the other, (1) take the adjoint, (2) replace $a,{a^{\dagger}}$ by ${a^{\dagger}},-a$ (an automorphism of $\mathcal{W}$, entailing the interchange of $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{N}$ and $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{A}$), and (3) negate $\lambda$ and interchange $L,R$.
The $R=0$ case of (57) was previously derived by Lang [14, 15] and Błasiak et al. [16]. They used a representation of $\mathcal{W}$ in which $a,{a^{\dagger}}$ are represented by $D_{x},x$, and exponentiated ${a^{\dagger}}^{L}a$ by integrating the vector field $x^{L}D_{x}$. Also, Penson et al. [48] obtained what amounted to the $R=1$ case of (57) by exploiting a Dobiński representation of generalized Bell polynomials.
More generally, Błasiak et al. [16] integrated $q(x)D_{x}+v(x)$, where $q,v\in\mathbb{C}[[x]]$. As ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ equals $({a^{\dagger}})^{L+R}a+\nobreak R({a^{\dagger}})^{L+R-1}$, by substituting $q(x)=x^{L+R}$ and $v(x)=Rx^{L+R-1}$ into their formulae one can derive (57) in full generality.
Błasiak et al. [17] also used Sheffer polynomials to normal order an exponentiated version of the operator $\tilde{M}$ mentioned above at the end of §3.2. Up to the taking of an adjoint, they derived a formula for ${\rm e}^{\lambda\tilde{M}^{\dagger}}$: they showed that
$\begin{split}&\exp\left\{\lambda\,\frac{1}{f^{\prime}({a^{\dagger}})}\left[a-\frac{g^{\prime}({a^{\dagger}})}{g({a^{\dagger}})}\right]\right\}\\ &\quad=\mathopen{:}\,\mathrel{\frac{g({a^{\dagger}})}{g\left(\bar{f}(f({a^{\dagger}})+\lambda)\right)}\exp{\Bigl{\{}\left[\bar{f}(f({a^{\dagger}})+\lambda)-{a^{\dagger}}\right]a\Bigr{\}}}}\,\mathclose{:}_{N},\end{split}$ (59)
where $g,f\in\mathbb{C}[[D]]$ are any pair that defines a Sheffer sequence. Equation (59) was not obtained in the power-series framework of the present paper, as its derivation employed a Fock space representation and projections onto coherent states, but it is a valid identity in $\mathbb{C}[[{a^{\dagger}}]][a][[\lambda]]$.
By examination, formally substituting $g(D)=D^{-R}$, $f(D)=-D^{-e}/e$ into (59) yields (57). But this additional derivation of eq. (57) is not rigorous, as these $g,f$ contain negative powers of $D$ and do not belong to $\mathbb{C}[[D]]$. At most, it is an analytic continuation of a rigorous derivation.
5 A new parametric family and the main result
The main result (Theorem 5.3) can now be approached. According to Theorem 4.4, the normally ordered ($s=\nobreak-1$) and anti-normally ordered ($s=\nobreak+1$) representations of the exponential operator ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ can be expressed in terms of $\mathcal{HS}(t,z;\allowbreak A,B,r)$, the EGF of the parametric family of exponential Riordan arrays of Hsu–Shiue type, for which there is an explicit expression. (Recall Proposition 3.4.) For any $s\in\mathbb{C}$, the $s$-ordered representation of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ will now be obtained in terms of a new family of exponential Riordan arrays, or family of Sheffer polynomial sequences, which is of independent interest.
5.1 The new parametric family
Any member of the Hsu–Shiue family can be viewed as the $n$-indexed sequence of row polynomials coming from an exponential Riordan array $\mathbf{R}[d_{\textit{HS}}(z),\allowbreak h_{\textit{HS}}(z)]$, for known elementary functions $d_{\textit{HS}}(z)$, $h_{\textit{HS}}(z)$, and its EGF $\mathcal{HS}(t,z)$ equals $d_{\textit{HS}}(z){\rm e}^{th_{\textit{HS}}(z)}$. Equivalently, the member is viewed in Sheffer terms as $\mathbf{S}[g_{\textit{HS}}(D),\allowbreak f_{\textit{HS}}(D)]$, where $g_{\textit{HS}},f_{\textit{HS}}]$ is the inverse in the exponential Riordan group $\mathfrak{R}$ of the element $d_{\textit{HS}},h_{\textit{HS}}]$. The function $D=h_{\textit{HS}}(z)$ has compositional inverse $z=f_{\textit{HS}}(D)$, and $g_{\textit{HS}}(D)$ equals $1/d_{\textit{HS}}(z)=1/d_{\textit{HS}}(f_{\textit{HS}}(D))$.
As noted in Sec. 3.3, the Sheffer functions $g_{\textit{HS}},f_{\textit{HS}}$ turn out to be minor variations of the Riordan functions $d_{\textit{HS}},h_{\textit{HS}}$: their parameters $A,B$ are interchanged and $r$ is negated. This is a special property of the Hsu–Shiue family.
When $s\neq\pm 1$ and $L+R>1$, the exponential Riordan EGF in terms of which the $s$-ordered representation of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ is expressed in Theorem 5.3 will come from a Riordan pair $[d,h]$ in which $d(z),\allowbreak h(z)$ are not elementary functions of $z$. But the functions in the inverse pair $g,f]=[d,h]^{-1}$ will be elementary, and one can view $[d,h]$ as $\bar{g},\bar{f}]\mathrel{:=}[g,f]^{-1}$, if necessary, and write the EGF as $\bar{g}(z){\rm e}^{t\bar{f}(z)}$. So the sequence of polynomials from which the EGF is computed can optionally be denoted more briefly and comprehensibly as $\mathbf{S}[g,f]$ (in Sheffer terms), than as $\mathbf{R}[d,h]$ (in exponential Riordan terms).
The following defines a new family of Sheffer polynomial sequences with this feature, which will appear in Theorem 5.3. It extends the Hsu–Shiue family, having the extra parameter $s$ and the pair $r,r^{\prime}$ instead of $r$. Why it is called the “two-point” Hsu–Shiue family will be clear from Proposition 5.2.
Firstly, define parametric functions $g_{T},f_{T}$ of $D$, or more accurately parametric Maclaurin series $g_{T},f_{T}\in\mathbb{C}[[D]]$, by
$g_{T}(D;B,r,r^{\prime};\,s)\mathrel{:=}\left(\frac{1+s}{2}\right)\left[\frac{1+(\frac{1-s}{2})B\,D}{1-(\frac{1+s}{2})B\,D}\right]^{-r/B}\!\!+\left(\frac{1-s}{2}\right)\left[\frac{1-(\frac{1+s}{2})B\,D}{1+(\frac{1-s}{2})B\,D}\right]^{r^{\prime}/B},$ (60)
$z=f_{T}(D;A,B;\,s)\mathrel{:=}\frac{[1-(\frac{1+s}{2})B\,D]^{-A/B}-1}{A}+\frac{[1+(\frac{1-s}{2})B\,D]^{-A/B}-1}{-A},$
with the cases $A=0$, $B=0$ handled by taking limits. These $g_{T},f_{T}$ have Maclaurin expansions that are series in $\mathbb{C}[[D]]$ of order 0 and $1$ respectively: $g_{T}=1\cdot D^{0}+\nobreak O(D^{1})$ and $f_{T}=1\cdot D^{1}+\nobreak O(D^{2})$.
Definition 5.1.
The parametric sequence $T_{n}(t)=T_{n}(t;\,A,B,r,r^{\prime};\,s)$ of two-point Hsu–Shiue polynomials, denoted by $\mathcal{T}(A,B,r,r^{\prime};\,s)$, is the Sheffer sequence $\mathbf{S}[g_{T}(\cdot;B,r,r^{\prime};s),\allowbreak f_{T}(\cdot;A,B;s)]$, with the EGF
$\mathcal{T}(t,z;\,A,B,r,r^{\prime};\,s)=\sum_{n=0}^{\infty}T_{n}(t;\,A,B,r,r^{\prime};\,s)\frac{z^{n}}{n!}=\bar{g}_{T}(z;\,s){\rm e}^{t\bar{f}_{T}(z;\,s)}.$
Here, the pair $\bar{g}_{T},\bar{f}_{T}]$ is the inverse in the exponential Riordan group $\mathfrak{R}$ of the parametric element $g_{T},f_{T}]$ defined in eq. (60). That is, $D=\bar{f}_{T}(z;s)$ is inverse to $z=f_{T}(D;s)$ and can be obtained by series reversion, and $\bar{g}_{T}=1/g_{T}(D;s)=1/g_{T}(\bar{f}_{T}(z;s))$. The notations $\mathbf{S}[g_{T},f_{T}]$ and $\mathbf{R}[\bar{g}_{T},\bar{f}_{T}]$ are equivalent.
It should be clear from the definition of $f_{T}$ in (60) that for general $s$ and $A,B$, the inverse $D=\bar{f}_{T}(z)$ of $z=f_{T}(D)$ is not an elementary function. But:
Proposition 5.2.
The sequence $\mathcal{T}(A,B,r,r^{\prime};s)$ of two-point Hsu–Shiue polynomials reduces as follows when $s=\pm 1$:
$\begin{aligned} &{T}_{n}(t;A,B,r,r^{\prime};\,-1)\\ &\quad=\mathbf{S}[d_{\textit{HS}}(\cdot;B,-r^{\prime}),h_{\textit{HS}}(\cdot;B,-A)]_{n}(t)=\mathbf{R}[d_{\textit{HS}}(\cdot;-A,r^{\prime}),h_{\textit{HS}}(\cdot;-A,B)]_{n}(t),\end{aligned}$
and
$\begin{aligned} &{T}_{n}(t;A,B,r,r^{\prime};\,+1)\\ &\quad=\mathbf{S}[d_{\textit{HS}}(\cdot;-B,-r),h_{\textit{HS}}(\cdot;-B,A)]_{n}(t)=\mathbf{R}[d_{\textit{HS}}(\cdot;A,r),h_{\textit{HS}}(\cdot;A,-B)]_{n}(t).\end{aligned}$
That is, individual polynomials reduce according to
${T}_{n}(t;\,A,B,r,r^{\prime};\,-1)={HS}_{n}(t;\,-A,B,r^{\prime}),$
${T}_{n}(t;\,A,B,r,r^{\prime};\,+1)={HS}_{n}(t;\,A,-B,r),$
and their EGF’s reduce according to
$\mathcal{T}(t,z;\,A,B,r,r^{\prime};\,-1)$ $=\mathcal{HS}(t,z;\,-A,B,r^{\prime})$
$=(1-Az)^{-r^{\prime}/A}\exp\left\{\frac{t}{B}\left[(1-Az)^{-B/A}-1\right]\right\}$
and
$\mathcal{T}(t,z;\,A,B,r,r^{\prime};\,+1)$ $=\mathcal{HS}(t,z;\,A,-B,r)$
$=(1+Az)^{r/A}\exp\left\{\frac{t}{B}\left[1-(1+Az)^{-B/A}\right]\right\}.$
Proof.
Set s=1,+1s=-1,+1 in the formulae (60) for $g_{T},f_{T}$, and compare them with the formulae (36) for $d_{\textit{HS}},h_{\textit{HS}}$. The formula for the parametric Hsu–Shiue EGF $\mathcal{HS}(t,z)$ was supplied in Proposition 3.4. ∎
So, exponential Riordan arrays of the newly defined two-point Hsu–Shiue type (and likewise their row polynomials and EGF’s) interpolate as $s$ varies over the $1$-simplex $1\leqslant s\leqslant+1$} between conventional Hsu–Shiue ones, to which they reduce when $s=\pm 1$. One could also define three-point interpolation, with a parameter vector that varies over a $2$-simplex, etc.
5.2 Main result
The following theorem states how the $s$-ordered representation of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ can be expressed in terms of the two-point Hsu–Shiue EGF $\mathcal{T}(t,z;A,B,r,r^{\prime};s)$, and subsumes Theorem 4.4, which provided only normal ($s=\nobreak-1$) and anti-normal ($s=\nobreak+1$) orderings. Examples are given in Sec. 6. The corollary to this theorem (Corollary 5.4) expresses $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}$ in terms of the row polynomials $T_{n}(t;\allowbreak A,B,r,r^{\prime};s)$, and subsumes Theorem 4.2.
Theorem 5.3.
For all integer $L,R\geqslant 0$ with $e=L+R-1\geqslant 0$, $G={\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ has for all $s\in\mathbb{C}$ the $s$-ordered representation
${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ $=\mathopen{:}\,\mathrel{\mathcal{T}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,e,1,-L,R;\,s)}\,\mathclose{:}_{s}$ (61)
$=\mathopen{:}\,\mathrel{\bar{g}_{T}(\lambda{a^{\dagger}}^{e};\,s)\exp\left[({a^{\dagger}}a)\bar{f}_{T}(\lambda{a^{\dagger}}^{e};\,s)\right]}\,\mathclose{:}_{s},$ (62)
where $\bar{g}_{T},\bar{f}_{T}]$ is the Riordan-group inverse of $g_{T},f_{T}]$, defined as in eq. (60), with parameters $(A,B,r,r^{\prime};s)=(e,1,-L,R;s)$. This reduces when $s=\pm 1$ to
$\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl{(}1-e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}$ $\textrm{if }s=-1$
and
$\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl{(}1+e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A}$ $\textrm{if }s=+1,$
where in both, the $e=0$ case is handled by taking the formal $e\to 0$ limit.
Proof.
The idea is this: $G={\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ equals $\mathopen{:}\,\mathrel{F^{(s)}(\alpha,\alpha^{*})}\,\mathclose{:}_{s}$, where $F^{(s)}(\alpha,\alpha^{*})$ is some function that satisfies the diffusion equation (23), a PDE that evolves $F^{(s)}$ toward smaller values of $s$. (In this proof, for clarity the commuting arguments of $F^{(s)}$ will be written as $\alpha,\alpha^{*}$ and not as $a,{a^{\dagger}}$. The notational convention in quantum optics that dependence on $\alpha,\alpha^{*}$ really means dependence on a single complex variable $\alpha=\allowbreak\alpha_{r}+\nobreak{\rm i}\alpha_{i}$ will briefly be followed.)
In principle, one could start with $F^{(+1)}=F^{(A)}$ and integrate the diffusion equation backward in $s$ toward $F^{(-1)}=F^{(N)}$, as one does in quantum optics when smoothing the Glauber–Sudarshan $P$-function ($s=\nobreak+1$) to obtain the Husimi $Q$-function ($s=\nobreak-1$) [2, 3]. The differential operator $\partial^{2}/\partial\alpha\partial\alpha^{*}$ in the PDE is to be interpreted as $\Delta/4$, with $\Delta=\partial^{2}/\partial\alpha_{r}^{2}+\partial^{2}/\partial\alpha_{i}^{2}$ the Laplacian on the complex $\alpha$-plane.
However, to avoid convergence questions the exponential operator $G$ is viewed in this paper as an element of $\mathcal{W}[[\lambda]]$, and $F^{(s)}$ is accordingly an element of $\mathcal{W}_{c}[[\lambda]]$. So $F^{(s)}$ is not a single function on the complex $\alpha$-plane (i.e., a function of $\alpha_{r},\alpha_{i}$ or informally of $\alpha,\alpha^{*}$), and the convolution or smoothing formula (24) cannot easily be used. Instead, $F^{(s)}$ at each value of $s$ is a power series in $\lambda$, with coefficients that are polynomial in $\alpha,\alpha^{*}$, and the series is formal: it is never summed. The following proof should be read with this in mind.
The statement of the theorem incorporates the $s=\pm 1$ expressions for $F^{(s)}$ from Theorem 4.4. To match their functional form, make the Ansatz
$F^{(s)}(\alpha,\alpha^{*})={V}(\alpha^{*}\alpha,\lambda{\alpha^{*}}^{e};\,s),$ (63)
where $\lambda\neq 0$ and the function $V(t,z;s)$ is to be determined. Changing the independent variables $\alpha,\alpha^{*}$ (really $\alpha_{r},\alpha_{i}$) to $t=\alpha^{*}\alpha$, $z=\lambda{\alpha^{*}}^{e}$ converts the diffusion equation (23) to an equation of diffusion type on the $(t,z)$-plane,
$\left[\frac{\partial}{\partial s}+\mathcal{X}\right]V(t,z;\,s)=0,\qquad\mathcal{X}\mathrel{:=}\frac{1}{2}\left(t\frac{\partial^{2}}{\partial t^{2}}+\frac{\partial}{\partial t}+ez\,\frac{\partial^{2}}{\partial t\,\partial z}\right),$ (64)
irrespective of $\lambda$; so that formally,
$V(\cdot,\cdot;\,s)=\exp\left[(s_{0}-s)\mathcal{X}\right]\,V(\cdot,\cdot;\,s_{0}),\qquad s<s_{0}.$ (65)
There is an obvious $z$-independent solution, namely
$V(t,z;\,s)\varpropto(s_{0}-s)^{-1}{\rm e}^{-2t/(s_{0}-s)},\qquad s<s_{0},$ (66)
for any real $s_{0}$. Depending on $t=\alpha^{*}\alpha$ but not on $z=\lambda{\alpha^{*}}^{e}$, for $s_{0}=+1$ this is in fact the $s$-ordered quasiprobability density on the complex $\alpha$-plane coming from the density operator of the vacuum state [2, §12.4.4], showing how the $P$-function ($s=\nobreak+1$), which is a $\delta$-function located at $\alpha=0$, broadens into a Gaussian as $s$ is decreased and the $Q$-function ($s=\nobreak-1$) is approached. But it is qualitatively different from the expressions $V(t=\alpha^{*}\alpha,\allowbreak z=\lambda{\alpha^{*}}^{e};\pm 1)$ appearing in Theorem 4.4.
To incorporate those functions, make the further Riordan/Sheffer Ansatz
$V(t,z;\,s)=\mathbf{R}[\bar{g}(\cdot;s),\bar{f}(\cdot;s)](t,z)=\mathbf{S}[g(\cdot;s),f(\cdot;s)](t,z),$ (67)
which says that
$V(t,z;\,s)=\bar{g}(z;\,s)\,{\rm e}^{t\bar{f}(z;\,s)},$ (68)
where as usual, $\bar{g}(z;s)=1\cdot z^{0}+O(z^{1})$ and $D=\bar{f}(z;s)=1\cdot z^{1}+O(z^{2})$, with $g(D;s)=1\cdot D^{0}+O(D^{1})$ and $z=f(D;s)=1\cdot D^{1}+O(D^{2})$. Substituting (68) into (64) and separating the terms proportional to $t^{0}$ and $t^{1}$ gives
$2\bar{g}_{s}+ez\bar{f}\bar{g}_{z}+(\bar{f}+ez\bar{f}_{z})g=0,$ (69a)
$2\bar{f}_{s}+ez\bar{f}\bar{f}_{z}+\bar{f}^{2}=0,$ (69b)
where subscripts denote partial derivatives. These are coupled nonlinear PDEs for $\bar{g}(z;s)$, $\bar{f}(z;s)$. The two-point boundary conditions
$\bar{g}(z;\,s=-1)=(1-ez)^{-R/e},$ $\bar{g}(z;\,s=+1)=(1+ez)^{-L/e},$ (70a)
$\bar{f}(z;\,s=-1)=(1-ez)^{-1/e}-1,$ $\bar{f}(z;\,s=+1)=1-(1+ez)^{-1/e}$ (70b)
come from Theorem 4.4.
The function (or more accurately formal infinite series) $z=f(D;s)$ is the inverse of $D=\bar{f}(z;s)$, computed by series reversion. With its aid, the independent variables $z,s$ in (69) can be replaced by $D,s$, and the dependent variables $\bar{g},\bar{f}$ by $g,f$, where $g=1/(\bar{g}\circ f)$, i.e., $\bar{g}=1/(g\circ\bar{f})$. This yields the PDEs
$-2f_{D}\,g_{s}+D^{2}f_{D}\,g_{D}+(Df_{D}+ef)g=0,$ (71a)
$-2f_{s}+D^{2}f_{D}+eDf=0,$ (71b)
which are satisfied by $f(D;s)$ and $g(D;s)$. The two-point boundary conditions on these dual functions follow from (70a),(70b) and are
$g(D;\,s=-1)=(1+D)^{-R},$ $g(D;\,s=+1)=(1-D)^{-L},$ (72a)
$f(D;\,s=-1)=\left[1-(1+D)^{-e}\right]/e,$ $f(D;\,s=+1)=\left[(1-D)^{-e}-1\right]/e.$ (72b)
The “dualized” PDEs (71) are linear and of first order, and can be solved by Lagrange’s method of characteristics.
The solution of (71b) is
$z=f(D;\,s)=D^{-e}F(\eta),$ (73)
a similarity solution in which $\eta\mathrel{:=}s-2/D$ is the similarity variable and the function $F(\eta)$ is arbitrary. To match the boundary conditions (72b), choose
$F(\eta)=\frac{\left(\frac{-1-\eta}{2}\right)^{-e}-\left(\frac{1-\eta}{2}\right)^{-e}}{e},$ (74)
so that
$z=f(D;\,s)=\frac{\left[1-(\frac{1+s}{2})D\right]^{-e}-{\left[1+(\frac{1-s}{2})D\right]^{-e}}}{e},$ (75)
with expansion $f(D;s)=1\cdot D^{1}+O(D^{2})$. With this choice of f(D;s)f(D;\,s), the PDE (71a) is solved in the same way. It has the similarity solution
$g(D;\,s)=\left[\left(\frac{1+s}{2}\right)+\left(\frac{1-s}{2}\right)\left(\frac{\eta+1}{\eta-1}\right)^{e+1}\right]G(\eta),$ (76)
where the function $G(\eta)$ is arbitrary. To match the boundary conditions (72a), choose
$G(\eta)=\left(\frac{\eta-1}{\eta+1}\right)^{L},$ (77)
so that
$g(D;\,s)=\left(\frac{1+s}{2}\right)\left[\frac{1+(\frac{1-s}{2})D}{1-(\frac{1+s}{2})D}\right]^{L}+\left(\frac{1-s}{2}\right)\left[\frac{1-(\frac{1+s}{2})D}{1+(\frac{1-s}{2})D}\right]^{R},$ (78)
with $g(D;s)=1\cdot D^{0}+O(D^{1})$. Comparing (75),(78) with the definitions of $f_{T},g_{T}$ given in (60) reveals that
$f(D;\,s)=f_{T}(D;e,1;\,s),\qquad g(D;\,s)=g_{T}(D;1,-L,R;\,s).$ (79)
Therefore
$V(t,z;\,s)=\mathcal{T}(t,z;\,e,1,-L,R;\,s),$ (80)
which is the substance of the theorem. ∎
Remark.
Wilcox [49] solves a PDE similar to (64) by normal ordering an exponentiated second-order differential operator analogous to the $\exp\left[(s_{0}-\nobreak s)\mathcal{X}\right]$ in eq. (65). In his paper the PDE comes first, but in the just-concluded proof of Theorem 5.3 the normal ordering (actually $s$-ordering) came first, and determined the PDE to be solved.
Corollary 5.4.
For all integer $L,R\geqslant 0$ with $e=L+R-1\geqslant 0$,
$({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}=\mathopen{:}\,\mathrel{T_{n}({a^{\dagger}}a;e,1,-L,R;s)}\,\mathclose{:}_{s},$
where $T_{n}(t;e,1,-L,R;s)$ is the case $(A,B,r,r^{\prime})=(e,1,-L,R)$ of the nn\!’th parametric Sheffer polynomial $\mathbf{S}[g_{T},f_{T}]_{n}(t)=\mathbf{R}[\bar{g}_{T},\bar{f}_{T}]_{n}(t)$, with $g_{T}(D;s)$ and $z=f_{T}(D;s)$ given in Definition 5.1, i.e., in eq. (60).
Proof.
Expand each side of the identity (61) in Theorem 5.3 as a formal power series in $\lambda$, and equate the coefficients of $\lambda^{n}$ on the two sides. ∎
The following is the Weyl ordering ($s=\nobreak 0$) case of the main result Theorem 5.3, which is the one that can be expressed most simply. The functions $\bar{g}(z)$, $D=\bar{f}(z)$ come from $g(D)$, $z=f(D)$, which in turn come from the two-point functions $g_{T}(D;s)$, $z=f_{T}(D;s)$ defined in (60), by setting $s=\nobreak 0$.
Theorem 5.5.
For all integer $L,R\geqslant 0$ with $e=L+R-1\geqslant 0$, one has the Weyl ordering
${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}=\mathopen{:}\,\mathrel{\bar{g}(\lambda{a^{\dagger}}^{e})\,\exp\left[({a^{\dagger}}a)\,\bar{f}(\lambda{a^{\dagger}}^{e})\right]}\,\mathclose{:}_{W},$
an identity in $\mathcal{W}[[\lambda]]$ in which $D=\bar{f}(z)$ is the compositional inverse of
$z=f(D)=\frac{(1-\frac{1}{2}D)^{-e}-1}{e}+\frac{(1+\frac{1}{2}D)^{-e}-1}{-e}$
with expansion $D=\bar{f}(z)=1\cdot z^{1}+O(z^{2})$, and $\bar{g}(z)=1/g(D)=1/g(\bar{f}(z))$ where
$g(D)=\frac{1}{2}\left[\frac{1+\frac{1}{2}D}{1-\frac{1}{2}D}\right]^{L}\!+\frac{1}{2}\left[\frac{1-\frac{1}{2}D}{1+\frac{1}{2}D}\right]^{R},$
with expansion $\bar{g}(z)=1\cdot z^{0}+O(z^{1})$ The case $e=0$ is handled by taking the formal $e\to 0$ limit.
6 Examples
The main result (Theorem 5.3) will now be applied, to produce several new $s$-ordered representations of exponential boson operators ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}\in\mathcal{W}[[\lambda]]$, and of $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}\in\nobreak\mathcal{W}$.
According to the theorem, for any $(L,R)$ one needs the EGF of a certain family of two-point Hsu–Shiue polynomials (of Sheffer type), denoted by $\mathcal{T}(t,z;\allowbreak A,B,r,r^{\prime};s)$. The EGF equals $\bar{g}_{T}(z){\rm e}^{t\bar{f}_{T}(z)}$, where $\bar{g}_{T},\bar{f}_{T}]$ is the inverse in the exponential Riordan group $\mathfrak{R}$ of the parametric $s$-dependent element $g_{T},f_{T}]$ defined in eq. (60). One simply substitutes $(A,B,r,r^{\prime})=(e,1,-L,R)$, where $e\mathrel{:=}\allowbreak L+\nobreak R-\nobreak 1$, and sets $(t,z)=({a^{\dagger}}a,\lambda{a^{\dagger}}^{e})$ (which really means $(t,z)=(\alpha^{*}\alpha,\lambda{\alpha^{*}}^{e})$), to obtain the desired $s$-ordered representation.
The only possible stumbling block is the computation of the inverse group element (the function $D=\bar{f}_{T}(z;s)$ is the compositional inverse of $z=f_{T}(D;s)$). This potential difficulty has already appeared in the statement of Theorem 5.5, without comment. The following subsections treat the cases when $e=0,1,2$.
6.1 The case $e=0$
This includes the operators ${\rm e}^{\lambda({a^{\dagger}}a)}$, ${\rm e}^{\lambda(a{a^{\dagger}})}$, with $(L,R)=(1,0)$, $(0,1)$. Substituting $(A,B,r,r^{\prime})=\allowbreak(e,1,-L,R)=\allowbreak(0,1,-1,0)$ into (60) gives
$g_{T}(D;\,s)=\frac{2}{2-D-sD},\qquad z=f_{T}(D;\,s)=\ln\left(\frac{2+D-sD}{2-D-sD}\right).$ (81)
(In deriving the latter, substituting $A=0$ means taking the $A\to 0$ limit.) The inverse $D=\bar{f}_{T}(z;s)$ and $\bar{g}_{T}(z;s)=1/g_{T}(D=\bar{f}_{T}(z;s);s)$ become
$\bar{g}_{T}(z;\,s)=\frac{2}{1+{\rm e}^{z}-s(1-{\rm e}^{z})},\qquad D=\bar{f}_{T}(z;\,s)=\frac{2({\rm e}^{z}-1)}{1+{\rm e}^{z}-s(1-{\rm e}^{z})}.$ (82)
Substituting into $\bar{g}_{T}(z;s){\rm e}^{t\bar{f}_{T}(z;s)}$ and setting $(t,z)=({a^{\dagger}}a,\lambda{a^{\dagger}}^{e})=\allowbreak({a^{\dagger}}a,\lambda)$ yields
${\rm e}^{\lambda({a^{\dagger}}a)}=\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}{a^{\dagger}}a\right]}\,\mathclose{:}_{s},$ (83)
which is precisely the original Cahill–Glauber $s$-ordering result, eq. (2). The case $(L,R)=(0,1)$ is similar.
6.2 The case $e=1$
This includes the operators ${\rm e}^{\lambda({a^{\dagger}}^{2}a)}$, ${\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}})}$, ${\rm e}^{\lambda(a{a^{\dagger}}^{2})}$, with $(L,R)=(2,0)$, $(1,1)$, $(0,2)$. Substituting $(A,B)=(e,1)=(1,1)$ into (60) gives
$g_{T}(D;\,s)=\left\{\begin{aligned} &\frac{4+D^{2}-s^{2}D^{2}}{(2-D-sD)^{2}},&\qquad&(L,R)=(2,0);\\ &\frac{4+D^{2}-s^{2}D^{2}}{(2-sD)^{2}-D^{2}},&\qquad&(L,R)=(1,1);\\ &\frac{4+D^{2}-s^{2}D^{2}}{(2+D-sD)^{2}},&\qquad&(L,R)=(0,2);\end{aligned}\right$. (84)
and
$z=f_{T}(D;\,s)=\frac{4D}{4-D^{2}}.$ (85)
Hence,
$\bar{g}_{T}(z;\,s)=\left\{\begin{aligned} &\frac{1+s}{1+2sz+z^{2}+(z+s)\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(2,0);\\ &\frac{1}{\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(1,1);\\ &\frac{1-s}{1+2sz+z^{2}-(z+s)\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(0,2);\end{aligned}\right$. (86)
and
$D=\bar{f}_{T}(z;\,s)=\frac{2z}{1+sz+\sqrt{1+2sz+z^{2}}}.$ (87)
Substituting into $\bar{g}_{T}(z;s){\rm e}^{t\bar{f}_{T}(z;s)}$, with $(t,z)=({a^{\dagger}}a,\lambda{a^{\dagger}}^{e})=({a^{\dagger}}a,\lambda{a^{\dagger}})$, yields
Theorem 6.1.
One has the $s$-ordered representations
$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{2}a)}&=\mathopen{:}\,\frac{1+s}{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}+(\lambda{a^{\dagger}}+s)\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\\ {\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}})}&=\mathopen{:}\,\frac{1}{\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\\ {\rm e}^{\lambda(a{a^{\dagger}}^{2})}&=\mathopen{:}\,\frac{1-s}{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}-(\lambda{a^{\dagger}}+s)\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\end{aligned}$ (88)
where
$\begin{aligned} \mathcal{E}\mathrel{:=}\exp\left[({a^{\dagger}}a)\,\frac{2\lambda{a^{\dagger}}}{1+s\lambda{a^{\dagger}}+\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\right],\end{aligned}$ (89)
for all $s\in\mathbb{C}$.
When $s=\nobreak-1$ and $s=\nobreak+1$, each of these three ordered representations reduces to a normal ordered and an anti-normal ordered one, which are the ones provided by Theorem 4.4. (To set $s=\nobreak-1$ in the first or $s=\nobreak+1$ in the third, a limit must be taken.) But, the three Weyl-ordered representations obtained by setting $s=\nobreak 0$ seem to be new to the literature.
From such identities as those of Theorem 6.1, it is possible to derive explicitly $s$-ordered series expansions of $({a^{\dagger}}^{2}a)^{n}$, $({a^{\dagger}}a{a^{\dagger}})^{n}$, $(a{a^{\dagger}}^{2})^{n}$; at least, when s=1,0,+1s=-1,0,+1. For the case of $({a^{\dagger}}a{a^{\dagger}})^{n}$ when $s=\pm 1$, this has already been done: see the Laguerre-polynomial representations in eqs. (55),(56) above, each term in which is explicitly normal, resp. anti-normal ordered. The expansion of $({a^{\dagger}}a{a^{\dagger}})^{n}$ when $s=\nobreak 0$ is of particular interest: it expresses $({a^{\dagger}}a{a^{\dagger}})^{n}$ as the symmetric (Weyl) quantization of a classical function on the complex $\alpha$-plane.
Proposition 6.2.
One has the Weyl-ordered ($s=\nobreak 0$) representation
$({a^{\dagger}}a{a^{\dagger}})^{n}=\sum_{{0\leqslant k\leqslant n}\atop{\text{$n-k$ even}}}(-1)^{(n-k)/2}\,2^{-(n-k)}\binom{n}{(n-k)/2}\,\frac{n!}{k!}\,\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}({a^{\dagger}}a)^{k}}\,\mathclose{:}_{W}.$
Remark.
The coefficients in this expansion, unlike those in (55),(56), are rational rather than integer-valued.
Proof.
Consider the $(L,R)=(1,1)$ case of Corollary 5.4 when $s=\nobreak 0$, or more concretely the $(L,R)=(1,1)$ case of Theorem 5.5, or the second identity in Theorem 6.1 when $s=\nobreak 0$. They all imply that
$({a^{\dagger}}a{a^{\dagger}})^{n}=\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\,\mathbf{R}[\bar{g}_{T}(z;\,0),\bar{f}_{T}(z;\,0)]_{n}(t={a^{\dagger}}a)}\,\mathclose{:}_{W},$ (90)
where $\bar{g}_{T}(z;0)$ comes from the $(L,R)=(1,1)$ case of (86), and $\bar{f}_{T}(z;0)$ comes similarly from (87). That is,
$({a^{\dagger}}a{a^{\dagger}})^{n}=\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\,\mathbf{R}\left[\frac{1}{\sqrt{1+z^{2}}},\frac{2z}{1+\sqrt{1+z^{2}}}\right]_{n}\!(t={a^{\dagger}}a)}\,\mathclose{:}_{W}.$ (91)
Up to normalization, the Riordan array $\mathbf{R}[d(z),h(z)]$ appearing here is familiar: versions appear as A108044 and A120616 in the OEIS [50]. The expression
$\mathbf{R}_{o}\left[\frac{1}{\sqrt{1+4z^{2}}},\frac{2z}{1+\sqrt{1+4z^{2}}}\right]_{n,k}\!=(-1)^{(n-k)/2}\binom{n}{(n-k)/2}$ (92)
for its matrix elements is known (the binomial coefficient is zero if $n-\nobreak k$ is odd). Here, $\mathbf{R}_{o}$ in distinction to $\mathbf{R}$ signifies an ordinary (non-exponential) Riordan array, meaning that in the definition of $\mathbf{R}_{o}[d(z),h(z)]_{n,k}$ the $n!/k!$ factor in the definition (27) used in this paper is absent. Altering the normalization and restoring the $n!/k!$ factor yields the claimed formula for $({a^{\dagger}}a{a^{\dagger}})^{n}$. ∎
6.3 The case $e=2$
This includes the operators ${\rm e}^{\lambda({a^{\dagger}}^{3}a)}$, ${\rm e}^{\lambda({a^{\dagger}}^{2}a{a^{\dagger}})}$, ${\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}}^{2})}$, ${\rm e}^{\lambda(a{a^{\dagger}}^{3})}$, with $(L,R)=(3,0)$, $(2,1)$, $(1,2)$, $(0,3)$. As with $e=0,1$, one substitutes $(A,B)=(e,1)=(2,1)$ and $L,R$ into the definitions in eq. (60) of the functions $g_{T}(D;s)$ and $z=f_{T}(D;s)$, and computes $\bar{g}_{T}(z;s),\allowbreak\bar{f}_{T}(z;s)]$ by inverting $g_{T},f_{T}]$ in the exponential Riordan group. The $s$-ordered representation of the operator is then $\bar{g}_{T}(\lambda{a^{\dagger}}^{e};s)\times\allowbreak\exp\left[({a^{\dagger}}a)\bar{f}_{T}(\lambda{a^{\dagger}}^{e};s)\right]$.
But for each of the four operators, the representation is quite complicated. The stumbling block is that according to (60), $z=f_{T}(D;s)$ is given by
$z=f_{T}(D;s)=\frac{8D(2-sD)}{\left[(2-sD)^{2}-D^{2}\right]^{2}},$ (93)
and solving for $D=\bar{f}_{T}(z;s)$ entails solving the quartic
$(1-s^{2})^{2}z\,D^{4}+8s(1-s^{2})z\,D^{3}-8\left[(1-3s^{2})z-s\right]D^{2}-16(1+2sz)\,D+16z=0.$ (94)
The solution $D=D(z;s)$ satisfying $D=1\cdot z^{1}+\nobreak O(z^{2})$ can be expressed in radicals by employing the quartic formula, but owing to its complexity it is not given here. If $s=\pm 1$, the quartic (94) degenerates to a more easily solved quadratic, which can be viewed as the reason why the normal and anti-normal orderings of each of the four operators, which are supplied by Theorem 4.4, are relatively simple. But for general $s$, and even when $s=\nobreak 0$ (the Weyl ordering case), there seems to be no simplification.
Whether it is possible to express $s$-ordered representations of operators ${\rm e}^{{a^{\dagger}}^{L}a{a^{\dagger}}^{R}}$ with $e=L+\nobreak R-\nobreak 1>2$ in terms of radicals, for general or even specialized $s$, is a question requiring further study.

### v2

@@
 The $s$-ordered form of any product of single-mode boson creation and annihilation operators, containing only a single annihilator, is computed explicitly.
-The $s$-ordering concept originated in quantum optics, and subsumes normal, symmetric (Weyl), and anti-normal ordering for any two operators satisfying a canonical commutation relation.
+The $s$-ordering concept originated in quantum optics, but subsumes normal, symmetric (Weyl), and anti-normal ordering for any two operators satisfying a canonical commutation relation.
 Because the $s$-ordering map can be viewed as producing a function of a complex variable, its inverse is a quantization map that takes such “classical” functions to quantum operators.
-The explicit $s$-ordered expressions are derived with the aid of a parametric family of Sheffer polynomial sequences (or equivalently an exponential Riordan array of polynomial coefficients), called the Hsu–Shiue family.
-To yield orderings interpolating between normal and anti-normal, this family is extended in an intricate way.
+The explicit $s$-ordered expressions are derived with the aid of a parametric family of Sheffer polynomial sequences (or equivalently a parametric exponential Riordan array of polynomial coefficients), called the Hsu–Shiue family.
+To yield orderings interpolating between normal and anti-normal, this family must be extended.
 keywords:
 Exponential boson operator , Heisenberg–Weyl algebra , operator ordering , s-ordering , normal ordering , Sheffer sequence , Riordan array
-MSC:
-[2020] 81S30 , 81S05 , 05A40
+2020 MSC:
+81S30 , 81S05 , 05A40
 1 Introduction
@@
 In the then new field of quantum optics, Cahill and Glauber [1] derived as identities in $\mathcal{W}[[\lambda]]$, the space of formal power series in $\lambda$ with coefficients in $\mathcal{W}$, the representations
-${\rm e}^{\lambda({a^{\dagger}}a)}=\left\{\begin{aligned} &\mathopen{:}\,\mathrel{\exp\left[\left({\rm e}^{\lambda}-1\right){a^{\dagger}}a\right]}\,\mathclose{:}_{N},\\ &\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}}{a^{\dagger}}a\right]}\,\mathclose{:}_{W},\\ &\mathopen{:}\,\mathrel{{\rm e}^{-\lambda}\exp\left[\left(1-{\rm e}^{\lambda}\right){a^{\dagger}}a\right]}\,\mathclose{:}_{A}.\end{aligned}\right$.
-(1)
+${\rm e}^{\lambda({a^{\dagger}}a)}=\left\{\begin{aligned} &\mathopen{:}\,\mathrel{\exp\left[\left({\rm e}^{\lambda}-1\right){a^{\dagger}}a\right]}\,\mathclose{:}_{N},\\ &\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}}{a^{\dagger}}a\right]}\,\mathclose{:}_{W},\\ &\mathopen{:}\,\mathrel{{\rm e}^{-\lambda}\exp\left[\left(1-{\rm e}^{-\lambda}\right){a^{\dagger}}a\right]}\,\mathclose{:}_{A}.\end{aligned}\right$.
+(1.1)
 […]
@@
 They derived the identity
-${\rm e}^{\lambda({a^{\dagger}}a)}=\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}{a^{\dagger}}a\right]}\,\mathclose{:}_{s},$ (2)
-which subsumes the preceding three.
+${\rm e}^{\lambda({a^{\dagger}}a)}=\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}{a^{\dagger}}a\right]}\,\mathclose{:}_{s},$ (1.2)
+which subsumes the preceding three, and has the quantization
+$\mathopen{:}\,\mathrel{{\rm e}^{\lambda({a^{\dagger}}a)}}\,\mathclose{:}_{s}=\frac{2}{2-\lambda-s\lambda}\left(\frac{2+\lambda-s\lambda}{2-\lambda-s\lambda}\right)^{{a^{\dagger}}a}\!=\left[1+\tfrac{1}{2}(1-s)\lambda\right]^{{a^{\dagger}}a}\left[1-\tfrac{1}{2}(1+s)\lambda\right]^{-a{a^{\dagger}}}$ (1.3)
+as its inverse.
 The $s$-ordering concept has proved useful in quantum mechanics on phase space, and specifically in the quasiprobability approach to quantum optics [2, 3] and other fields of quantum physics [4].
@@
 Then, $F^{(A)}$ can be viewed as a quasiprobability density on $\mathbb{C}\ni\alpha$, from which quantum-mechanical expectations of normal ordered operators, with respect to the state, can be computed by integration in a formally classical way.
-This $F^{(A)}$, which is obtained (by setting $s=\nobreak+1$) from a general $s$-parametrized representation of $\hat{\rho}$, which is denoted by $F^{(s)}$, may not be non-negative on the $\alpha$-plane; in which case it cannot be interpreted as a classical probability density, and the state is said to be non-classical.
+This $F^{(A)}$, which is obtained (by setting $s=\penalty 10000+1$) from a general $s$-parametrized representation of $\hat{\rho}$, which is denoted by $F^{(s)}$, may not be non-negative on the $\alpha$-plane; in which case it cannot be interpreted as a classical probability density, and the state is said to be non-classical.
 But non-classicality can be quantified: as $s$ is lowered from +1+1 toward 0 and 1-1, regions of negativity (if any) of $F^{(s)}$ and its marginals shrink at some rate and vanish [6].
 Quantum optics models in which the time evolution of $s$-parametrized quasiprobability densities has been studied include the widely applicable Jaynes–Cummings model of a two-level atom (a qubit) interacting with a single light mode of a cavity [7], and the Drummond–Walls model of dispersive optical bistability [8].
-Plots of evolving densities, at multiple alternative values of $s$, reveal much about model dynamics.
+Plots of densities evolving with time, at multiple alternative values of $s$, reveal much about model dynamics.
 In this paper, we go beyond previous work on the foundations of $s$-ordering by finding closed-form, analytic expressions of $s$-ordered representations denoted by $F^{(s)}$ (i.e., $s$-dependent functions on the complex $\alpha$-plane), for certain single-mode exponential boson operators of the form ${\rm e}^{\lambda\omega}$, where $\omega\in\mathcal{W}$.
-That is, for each $s$ we find an $F^{(s)}$ with $s$-quantization equal to $\omega$.
-The case most amenable to an analytic treatment is when $\omega$ is a monomial or “boson string”: a product of $a$’s and $a^{\dagger}$}’s, i.e., a word over $\{a,{a^{\dagger}}\}$.
-The “single annihilator” word $\omega={a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ will be treated, where $e=L+\nobreak R-\nobreak 1\geqslant 0$ is the excess of $a^{\dagger}$}’s over $a$’s.
-The Cahill–Glauber result, eq.
-(2) above, is the $s$-representation of ${\rm e}^{\lambda\omega}$ when $(L,R)=(1,0)$ and $\omega={a^{\dagger}}a$.
+That is, for each $s$ we find an $F^{(s)}$ for which the $s$-quantization $\mathopen{:}\,\mathrel{F^{(s)}}\,\mathclose{:}_{s}$ equals ${\rm e}^{\lambda\omega}$.
+[…]
 The chief tool is the theory of Sheffer polynomials [9, 10], or equivalently the theory of exponential Riordan arrays (infinite triangular arrays of polynomial coefficients) [11, 12, 13].
 A Sheffer polynomial sequence is a sequence of polynomials $s_{n}(t)$ with $s_{0}(t)\equiv 1$ and $\deg s_{n}=n$, $n\geqslant 0$, which can be constructed in any of several ways, including the use of differential ladder operators that increment and decrement $n$.
-By extending previous work of Lang [14, 15] and Błasiak and collaborators [16, 17], we first derive normal and anti-normal representations that subsume the first and third identities in (1), namely
-$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,-e,1,R)}\,\mathclose{:}_{N}\\ &=\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl{(}1-e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}\end{aligned}$ (3)
+By extending previous work of Lang [14, 15] and Błasiak and collaborators [16, 17], we first derive normal and anti-normal representations that subsume the first and third identities in (1.1), namely
+$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,-e,1,R)}\,\mathclose{:}_{N}\\ &=\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl(1-e\lambda{a^{\dagger}}^{e}\bigr)^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}\end{aligned}$ (1.4)
 and
-$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}(a{a^{\dagger}},\lambda{a^{\dagger}}^{e};\,e,-1,-L)}\,\mathclose{:}_{A}\\ &=\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl{(}1+e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A},\end{aligned}$ (4)
+$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}(a{a^{\dagger}},\lambda{a^{\dagger}}^{e};\,e,-1,-L)}\,\mathclose{:}_{A}\\ &=\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl(1+e\lambda{a^{\dagger}}^{e}\bigr)^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A},\end{aligned}$ (1.5)
 […]
-Our main result (Theorem 5.3) is the explicit $s$-ordering of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$, valid for arbitrary $s\in\mathbb{C}$.
-It subsumes eqs.
-(3) and (4), which are its $s=\nobreak-1$ and $s=\nobreak+1$ special cases, and includes a new and explicit Weyl ordering formula ($s=\nobreak 0$); for the latter, see Theorem 5.5.
-The general $s$-ordered representation is expressed in terms of what will be called a two-point Hsu–Shiue EGF, interpolating in an intricate sense between the EGF’s in (3) and (4).
+[…]
 Many examples, including a full treatment of the case $e=1$, are given.
@@
 Theorem 4.4 subsumes eqs.
-(3) and (4) above, and serves as a lemma in the derivation by interpolation of the main result, in Sec.
+(1.4) and (1.5) above, and serves as a lemma in the derivation by interpolation of the main result, in Sec.
 5.
@@
 $\mathcal{W}$ has a unit (multiplicative identity), written $1$, which is the empty product multiplied by the scalar $1$.
-The representation as polynomials in $a,{a^{\dagger}}$ is not unique, as the product on $\mathcal{W}$ satisfies $a,{a^{\dagger}}]=1$.
-$\mathcal{W}$ can be identified with $\mathbb{C}\langle a,{a^{\dagger}}\rangle/\allowbreak(a{a^{\dagger}}-\nobreak{a^{\dagger}}a-\nobreak 1)$, the quotient of the free algebra $\mathbb{C}\langle a,{a^{\dagger}}\rangle$ by the two-sided ideal generated by $a{a^{\dagger}}-\nobreak\nobreak{a^{\dagger}}a-\nobreak 1$.
-Thus each element of $\mathcal{W}$ is an equivalence class of polynomials in $a,{a^{\dagger}}$, though a distinguished representative can be singled out by imposing an ordering criterion.
+[…]
 One can also view $\mathcal{W}$ as $\mathbb{C}[{a^{\dagger}}][a]$, the ring of polynomials in $a$ with coefficients that are polynomials in $a^{\dagger}$}, equipped with a product operation that comes from $a,{a^{\dagger}}]=1$.
@@
 Let $G\in\mathcal{W}$, i.e., let $G(a,{a^{\dagger}})$ be a polynomial in the noncommuting $a,{a^{\dagger}}$ that represents $G$ (and is not unique).
-One can define two c-number or classical representations (sometimes called transforms) associated to $G$, denoted by $F^{(N)},F^{(A)}\in\mathcal{W}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$, the ring of polynomials in commuting indeterminates $\alpha,\alpha^{*}$, which is $\mathbb{C}\langle\alpha,\alpha^{*}\rangle/\allowbreak(\alpha\alpha^{*}-\nobreak\alpha^{*}\alpha)$.
+One can define two c-number or classical representations (sometimes called transforms) associated to $G$, denoted by $F^{(N)},F^{(A)}\in\mathcal{W}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$, the ring of polynomials in commuting indeterminates $\alpha,\alpha^{*}$, which is $\mathbb{C}\langle\alpha,\alpha^{*}\rangle/\allowbreak(\alpha\alpha^{*}-\penalty 10000\alpha^{*}\alpha)$.
 They are defined by requiring that
-$G(a,{a^{\dagger}})$ $=\mathopen{:}\,\mathrel{F^{(N)}(\alpha,\alpha^{*})}\,\mathclose{:}_{N},$ (5a)
-$G(a,{a^{\dagger}})$ $=\mathopen{:}\,\mathrel{F^{(A)}(\alpha,\alpha^{*})}\,\mathclose{:}_{A},$ (5b)
+$G(a,{a^{\dagger}})$ $=\mathopen{:}\,\mathrel{F^{(N)}(\alpha,\alpha^{*})}\,\mathclose{:}_{N},$ (2.1a)
+$G(a,{a^{\dagger}})$ $=\mathopen{:}\,\mathrel{F^{(A)}(\alpha,\alpha^{*})}\,\mathclose{:}_{A},$ (2.1b)
 as equalities in $\mathcal{W}$.
@@
 Here, the Weyl quantization $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{W}$ maps any monomial ${\alpha^{*}}^{n}\alpha^{m}$ to the average of the $\binom{n+m}{n}=\binom{n+m}{m}$ distinct orderings of a product of $n$ $a^{\dagger}$}’s and $m$ $a$’s.
-(An equivalent definition is that in $\mathcal{W}$, $\mathopen{:}\,\mathrel{(\mu\alpha+\nobreak\nu\alpha^{*})^{p}}\,\mathclose{:}_{W}$ must equal $(\mu a+\nobreak\nu{a^{\dagger}})^{p}$.) For instance,
-$\mathopen{:}\,\mathrel{{\alpha^{*}}^{2}\alpha}\,\mathclose{:}_{W}=({a^{\dagger}}^{2}a+{a^{\dagger}}a{a^{\dagger}}+a{a^{\dagger}}^{2})/3,$ (6)
-where the right-hand side, an element of $\mathcal{W}$, can be rewritten with the aid of $a,{a^{\dagger}}]=1$ in various ways, including ${a^{\dagger}}^{2}a+\nobreak{a^{\dagger}}$, ${a^{\dagger}}a{a^{\dagger}}$, $a{a^{\dagger}}^{2}-\nobreak{a^{\dagger}}$.
+(An equivalent definition is that in $\mathcal{W}$, $\mathopen{:}\,\mathrel{(\mu\alpha+\penalty 10000\nu\alpha^{*})^{p}}\,\mathclose{:}_{W}$ must equal $(\mu a+\penalty 10000\nu{a^{\dagger}})^{p}$.) For instance,
+$\mathopen{:}\,\mathrel{{\alpha^{*}}^{2}\alpha}\,\mathclose{:}_{W}=({a^{\dagger}}^{2}a+{a^{\dagger}}a{a^{\dagger}}+a{a^{\dagger}}^{2})/3,$ (2.2)
+where the right-hand side, an element of $\mathcal{W}$, can be rewritten with the aid of $a,{a^{\dagger}}]=1$ in various ways, including ${a^{\dagger}}^{2}a+\penalty 10000{a^{\dagger}}$, ${a^{\dagger}}a{a^{\dagger}}$, $a{a^{\dagger}}^{2}-\penalty 10000{a^{\dagger}}$.
 So
 $G(a,{a^{\dagger}})={a^{\dagger}}a{a^{\dagger}}\quad\Longrightarrow\quad\left\{\begin{aligned} F^{(N)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha+\alpha^{*},\\ F^{(W)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha,\\ F^{(A)}(\alpha,\alpha^{*})&={\alpha^{*}}^{2}\alpha-\alpha^{*},\end{aligned}\right$.
-(7)
+(2.3)
 and if $G(a,{a^{\dagger}})={a^{\dagger}}^{2}a$, then $\alpha^{*}$ would be subtracted from each $F(\alpha,\alpha^{*})$, etc.
 It is useful to define a formal power series $D(\beta,\beta^{*})\in\mathcal{W}[[\beta,\beta^{*}]]$ by
-$D(\beta,\beta^{*})$ $=\exp(\beta{a^{\dagger}}-\beta^{*}a)=\sum_{p=0}^{\infty}\frac{1}{p!}(\beta{a^{\dagger}}-\beta^{*}a)^{p}$ (8)
+$D(\beta,\beta^{*})$ $=\exp(\beta{a^{\dagger}}-\beta^{*}a)=\sum_{p=0}^{\infty}\frac{1}{p!}(\beta{a^{\dagger}}-\beta^{*}a)^{p}$ (2.4)
 $=\sum_{n,m=0}^{\infty}\frac{\beta^{n}(-\beta^{*})^{m}}{n!\,m!}\,\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{W},$
 so that
-$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{W}=\frac{\partial^{n+m}}{\partial\beta^{n}\partial(-\beta^{*})^{m}}D(\beta,\beta^{*})\Bigm{|}_{(\beta,\beta^{*})=(0,0)}.$ (9)
+$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{W}=\frac{\partial^{n+m}}{\partial\beta^{n}\partial(-\beta^{*})^{m}}D(\beta,\beta^{*})\Bigm|_{(\beta,\beta^{*})=(0,0)}.$ (2.5)
 In quantum optics, $D(\beta,\beta^{*})$ appears as the displacement operator.
@@
 (Not least, in the exponentiation of $\mathcal{W}$ to the complexified Heisenberg–Weyl group [20].) For greater rigor, one could instead introduce commuting quantities $\alpha_{r},\alpha_{i}$ and $\beta_{r},\beta_{i}$, the real and imaginary parts of $\alpha$ and $\beta$, to serve as real coordinates on the complex $\alpha$- and $\beta$-planes [2].
-Define an $s$-parametrized deformation of the exponential operator $D(\beta,\beta^{*})$, namely $D(\beta,\beta^{*};s)={\rm e}^{s\beta(-\beta^{*})/2})D(\beta,\beta^{*})$, where $s\in\mathbb{C}$, and for each $n,m$, define the $s$-quantization $\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}$ of ${\alpha^{*}}^{n}\alpha^{m}$ by
-$D(\beta,\beta^{*};s)=\sum_{n,m=0}^{\infty}\frac{\beta^{n}(-{\beta^{*}})^{m}}{n!\,m!}\,\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s},$ (10)
+Define an $s$-parametrized deformation of the exponential operator $D(\beta,\beta^{*})$, namely $D(\beta,\beta^{*};s)={\rm e}^{s\beta(-\beta^{*})/2}D(\beta,\beta^{*})$, where $s\in\mathbb{C}$; and for each $n,m$, define the $s$-quantization $\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}$ of ${\alpha^{*}}^{n}\alpha^{m}$ by
+$D(\beta,\beta^{*};s)=\sum_{n,m=0}^{\infty}\frac{\beta^{n}(-{\beta^{*}})^{m}}{n!\,m!}\,\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s},$ (2.6)
 so that
-$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}=\frac{\partial^{n+m}}{\partial\beta^{n}\partial(-\beta^{*})^{m}}D(\beta,\beta^{*};s)\Bigm{|}_{(\beta,\beta^{*})=(0,0)}.$ (11)
+$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}=\frac{\partial^{n+m}}{\partial\beta^{n}\partial(-\beta^{*})^{m}}D(\beta,\beta^{*};s)\Bigm|_{(\beta,\beta^{*})=(0,0)}.$ (2.7)
 Equivalently, one requires that
-$D(\beta,\beta^{*})=\exp(\beta{a^{\dagger}}-\beta^{*}a)={\rm e}^{s|\beta|^{2}}\mathopen{:}\,\mathrel{{\rm e}^{\beta\alpha^{*}-\beta^{*}\alpha}}\,\mathclose{:}_{s},$ (12)
+$D(\beta,\beta^{*})=\exp(\beta{a^{\dagger}}-\beta^{*}a)={\rm e}^{s|\beta|^{2}/2}\mathopen{:}\,\mathrel{{\rm e}^{\beta\alpha^{*}-\beta^{*}\alpha}}\,\mathclose{:}_{s},$ (2.8)
 as an equality in $\mathcal{W}[[\beta,\beta^{*}]]$.
-As an example of (11), one finds with the aid of $a,{a^{\dagger}}]=1$ that
-$\mathopen{:}\,\mathrel{{\alpha^{*}}^{2}\alpha}\,\mathclose{:}_{s}={a^{\dagger}}a{a^{\dagger}}+s\,{a^{\dagger}}.$ (13)
+As an example of (2.7), one finds with the aid of $a,{a^{\dagger}}]=1$ that
+$\mathopen{:}\,\mathrel{{\alpha^{*}}^{2}\alpha}\,\mathclose{:}_{s}={a^{\dagger}}a{a^{\dagger}}+s\,{a^{\dagger}}.$ (2.9)
 The $s$-quantization operator $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{s}$ extends by linearity from monomials in the (commuting) indeterminates $\alpha,\alpha^{*}$, to polynomials in same, i.e., to elements of ${\mathcal{W}}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$, and to elements of $\mathcal{W}_{c}[[\lambda]]=\mathbb{C}[\alpha,\alpha^{*}][[\lambda]]$, as well.
 Thus, one can define a unique $s$-ordered representation or $s$-transform of any $G=G(a,{a^{\dagger}})$ in $\mathcal{W}$, denoted by $F^{(s)}(\alpha,\alpha^{*})$, by requiring that
-$G(a,{a^{\dagger}})=\mathopen{:}\,\mathrel{F^{(s)}(\alpha,\alpha^{*})}\,\mathclose{:}_{s}.$ (14)
+$G(a,{a^{\dagger}})=\mathopen{:}\,\mathrel{F^{(s)}(\alpha,\alpha^{*})}\,\mathclose{:}_{s}.$ (2.10)
 So, for example,
-$G(a,{a^{\dagger}})={a^{\dagger}}a{a^{\dagger}}\quad\Longrightarrow\quad F^{(s)}(\alpha,\alpha^{*})={\alpha^{*}}^{2}\alpha-s\,\alpha^{*}.$ (15)
-When $s=\nobreak 0$, $F^{(s)}$ reduces to the Weyl-ordered representation $F^{(W)}$. It is also not difficult to show, using the Baker–Campbell–Hausdorff formula
-${\rm e}^{A}{\rm e}^{B}={\rm e}^{A+B+[A,B]/2},$ (16)
-which holds as an equality between power series if $[A,B]$ commutes with $A$ and $B$, that when $s=\nobreak-1$ and $s=\nobreak+1$, $F^{(s)}$ reduces respectively to the normal and anti-normal ordered representations, $F^{(N)}$ and $F^{(A)}$.
-The s=1,0,+1s=-1,0,+1 cases of (15) are reflected in eq.
-(7).
+$G(a,{a^{\dagger}})={a^{\dagger}}a{a^{\dagger}}\quad\Longrightarrow\quad F^{(s)}(\alpha,\alpha^{*})={\alpha^{*}}^{2}\alpha-s\,\alpha^{*}.$ (2.11)
+[…]
+${\rm e}^{A}{\rm e}^{B}={\rm e}^{A+B+[A,B]/2},$ (2.12)
+which holds as an equality between power series if $[A,B]$ commutes with $A$ and $B$, that when $s=\penalty 10000-1$ and $s=\penalty 10000+1$, $F^{(s)}$ reduces respectively to the normal and anti-normal ordered representations, $F^{(N)}$ and $F^{(A)}$.
+The s=1,0,+1s=-1,0,+1 cases of (2.11) are reflected in eq.
+(2.3).
 In modern quantum optics, the most commonly encountered c-number representations of any element of the Heisenberg–Weyl algebra are the classical functions $F^{(-1)}=F^{(N)}$, $F^{(0)}=F^{(W)}$, $F^{(+1)}=F^{(A)}$ on $\mathbb{C}$ that arise from $G=\hat{\rho}$, where $\hat{\rho}$ is a specified (single-mode) density operator that is an element of $\mathcal{W}$, or an extension of $\mathcal{W}$ that includes formal power series in $a,{a^{\dagger}}$ [2, 21, 22].
@@
 In the original paper of Cahill–Glauber [1] and a few more recent ones [26, 27, 28], s=+1,1s=+1,-1 correspond to normal and anti-normal ordering; e.g., for a density operator, to $Q$ and $P$.
-But usually, they correspond to anti-normal and normal ordering (i.e., to $P$ and $Q$).
+But usually, they correspond to anti-normal and normal ordering (e.g., to $P$ and $Q$).
 See, e.g., [2, Chapter 12], [6, 29], and [3, Chapter 3].
 The latter convention has been adopted here, so the sign of $s$ in eq.
-(2), taken from [1], has been changed.
+(1.2), taken from [1], has been changed.
 Note also that in some works (e.g., [21] and [24, Table V]), the meanings of the so-called reciprocal orderings $F^{(N)}$ and $F^{(A)}$, related by negation of $s$, are interchanged.)
-Differentiating the exponential operator $D(\beta,\beta^{*};s)={\rm e}^{(s^{\prime}-s)|\beta|^{2}/2}D(\beta,\beta^{*};s^{\prime})$ with respect to $\beta,\beta^{*}$, in a procedure that can be traced to Wilcox [30], and using (11), yields the $s$-ordering conversion formula
-$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}=\sum_{k=0}^{\min(m,n)}k!\,\binom{n}{k}\binom{m}{k}\left(\frac{s-s^{\prime}}{2}\right)^{k}\mathopen{:}\,\mathrel{{\alpha^{*}}^{n-k}\alpha^{m-k}}\,\mathclose{:}_{s^{\prime}}.$ (17)
-The simplest case of (17) is the contraction
-$\mathopen{:}\,\mathrel{\alpha^{*}\alpha}\,\mathclose{:}_{s}-\mathopen{:}\,\mathrel{\alpha^{*}\alpha}\,\mathclose{:}_{s^{\prime}}=\frac{s-s^{\prime}}{2},$ (18)
-from which the general formula (17) can be derived, much as Wick’s theorem is derived [28]. As another example, the $s=\nobreak+1$, $s^{\prime}=-1$ case of (17) is
-$a^{m}{a^{\dagger}}^{n}=\sum_{k=0}^{\min(m,n)}k!\,\binom{n}{k}\binom{m}{k}{a^{\dagger}}^{n-k}a^{m-k},$ (19)
-an identity in $\mathcal{W}$ that converts an anti-normal ordering ($s=\nobreak+1$) to a normal one ($s=\nobreak-1$).
-Thus one can convert between the coefficients of an anti-normal and a normal series, $\sum_{m,n}f_{m,n}^{(A)}a^{m}{a^{\dagger}}^{n}$ and $\sum_{m,n}f_{n,m}^{(N)}{a^{\dagger}}^{n}a^{m}$ [31].
-Equation (17) can be written in terms of a two-variable Hermite polynomial [27, 28], as was recognized early (when $s\in\{0,\pm 1\}$) by Agarwal and Wolf [24, Table V]. It can also be written as
-$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}$ $=\sum_{k=0}^{\min(m,n)}\frac{1}{k!}\left(\frac{s-s^{\prime}}{2}\right)^{k}\mathopen{:}\,\mathrel{\left(\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right)^{k}{{\alpha^{*}}^{n}\alpha^{m}}}\,\mathclose{:}_{s^{\prime}}$ (20)
+Differentiating the exponential operator $D(\beta,\beta^{*};s)={\rm e}^{(s^{\prime}-s)|\beta|^{2}/2}D(\beta,\beta^{*};s^{\prime})$ with respect to $\beta,\beta^{*}$, in a procedure that can be traced to Wilcox [30], and using (2.7), yields the $s$-ordering conversion formula
+$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}=\sum_{k=0}^{\min(m,n)}k!\,\binom{n}{k}\binom{m}{k}\left(\frac{s-s^{\prime}}{2}\right)^{k}\mathopen{:}\,\mathrel{{\alpha^{*}}^{n-k}\alpha^{m-k}}\,\mathclose{:}_{s^{\prime}}.$ (2.13)
+The $n=m=1$ case of (2.13) is the contraction
+$\mathopen{:}\,\mathrel{\alpha^{*}\alpha}\,\mathclose{:}_{s}-\mathopen{:}\,\mathrel{\alpha^{*}\alpha}\,\mathclose{:}_{s^{\prime}}=\frac{s-s^{\prime}}{2},$ (2.14)
+from which the general formula (2.13) can be derived, much as Wick’s theorem is derived [28]. As another example, the $s=\penalty 10000+1$, $s^{\prime}=-1$ case of (2.13) is
+$a^{m}{a^{\dagger}}^{n}=\sum_{k=0}^{\min(m,n)}k!\,\binom{n}{k}\binom{m}{k}{a^{\dagger}}^{n-k}a^{m-k},$ (2.15)
+an identity in $\mathcal{W}$ that converts an anti-normal ordering ($s=\penalty 10000+1$) to a normal one ($s=\penalty 10000-1$).
+Thus, one can convert between the coefficients of an anti-normal and a normal series, $\sum_{m,n}f_{m,n}^{(A)}a^{m}{a^{\dagger}}^{n}$ and $\sum_{m,n}f_{n,m}^{(N)}{a^{\dagger}}^{n}a^{m}$ [31].
+Equation (2.13) can be written in terms of a two-variable Hermite polynomial [27, 28], as was recognized early (when $s\in\{0,\pm 1\}$) by Agarwal and Wolf [24, Table V]. It can also be written as
+$\mathopen{:}\,\mathrel{{\alpha^{*}}^{n}\alpha^{m}}\,\mathclose{:}_{s}$ $=\sum_{k=0}^{\min(m,n)}\frac{1}{k!}\left(\frac{s-s^{\prime}}{2}\right)^{k}\mathopen{:}\,\mathrel{\left(\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right)^{k}{{\alpha^{*}}^{n}\alpha^{m}}}\,\mathclose{:}_{s^{\prime}}$ (2.16)
 $=\mathopen{:}\,\mathrel{\exp\left[\left(\frac{s-s^{\prime}}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]{{\alpha^{*}}^{n}\alpha^{m}}}\,\mathclose{:}_{s^{\prime}}.$
 This extends by linearity to
-$\mathopen{:}\,\mathrel{F}\,\mathclose{:}_{s}=\mathopen{:}\,\mathrel{\exp\left[\left(\frac{s-s^{\prime}}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]F}\,\mathclose{:}_{s^{\prime}}$ (21)
+$\mathopen{:}\,\mathrel{F}\,\mathclose{:}_{s}=\mathopen{:}\,\mathrel{\exp\left[\left(\frac{s-s^{\prime}}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]F}\,\mathclose{:}_{s^{\prime}}$ (2.17)
 for any $F\in\mathcal{W}_{c}$, i.e., for any polynomial $F=F(\alpha,\alpha^{*})$.
-Suppose now that $F^{(s)},F^{(s^{\prime})}\in{\mathcal{W}}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$ are the $s$-ordered and $s^{\prime}$-ordered “c-number” versions of some $G\in\mathcal{W}$, computed from $G$ by the requirement (14).
-By (20), they must be related to each other by the dual relation
-$F^{(s)}(\alpha,\alpha^{*})=\exp\left[\left(\frac{s^{\prime}-s}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]F^{(s^{\prime})}(\alpha,\alpha^{*}),$ (22)
+Suppose now that $F^{(s)},F^{(s^{\prime})}\in{\mathcal{W}}_{c}=\mathbb{C}[\alpha,\alpha^{*}]$ are the $s$-ordered and $s^{\prime}$-ordered “c-number” versions of some $G\in\mathcal{W}$, computed from $G$ by the requirement (2.10).
+By (2.16), they must be related to each other by the dual relation
+$F^{(s)}(\alpha,\alpha^{*})=\exp\left[\left(\frac{s^{\prime}-s}{2}\right)\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}\right]F^{(s^{\prime})}(\alpha,\alpha^{*}),$ (2.18)
 in which the exponentiated second-order differential operator has been inverted.
-In the formally infinite series in (22) obtained by expanding the exponential in a power series, only a finite number of terms will be nonzero.
+In the formally infinite series in (2.18) obtained by expanding the exponential in a power series, only a finite number of terms will be nonzero.
 Equivalently, $F=F^{(s)}(\alpha,\alpha^{*})$ satisfies the diffusion or heat equation
-$\frac{\partial F}{\partial(-s)}=\frac{1}{2}\,\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}F,$ (23)
-an unexponentiated version of (22). (See [2, Chapter 12].)
-Equation (22) can be extended to the case when $F^{(s)},F^{(s^{\prime})}$ are convergent power series in $\alpha,\alpha^{*}$ rather than polynomials, at least formally.
-The nicest case is when each of $F^{(s)},F^{(s^{\prime})}$, regarded as a Maclaurin expansion at $(\alpha,\alpha^{*})=(0,0)$, i.e., $(\alpha_{r},\alpha_{i})=(0,0)$, converges globally and defines an analytic function of the real variables $\alpha_{r},\alpha_{i}$, or even a complex analytic function of $\alpha=\alpha_{r}+\nobreak{\rm i}\alpha_{i}$.
+$\frac{\partial F}{\partial(-s)}=\frac{1}{2}\,\frac{\partial^{2}}{\partial\alpha\partial\alpha^{*}}F,$ (2.19)
+an unexponentiated version of (2.18). (See [2, Chapter 12].)
+Equation (2.18) can be extended to the case when $F^{(s)},F^{(s^{\prime})}$ are convergent power series in $\alpha,\alpha^{*}$ rather than polynomials, at least formally.
+The nicest case is when each of $F^{(s)},F^{(s^{\prime})}$, regarded as a Maclaurin expansion at $(\alpha,\alpha^{*})=(0,0)$, i.e., $(\alpha_{r},\alpha_{i})=(0,0)$, converges globally and defines an analytic function of the real variables $\alpha_{r},\alpha_{i}$, or even a complex analytic function of $\alpha=\alpha_{r}+\penalty 10000{\rm i}\alpha_{i}$.
 Because $\partial^{2}/\partial\alpha\partial\alpha^{*}$ equals $\Delta/4$, where $\Delta=\partial^{2}/\partial\alpha_{r}^{2}+\partial^{2}/\partial\alpha_{i}^{2}$ is the Laplacian on the complex $\alpha$-plane coordinatized by $\alpha_{r},\alpha_{i}$, one has
-$F^{(s)}(\alpha,\alpha^{*})$ $={\rm e}^{(s^{\prime}-s)\Delta/8}F^{(s^{\prime})}(\alpha,\alpha^{*})$ (24)
+$F^{(s)}(\alpha,\alpha^{*})$ $={\rm e}^{(s^{\prime}-s)\Delta/8}F^{(s^{\prime})}(\alpha,\alpha^{*})$ (2.20)
 $\varpropto\int F^{(s^{\prime})}(\beta,\beta^{*})\,{\rm e}^{-2|\beta-\alpha|^{2}/(s^{\prime}-s)}\,{\rm d}^{2}\beta.$
-So in this case, when $s<s^{\prime}$ it may be possible to compute $F^{(s)}$ from $F^{(s^{\prime})}$ by a smoothing Weierstrass transform operation on the complex $\alpha$-plane: convolution with a Gaussian kernel of variance $(s^{\prime}-\nobreak s)/4$.
-This operation is familiar from quantum optics [2, 3, 4, 20]: among the quasiprobability densities $F^{(s)}$ obtained when $G=G(a,{a^{\dagger}})$ is a density operator $\hat{\rho}$, the Wigner $W$ function ($s=\nobreak 0$) is a smoothed version of the Glauber–Sudarshan $P$ function ($s=\nobreak+1$), and the Husimi $Q$ function ($s=\nobreak-1$) is a smoothed version of the $W$ function.
-The smoothing of $P$ to $W$ and $Q$ has been extended from functions of $\alpha\in\mathbb{C}$ to tempered and other distributions on the complex $\alpha$-plane, and the variation with $s$ of the mathematical properties and physical interpretation of smoothed quasiprobability densities has been explored [2, 6, 20].
-In the proof of the main result of this paper (Theorem 5.3), the unexponentiated form (23), and not the exponentiated forms (22),(24), will be employed to relate $F^{(s)}$, $F^{(s^{\prime})}\in\mathcal{W}_{c}[[\lambda]]$, to obviate consideration of series convergence.
+So in this case, when $s<s^{\prime}$ it may be possible to compute $F^{(s)}$ from $F^{(s^{\prime})}$ by a smoothing Weierstrass transform operation on the complex $\alpha$-plane: convolution with a Gaussian kernel of variance $(s^{\prime}-\penalty 10000s)/4$.
+[…]
+In the proof of the main result of this paper (Theorem 5.3), the unexponentiated form (2.19), and not the exponentiated forms (2.18),(2.20), will be employed to relate $F^{(s)}$, $F^{(s^{\prime})}\in\mathcal{W}_{c}[[\lambda]]$, to obviate consideration of series convergence.
 3 Riordan, Sheffer, and Hsu–Shiue polynomials
@@
 Let $d,h\in\mathbb{C}[[z]]$ be formal power series in $z$ of order 0 and $1$ respectively, meaning that their first nonzero terms are the $z^{0}$ and $z^{1}$ terms; so $1/d$ and the compositional inverse $\bar{h}$ both exist as formal power series.
-(For instance, $d,h$ may be Maclaurin expansions of functions $d(z),h(z)$ analytic at $z=0$.) One usually requires that $d$ be monic, $d=1\cdot z^{0}+\nobreak O(z^{1})$, and often that $h=1\cdot z^{1}+\nobreak O(z^{2})$.
+(For instance, $d,h$ may be Maclaurin expansions of functions $d(z),h(z)$ analytic at $z=0$.) One usually requires that $d$ be monic, $d=1\cdot z^{0}+\penalty 10000O(z^{1})$, and often that $h=1\cdot z^{1}+\penalty 10000O(z^{2})$.
 Then, a polynomial sequence $s_{n}(t)$, $n\geqslant 0$, where $\deg s_{n}=n$, with EGF $\mathcal{S}(t,z)=d(z){\rm e}^{th(z)}$ that is an element of $\mathbb{C}[t][[z]]$, is defined by
-$\sum_{n=0}^{\infty}s_{n}(t)\frac{z^{n}}{n!}=\mathcal{S}(t,z),$ (25)
+$\sum_{n=0}^{\infty}s_{n}(t)\frac{z^{n}}{n!}=\mathcal{S}(t,z),$ (3.1)
 with $s_{0}(t)\equiv 1$.
-The sequence $s_{n}(t)$ is said to be exponential Riordan for the ordered pair $[d,h]$, and is denoted briefly by $\mathbf{R}[d,h]$.
+The sequence $s_{n}(t)$ is said to be exponential Riordan for the ordered pair $[d,h]$, and will be denoted briefly by $\mathbf{R}[d,h]$.
 Such notations as $s_{n}(t)=\mathbf{R}[d,h]_{n}(t)$ and $\mathcal{S}(t,z)=\mathbf{R}[d,h](t,z)$ are also used.
 Equivalently, one can define a triangular array of coefficients $s_{n,k}$, $0\leqslant k\leqslant n$, by
-$\sum_{n=0}^{\infty}\sum_{k=0}^{n}s_{n,k}t^{k}\frac{z^{n}}{n!}=\mathcal{S}(t,z),$ (26)
+$\sum_{n=0}^{\infty}\sum_{k=0}^{n}s_{n,k}t^{k}\frac{z^{n}}{n!}=\mathcal{S}(t,z),$ (3.2)
 so that
-$s_{n,k}=\frac{n!}{k!}\,[z^{n}]\left[d(z)h(z)^{k}\right],$ (27)
+$s_{n,k}=\frac{n!}{k!}\,[z^{n}]\left[d(z)h(z)^{k}\right],$ (3.3)
 where $z^{n}]$ extracts the coefficient of $z^{n}$ from the element of $\mathbb{C}[[z]]$ that follows.
@@
 The array $s_{n,k}$ can be viewed as the lower-triangular part of an infinite matrix, and is called an exponential Riordan array.
-The notation $s_{n,k}=\mathbf{R}[d,h]_{n,k}$ is also used, with $\mathbf{R}[d,h]$ alternatively signifying the matrix as a whole.
+The notation $s_{n,k}=\mathbf{R}[d,h]_{n,k}$ will also be used, with $\mathbf{R}[d,h]$ alternatively signifying the matrix as a whole.
 Row-finite lower-triangular matrices form a group under matrix multiplication, and exponential Riordan arrays turn out to form a subgroup, which will be denoted by $\mathfrak{R}$.
@@
 Proposition 3.1 ([13]).
-In the exponential Riordan group $\mathfrak{R}$ the product operation is $\mathbf{R}[d_{1},h_{1}]\,\mathbf{R}[d_{2},h_{2}]=\allowbreak\mathbf{R}[(d_{2}\circ\nobreak h_{1})d_{1},\allowbreak h_{2}\circ\nobreak h_{1}]$, and the inversion operation is $\mathbf{R}[d,h]^{-1}=\allowbreak\mathbf{R}[1/(d\circ\nobreak\bar{h}),\bar{h}]$, where $\bar{h}$ is the compositional inverse of $h$, i.e., $\bar{h}(h(z))=\nobreak z$.
+In the exponential Riordan group $\mathfrak{R}$ the product operation is $\mathbf{R}[d_{1},h_{1}]\,\mathbf{R}[d_{2},h_{2}]=\allowbreak\mathbf{R}[(d_{2}\circ\penalty 10000h_{1})d_{1},\allowbreak h_{2}\circ\penalty 10000h_{1}]$, and the inversion operation is $\mathbf{R}[d,h]^{-1}=\allowbreak\mathbf{R}[1/(d\circ\penalty 10000\bar{h}),\bar{h}]$, where $\bar{h}$ is the compositional inverse of $h$, i.e., $\bar{h}(h(z))=\penalty 10000z$.
 The multiplicative identity is $[1,z]$, which signifies the infinite identity matrix.
 For matrix inverses, the notation
-$\mathbf{R}[d,h]^{-1}\mathrel{=:}\mathbf{R}\overline{[d,h]}=\mathbf{R}[\bar{d},\bar{h}],$ (28)
+$\mathbf{R}[d,h]^{-1}\mathrel{=:}\mathbf{R}\overline{[d,h]}=\mathbf{R}[\bar{d},\bar{h}],$ (3.4)
 where $\bar{d}=1/(d\circ\bar{h})$, can be used. Note that $\bar{d},\bar{h}$, like $d,h$, are formal power series of orders 0 and $1$.
-The generating functions of the theory facilitate the computation of matrix–vector products.
+A quantum-optics example of inversion in $\mathfrak{R}$ is afforded by the Cahill–Glauber identity (1.2) and its inverse (1.3): expansion in the parameter $\lambda$ reveals that they are based respectively on the arrays
+$\mathbf{R}\left[\frac{2}{1+{\rm e}^{z}-s(1-{\rm e}^{z})},\frac{2({\rm e}^{z}-1)}{1+{\rm e}^{z}-s(1-{\rm e}^{z})}\right],\quad\mathbf{R}\left[\frac{2}{2-z-sz},\ln\left(\frac{2+z-sz}{2-z-sz}\right)\right]$ (3.5)
+in $\mathfrak{R}$, which are inverses of each other.
+The generating functions of the Riordan-array theory facilitate the computation of matrix–vector products, as follows.
 Proposition 3.2 ([13]).
 For any column vector $u=(u_{k})$ with EGF $u(t)=\sum_{k=0}^{\infty}u_{k}t^{k}/k!$, an element of $\mathbb{C}[[t]]$, and for any exponential Riordan array $s_{n,k}$ equaling $\mathbf{R}[d,h]_{n,k}$, define the column vector $v=(v_{n})$ by
-$v_{n}=\sum_{k=0}^{n}s_{n,k}u_{k},$ (29)
+$v_{n}=\sum_{k=0}^{n}s_{n,k}u_{k},$ (3.6)
 or symbolically by $v=\mathbf{R}[d,h]u$.
@@
 So a polynomial sequence $s_{n}$ is Sheffer for $[g,f]$, where $g,f\in\mathbb{C}[[D]]$ are of order 0 and $1$ respectively, and $\bar{g},\bar{f}\in\mathbb{C}[[z]]$ are likewise, if
-$\sum_{n=0}^{\infty}s_{n}(t)\frac{z^{n}}{n!}=\bar{g}(z){\rm e}^{t\bar{f}(z)}=\frac{1}{g(\bar{f}(z))}{\rm e}^{t\bar{f}(z)},$ (30)
+$\sum_{n=0}^{\infty}s_{n}(t)\frac{z^{n}}{n!}=\bar{g}(z){\rm e}^{t\bar{f}(z)}=\frac{1}{g(\bar{f}(z))}{\rm e}^{t\bar{f}(z)},$ (3.7)
 and one writes $s_{n}=\mathbf{S}[g,f]_{n}=\mathbf{R}[\bar{g},\bar{f}]_{n}$ and $s_{n,k}=\mathbf{S}[g,f]_{n,k}=\mathbf{R}[\bar{g},\bar{f}]_{n,k}$, etc.
-Many well-known polynomial sequences can be viewed as Sheffer sequences [9, 38], and there is some overlap with the orthogonal polynomials of theoretical physics.
-An example is the sequence of Hermite polynomials $\mathop{{}\it He}\nolimits_{n}(t)$, which is $\mathbf{S}[{\rm e}^{D^{2}/2},D]$ or equivalently $\mathbf{R}[{\rm e}^{-z^{2}/2},z]$.
-The $\mathop{{}\it He}\nolimits_{n}(t)$ are orthogonal with respect to an appropriate measure (namely, ${\rm e}^{-t^{2}/2}\,{\rm d}t$), as they happen to satisfy a three-term recurrence relation.
-The squared Hermite polynomials $\mathop{{}\it He}\nolimits_{n}(\sqrt{t})]^{2}$ turn out to be $\mathbf{S}[\sqrt{1-\nobreak 2D}/\allowbreak(1-\nobreak D),\allowbreak D/(1-\nobreak D)]$ or $\mathbf{R}[1/\sqrt{1-\nobreak z^{2}},\allowbreak z/\allowbreak(1+\nobreak z)]$, but do not satisfy such a relation.
-For a Sheffer sequence $\mathbf{S}[g,f]$, the formal power series $z=f(D)$ may be the Maclaurin expansion of an elementary function of $D$, without the compositional inverse $D=\bar{f}(z)$ appearing in the EGF (30) having the same property.
-An example is the sequence of Abel polynomials $A_{n}(t)=t(t-\nobreak n)^{n-1}$, which turns out to be ${\bf S}[1,D{\rm e}^{D}]$.
-(See [9].) But the compositional inverse of $z=f(D)=D{\rm e}^{D}$ is the Lambert $W$ function $D=\bar{f}(z)=\allowbreak W_{L}(z)=\allowbreak z-\nobreak z^{2}+\nobreak\frac{3}{2}z^{3}+\nobreak\cdots$, which is not an elementary function.
+[…]
+For a Sheffer sequence $\mathbf{S}[g,f]$, the formal power series $z=f(D)$ may be the Maclaurin expansion of an elementary function of $D$, without the compositional inverse $D=\bar{f}(z)$ appearing in the EGF (3.7) having the same property.
+An example is the sequence of Abel polynomials $A_{n}(t)=t(t-\penalty 10000n)^{n-1}$, which turns out to be ${\bf S}[1,D{\rm e}^{D}]$.
+(See [9].) But the compositional inverse of $z=f(D)=D{\rm e}^{D}$ is the Lambert $W$ function $D=\bar{f}(z)=\allowbreak W_{L}(z)=\allowbreak z-\penalty 10000z^{2}+\penalty 10000\frac{3}{2}z^{3}+\penalty 10000\cdots$, which is not an elementary function.
 It is more convenient to refer to the Abel sequence as ${\bf S}[1,D{\rm e}^{D}]$ than as ${\bf R}[1,W_{L}(z)]$.
@@
 For any pair $g,f\in\mathbb{C}[[D]]$ of orders $0,1$ that defines a Sheffer sequence $s_{n}=\mathbf{S}[g,f]_{n}$ of polynomials in $t$, the Heisenberg–Weyl algebra generated by $P,M$ with $[P,M]=1$ has a representation on $\mathbb{C}[t]$, the space of polynomials in $t$, given by lowering and raising operators
-$P=g(D),\qquad M=\left[t-\frac{g^{\prime}(D)}{g(D)}\right]\frac{1}{f^{\prime}(D)},$
+$P=f(D),\qquad M=\left[t-\frac{g^{\prime}(D)}{g(D)}\right]\frac{1}{f^{\prime}(D)},$
 the action of which on the Sheffer polynomials is
-$Ps_{n}=ns_{n-1},\qquad Ms_{n}=s_{n+1}.$ (31)
+$Ps_{n}=ns_{n-1},\qquad Ms_{n}=s_{n+1}.$ (3.8)
 Here, $D$ is interpreted as $D_{t}={\rm d}/{\rm d}t$, and though $P,M$ may be of infinite order in $D$, as differential operators they map $\mathbb{C}[t]$ into itself.
@@
 $t^{0}]\,g(D_{t}){\rm e}^{\lambda f(D_{t})}s_{n}(t)=\lambda^{n},$
-an equality in $\mathbb{C}[t][[\lambda]]$ which like (31) can be used as an alternative definition of the polynomials $s_{n}$ [9].
+an equality in $\mathbb{C}[t][[\lambda]]$ which like (3.8) can be used as an alternative definition of the polynomials $s_{n}$ [9].
 The exponentiated raising operator ${\rm e}^{\lambda M}$ is of interest.
-Each non-negative power $M^{n}$ is an element of $\mathbb{C}[t][[D_{t}]]$, and because $M^{n}s_{0}=s_{n}$ and $s_{0}(t)\equiv 1$, it follows from (30) that
-$\left[{\rm e}^{\lambda M}1\right](t)=\bar{g}(\lambda){\rm e}^{t\bar{f}(\lambda)},$ (32)
+Each non-negative power $M^{n}$ is an element of $\mathbb{C}[t][[D_{t}]]$, and because $M^{n}s_{0}=s_{n}$ and $s_{0}(t)\equiv 1$, it follows from (3.7) that
+$\left[{\rm e}^{\lambda M}1\right](t)=\bar{g}(\lambda){\rm e}^{t\bar{f}(\lambda)},$ (3.9)
 an equality in $\mathbb{C}[t][[\lambda]]$. The right-hand side is the Sheffer EGF at $z=\lambda$.
@@
 They make up a triangular array indexed by $n,k$ with $0\leqslant k\leqslant n$, and satisfy
-$x^{n}=\sum_{k=0}^{n}\left\{{n\atop k}\right\}x^{\underline{k}},$ (33)
-where $x^{\underline{k}}=x(x-1)\dots(x-k+1)$ is the falling factorial with unit “stride.” So, they are coefficients of connection between bases of $\mathbb{C}[x]$.
-The sequence of row polynomials $\sum_{k=0}^{n}\left\{{n\atop k}\right\}t^{k}$, $n\geqslant 0$, which could be called Stirling polynomials (of the second kind), but are more often called Touchard, Bell, or exponential polynomials, is the exponential Riordan sequence ${\bf R}[1,\allowbreak{\rm e}^{z}-\nobreak 1]$ or equivalently the Sheffer sequence ${\bf S}[1,\allowbreak\log(1+\nobreak D)]$.
+$x^{n}=\sum_{k=0}^{n}\left\{{n\atop k}\right\}x^{\underline{k}},$ (3.10)
+[…]
 The generalized Stirling numbers of Hsu–Shiue [34] are a family $\mathop{{}\it HS}\nolimits_{n,k}=\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)$ that subsumes the classical Stirling numbers and many previous generalizations.
 For $A,B,r\in\mathbb{C}$, they are defined by
-$(x+r)^{\underline{n},A}=\sum_{k=0}^{n}HS_{n,k}(A,B,r)\,x^{\underline{k},B},$ (34)
-where $(x+r)^{\underline{n},A}$, $x^{\underline{k},B}$ are falling factorials with stride not necessarily equal to unity, e.g., $x^{\underline{k},B}=x(x-\nobreak B)\dots[x-\nobreak(k-\nobreak 1)B]$.
-The $\mathop{{}\it HS}\nolimits_{n,k}$ reduce to $\left\{{n\atop k}\right\}$ if $(A,B,r)=(0,1,0)$.
-The parametric row polynomials
-$\mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)t^{k},\qquad n\geqslant 0,$ (35)
+$(x+r)^{\underline{n},A}=\sum_{k=0}^{n}HS_{n,k}(A,B,r)\,x^{\underline{k},B},$ (3.11)
+[…]
+$\mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)t^{k},\qquad n\geqslant 0,$ (3.12)
 will be called Hsu–Shiue polynomials. (Note that $\mathop{{}\it HS}\nolimits_{0}(t)\equiv 1$ and $\deg\mathop{{}\it HS}\nolimits_{n}=n$ for all $n$.)
 By examination, this $n$-indexed sequence of polynomials is exponential Riordan for the pair $d_{\textit{HS}}(\cdot;A,r),h_{\textit{HS}}(\cdot;A,B)]$, where if $A,B\neq 0$,
-$d_{\textit{HS}}(z;A,r)\mathrel{:=}(1+Az)^{r/A},$ (36)
+$d_{\textit{HS}}(z;A,r)\mathrel{:=}(1+Az)^{r/A},$ (3.13)
 $h_{\textit{HS}}(z;A,B)\mathrel{:=}\frac{(1+Az)^{B/A}-1}{B},$
@@
 (The cases $A=0$, $B=0$ are handled by taking limits.) One way of confirming this is to verify that eq.
-(34), viewed as an equality between a column vector indexed by $n$ and a matrix–vector product, is consistent with Proposition 3.2.
+(3.11), viewed as an equality between a column vector indexed by $n$ and a matrix–vector product, is consistent with Proposition 3.2.
 It is easily checked that
-$\overline{[d_{\textit{HS}}(\cdot;A,r),h_{\textit{HS}}(\cdot;A,B)]}={[d_{\textit{HS}}(\cdot;B,-r),h_{\textit{HS}}(\cdot;B,A)]},$ (37)
+$\overline{[d_{\textit{HS}}(\cdot;A,r),h_{\textit{HS}}(\cdot;A,B)]}={[d_{\textit{HS}}(\cdot;B,-r),h_{\textit{HS}}(\cdot;B,A)]},$ (3.14)
 which is a statement of duality: as infinite lower-triangular matrices and as elements of the exponential Riordan group $\mathfrak{R}$, Hsu–Shiue arrays $\mathcal{HS}(A,B,r)$ and $\mathcal{HS}(B,A,-r)$ are inverses of each other.
@@
 [34].) So, one can write
-$\begin{aligned} \mathcal{HS}(A,B,r)&={\bf R}[d_{\textit{HS}}(z;A,r),h_{\textit{HS}}(z;A,B)]\\ &={\bf S}[d_{\textit{HS}}(D;B,-r),h_{\textit{HS}}(D;B,A)]\end{aligned}$ (38)
+$\begin{aligned} \mathcal{HS}(A,B,r)&={\bf R}[d_{\textit{HS}}(z;A,r),h_{\textit{HS}}(z;A,B)]\\ &={\bf S}[d_{\textit{HS}}(D;B,-r),h_{\textit{HS}}(D;B,A)]\end{aligned}$ (3.15)
 and
-$\begin{aligned} \mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)&={\bf R}[d_{\textit{HS}}(z;A,r),h_{\textit{HS}}(z;A,B)]_{n}(t)\\ &={\bf S}[d_{\textit{HS}}(D;B,-r),h_{\textit{HS}}(D;B,A)]_{n}(t).\end{aligned}$ (39)
+$\begin{aligned} \mathop{{}\it HS}\nolimits_{n}(t;\,A,B,r)&={\bf R}[d_{\textit{HS}}(z;A,r),h_{\textit{HS}}(z;A,B)]_{n}(t)\\ &={\bf S}[d_{\textit{HS}}(D;B,-r),h_{\textit{HS}}(D;B,A)]_{n}(t).\end{aligned}$ (3.16)
 Regardless of whether one uses the exponential Riordan or the Sheffer notation to express the $\mathop{{}\it HS}\nolimits_{n}$, the following holds.
@@
 If $B\neq 0$, the $\mathop{{}\it HS}\nolimits_{n,k}$ can also be computed from
-$\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)=\frac{1}{B^{k}k!}\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}(Bj+r)^{\underline{n},A},$ (40)
-which comes from eq. (34) by finite-differencing to extract coefficients [40]. And for any $A,B,r$, the $\mathop{{}\it HS}\nolimits_{n,k}$ can be computed from the recurrence (42), below.
+$\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)=\frac{1}{B^{k}k!}\sum_{j=0}^{k}(-1)^{k-j}\binom{k}{j}(Bj+r)^{\underline{n},A},$ (3.17)
+which comes from eq. (3.11) by finite-differencing to extract coefficients [40]. And for any $A,B,r$, the $\mathop{{}\it HS}\nolimits_{n,k}$ can be computed from the recurrence (3.19), below.
 Proposition 3.5.
@@
     The first-order partial differential equation (PDE)
-    $\left[(-Az-1)\frac{\partial}{\partial z}+Bt\,\frac{\partial}{\partial t}+(r+t)\right]\mathcal{HS}=0,$ (41)
+    $\left[(-Az-1)\frac{\partial}{\partial z}+Bt\,\frac{\partial}{\partial t}+(r+t)\right]\mathcal{HS}=0,$ (3.18)
     on a neighborhood of $(t,z)=(0,0)$, with the initial condition $\mathcal{HS}(t,0)\equiv 1$.
@@
     The triangular recurrence
-    $\mathop{{}\it HS}\nolimits_{n+1,k+1}=\left[-An+B(k+1)+r\right]\mathop{{}\it HS}\nolimits_{n,k+1}+\mathop{{}\it HS}\nolimits_{n,k},$ (42)
+    $\mathop{{}\it HS}\nolimits_{n+1,k+1}=\left[-An+B(k+1)+r\right]\mathop{{}\it HS}\nolimits_{n,k+1}+\mathop{{}\it HS}\nolimits_{n,k},$ (3.19)
     with the initial condition that the apex element $\mathop{{}\it HS}\nolimits_{0,0}$ equal unity, and the convention that $\mathop{{}\it HS}\nolimits_{n,k}=0$ if $k<0$ or $k>n$.
@@
 The PDE (with initial condition) can be solved on a neighborhood of $(t,z)=(0,0)$ by the Lagrange–Charpit method of characteristics, yielding the same analytic formula for $\mathcal{HS}(t,z)$.
-And, it is easily checked that $\mathcal{HS}(t,z)$, a function analytic at $(0,0)$ with bivariate Maclaurin expansion in $\mathbb{C}[t][[z]]$, satisfies the PDE if and only if its expansion coefficients $\mathop{{}\it HS}\nolimits_{n,k}$ satisfy the recurrence (42).
+And, it is easily checked that $\mathcal{HS}(t,z)$, a function analytic at $(0,0)$ with bivariate Maclaurin expansion in $\mathbb{C}[t][[z]]$, satisfies the PDE if and only if its expansion coefficients $\mathop{{}\it HS}\nolimits_{n,k}$ satisfy the recurrence (3.19).
 ∎
-For some choices of $A,B,r$, explicit expressions for the polynomial coefficients $\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)$ and the polynomials $\mathop{{}\it HS}\nolimits_{n}(t;A,B,r)$ can be obtained from the summation formula (40) or the recurrence (42).
+For some choices of $A,B,r$, explicit expressions for the polynomial coefficients $\mathop{{}\it HS}\nolimits_{n,k}(A,B,r)$ and the polynomials $\mathop{{}\it HS}\nolimits_{n}(t;A,B,r)$ can be obtained from the summation formula (3.17) or the recurrence (3.19).
 For instance,
-$\mathop{{}\it HS}\nolimits_{n,k}(B,B,r)=\binom{n}{k}r^{\underline{n-k},B},$ (43a)
-$\mathop{{}\it HS}\nolimits_{n,k}(-B,B,r)=\binom{n}{k}(Bk+r)^{\overline{n-k},B},$ (43b)
-with an obvious notation for the rising factorial. Notice that by (43b),
-$\mathop{{}\it HS}\nolimits_{n,k}(-1,1,1)={\binom{n}{k}}\frac{n!}{k!},$ $\mathop{{}\it HS}\nolimits_{n,k}(1,-1,-1)=(-1)^{n-k}{\binom{n}{k}}\frac{n!}{k!},$ (44)
+$\mathop{{}\it HS}\nolimits_{n,k}(B,B,r)=\binom{n}{k}r^{\underline{n-k},B},$ (3.20a)
+$\mathop{{}\it HS}\nolimits_{n,k}(-B,B,r)=\binom{n}{k}(Bk+r)^{\overline{n-k},B},$ (3.20b)
+with an obvious notation for the rising factorial. Notice that by (3.20b),
+$\mathop{{}\it HS}\nolimits_{n,k}(-1,1,1)={\binom{n}{k}}\frac{n!}{k!},$ $\mathop{{}\it HS}\nolimits_{n,k}(1,-1,-1)=(-1)^{n-k}{\binom{n}{k}}\frac{n!}{k!},$ (3.21)
 and
-$\mathop{{}\it HS}\nolimits_{n}(t;\,-1,1,1)=n!\,L_{n}(-t),$ $\mathop{{}\it HS}\nolimits_{n}(t;\,1,-1,-1)=(-1)^{n}n!\,L_{n}(t),$ (45)
+$\mathop{{}\it HS}\nolimits_{n}(t;\,-1,1,1)=n!\,L_{n}(-t),$ $\mathop{{}\it HS}\nolimits_{n}(t;\,1,-1,-1)=(-1)^{n}n!\,L_{n}(t),$ (3.22)
 where $L_{n}$ is the $n$’th Laguerre polynomial.
-By examination, negating the parameters $A,B,r$ of any exponential Riordan array of Hsu–Shiue type multiplies $\mathop{{}\it HS}\nolimits_{n,k}$ by $(-1)^{n-k}$ and alters $\mathop{{}\it HS}\nolimits_{n}(t)$ to $(-1)^{n}\mathop{{}\it HS}\nolimits_{n}(-t)$.
-One can see from (39) that the polynomial sequence $n!\,L_{n}(-t)$ is the Sheffer sequence $\mathbf{S}[1/(1+\nobreak D),\allowbreak D/\allowbreak(1+\nobreak D)]$, or equivalently $\mathbf{R}[1/(1-\nobreak z),z/\allowbreak(1-\nobreak z)]$; and $(-1)^{n}n!\,L_{n}(t)$ is $\mathbf{S}[1/(1-\nobreak D),\allowbreak D/\allowbreak(1-\nobreak D)]$ or $\mathbf{R}[1/(1+\nobreak z),z/\allowbreak(1+\nobreak z)]$.
+By examination, multiplying the parameters $A,B,r$ of any exponential Riordan array $\mathcal{HS}=(\mathop{{}\it HS}\nolimits_{n,k})$ of Hsu–Shiue type by $c\neq 0$ multiplies $\mathop{{}\it HS}\nolimits_{n,k}$ by $c^{n-k}$ and alters $\mathop{{}\it HS}\nolimits_{n}(t)$ to $c^{n}\mathop{{}\it HS}\nolimits_{n}(t/c)$.
+As an application, the $m=n$ case of the $s$-ordering conversion formula (2.13) can be rewritten as
+$\mathopen{:}\,\mathrel{({a^{\dagger}}a)^{n}}\,\mathclose{:}_{s}=\mathopen{:}\,\mathrel{\delta^{n}n!\,L_{n}(-{a^{\dagger}}a/\delta)}\,\mathclose{:}_{s^{\prime}}=\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-\delta,\delta,\delta)}\,\mathclose{:}_{s^{\prime}},$ (3.23)
+where $\delta=(s-s^{\prime})/2$.
+The arrays $\mathcal{HS}(-\delta,\delta,\delta)$, $\delta\in\mathbb{C}$, make up an ‘$s$-shifting’ subgroup of $\mathfrak{R}$ that is isomorphic to the additive group of $\mathbb{C}$.
+One can see from (3.16) that the polynomial sequence $n!\,L_{n}(-t)$ is the Sheffer sequence $\mathbf{S}[1/(1+\penalty 10000D),\allowbreak D/\allowbreak(1+\penalty 10000D)]$, or equivalently $\mathbf{R}[1/(1-\penalty 10000z),z/\allowbreak(1-\penalty 10000z)]$; and $(-1)^{n}n!\,L_{n}(t)$ is $\mathbf{S}[1/(1-\penalty 10000D),\allowbreak D/\allowbreak(1-\penalty 10000D)]$ or $\mathbf{R}[1/(1+\penalty 10000z),z/\allowbreak(1+\penalty 10000z)]$.
 (Cf.
-[9].)
+[9].) The just-mentioned $1$-parameter subgroup of $\mathfrak{R}$ comprises $\mathbf{S}[1/(1+\penalty 10000\delta D],\allowbreak D/(1+\penalty 10000\delta D)]$ or $\allowbreak\mathbf{R}[1/(1-\penalty 10000\delta z),\allowbreak z/(1-\penalty 10000\delta z)]$, $\delta\in\mathbb{C}$.
 An unrelated fact will be needed below: besides $\mathop{{}\it HS}\nolimits_{n,k}(0,1,0)=\left\{{n\atop k}\right\}$, one can prove by induction that $\mathop{{}\it HS}\nolimits_{n,k}(0,1,1)=\left\{{{n+1}\atop{k+1}}\right\}$.
@@
 The normal and anti-normal results, given in Theorem 4.4 below, appeared in the Introduction as eqs.
-(3),(4).
+(1.4),(1.5).
 They will follow from Theorem 4.2, a result of the author on the $n$’th power of ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ [41, Thm.
 6.5], a short proof of which is supplied, to make this paper self-contained.
-As will be sketched, they also follow from the results of Lang [14, 15] and Błasiak et al.
-[16, 17] on exponentiated vector fields and on applications of Sheffer sequences to ordering problems.
-That Fan et al.
-[42] obtained the $R=\nobreak 0$ case of eq.
-(3) by the formal IWOP (integration within ordered products) technique [43] must also be mentioned.
-As motivation for introducing the coefficients of Hsu–Shiue and other Sheffer polynomials, consider the following identities. Firstly, the two in $\mathcal{W}[[\lambda]]$,
-${\rm e}^{\lambda({a^{\dagger}}a)}$ $=\mathopen{:}\,\mathrel{\exp\left[\left({\rm e}^{\lambda}-1\right){a^{\dagger}}a\right]}\,\mathclose{:}_{N}$ (46a)
-$=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda;\,0,1,0)}\,\mathclose{:}_{N}=\mathopen{:}\,\mathrel{\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,0,1,0)\frac{\lambda^{n}}{n!}}\,\mathclose{:}_{N}$ (46b)
+[…]
+To introduce further the coefficients of Hsu–Shiue and other Sheffer polynomials, consider the following identities. Firstly, the two in $\mathcal{W}[[\lambda]]$,
+${\rm e}^{\lambda({a^{\dagger}}a)}$ $=\mathopen{:}\,\mathrel{\exp\left[\left({\rm e}^{\lambda}-1\right){a^{\dagger}}a\right]}\,\mathclose{:}_{N}$ (4.1a)
+$=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda;\,0,1,0)}\,\mathclose{:}_{N}=\mathopen{:}\,\mathrel{\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,0,1,0)\frac{\lambda^{n}}{n!}}\,\mathclose{:}_{N}$ (4.1b)
 and
-${\rm e}^{\lambda(a{a^{\dagger}})}$ $=\mathopen{:}\,\mathrel{{\rm e}^{\lambda}\exp\left[\left({\rm e}^{\lambda}-1\right)a{a^{\dagger}}\right]}\,\mathclose{:}_{N}$ (46a)
-$=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda;\,0,1,1)}\,\mathclose{:}_{N}=\mathopen{:}\,\mathrel{\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,0,1,1)\frac{\lambda^{n}}{n!}}\,\mathclose{:}_{N}.$ (46b)
-Here, (46a) is the Cahill–Glauber normal ordering shown in (1), and (46a) is an immediate consequence.
+${\rm e}^{\lambda(a{a^{\dagger}})}$ $=\mathopen{:}\,\mathrel{{\rm e}^{\lambda}\exp\left[\left({\rm e}^{\lambda}-1\right)a{a^{\dagger}}\right]}\,\mathclose{:}_{N}$ (4.1a)
+$=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda;\,0,1,1)}\,\mathclose{:}_{N}=\mathopen{:}\,\mathrel{\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,0,1,1)\frac{\lambda^{n}}{n!}}\,\mathclose{:}_{N}.$ (4.1b)
+Here, (4.1a) is the Cahill–Glauber normal ordering shown in (1.1), and (4.1a) is an immediate consequence.
 The rewritings in terms of $\mathcal{HS}(t,z)$ come from the $A\to 0$ limit of the EGF formula in Proposition 3.4.
 Secondly, the two in $\mathcal{W}$,
-$({a^{\dagger}}a)^{n}$ $=\sum_{k=0}^{n}\left\{{{n}\atop{k}}\right\}{a^{\dagger}}^{k}a^{k}=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(0,1,0)\,{a^{\dagger}}^{k}a^{k}$ (47)
+$({a^{\dagger}}a)^{n}$ $=\sum_{k=0}^{n}\left\{{{n}\atop{k}}\right\}{a^{\dagger}}^{k}a^{k}=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(0,1,0)\,{a^{\dagger}}^{k}a^{k}$ (4.2)
 and
-$(a{a^{\dagger}})^{n}$ $=\sum_{k=0}^{n}\left\{{{n+1}\atop{k+1}}\right\}{a^{\dagger}}^{k}a^{k}=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(0,1,1)\,{a^{\dagger}}^{k}a^{k}.$ (48)
-These two are due to Katriel [44], who showed for the first time how Stirling numbers arise in boson algebra.
+$(a{a^{\dagger}})^{n}$ $=\sum_{k=0}^{n}\left\{{{n+1}\atop{k+1}}\right\}{a^{\dagger}}^{k}a^{k}=\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(0,1,1)\,{a^{\dagger}}^{k}a^{k}.$ (4.3)
+These are due to Katriel [44], who showed for the first time how Stirling numbers arise in boson algebra.
 Each can also be proved combinatorially, by applying Wick’s theorem and enumerating contractions [45, §9.3].
-Clearly (46b),(46b) follow from (47),(48) by multiplying by $\lambda^{n}/n!$ and summing over $n$.
+More generally, one can expand any polynomial in ${a^{\dagger}}a$ or $a{a^{\dagger}}$ in the monomials ${a^{\dagger}}^{k}a^{k}$ or $a^{k}{a^{\dagger}}^{k}$ [46].
+Clearly (4.1b),(4.1b) follow from (4.2),(4.3) by multiplying by $\lambda^{n}/n!$ and summing over $n$.
 What is needed for a generalization to the exponential operator ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ is a formula for $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}$.
@@
 If $\omega$ is any “single annihilator” word over the alphabet $\{a,{a^{\dagger}}\}$, i.e., $\omega={a^{\dagger}}^{L}a{a^{\dagger}}^{R}$, and $e=L+R-1\geqslant 0$, then for all $n\geqslant 0$,
-$\omega^{n}=\left({a^{\dagger}}^{e}\right)^{n}\left[\sum_{k=0}^{n}\mathbf{R}[d,h]_{n,k}\,{a^{\dagger}}^{k}a^{k}\right],$ (49)
+$\omega^{n}=\left({a^{\dagger}}^{e}\right)^{n}\left[\sum_{k=0}^{n}\mathbf{R}[d,h]_{n,k}\,{a^{\dagger}}^{k}a^{k}\right],$ (4.4)
 an identity in $\mathcal{W}$, where $\mathbf{R}[d,h]$ is some $\omega$-specific exponential Riordan array.
 What is difficult is finding explicitly the $(L,R)$-dependent series $d,h\in\mathbb{C}[[z]]$ that define the Riordan array.
-Identities (46a),(46b) suggest that $\mathbf{R}[d,h]$ should be of Hsu–Shiue type.
-But the derivation of the identity (49) need not be combinatorial, or based on a generalization of Wick’s theorem.
+Identities (4.1a),(4.1b) suggest that $\mathbf{R}[d,h]$ should be of Hsu–Shiue type.
+But the derivation of the identity (4.4) need not be combinatorial, or based on a generalization of Wick’s theorem.
 Combinatorial interpretations of the reordered coefficients of such boson strings as ${a^{\dagger}}^{r_{n}}a^{s_{n}}\dots a^{\dagger}{}^{r_{1}}a^{s_{1}}\in\mathcal{W}$ are known [18].
@@
 (See, e.g., [35, Chapter 4].) The following two parametric identities in $\mathcal{W}$ are extracted from [41, Theorem 6.5], and have a non-combinatorial proof.
-The first subsumes Katriel’s formulae, (47) and (48).
+The first subsumes Katriel’s formulae, (4.2) and (4.3).
 Theorem 4.2.
 For all integer $L,R\geqslant 0$ with $e=L+R-1\geqslant 0$,
-$\begin{aligned} \left({a^{\dagger}}^{L}a{a^{\dagger}}^{R}\right)^{n}&=\left({a^{\dagger}}^{e}\right)^{n}\left[\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(-e,1,R)\,{a^{\dagger}}^{k}a^{k}\right]\\ &=\mathopen{:}\,\mathrel{\left({a^{\dagger}}^{e}\right)^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-e,1,R)}\,\mathclose{:}_{N}\end{aligned}$ (50)
+$\begin{aligned} \left({a^{\dagger}}^{L}a{a^{\dagger}}^{R}\right)^{n}&=\left({a^{\dagger}}^{e}\right)^{n}\left[\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(-e,1,R)\,{a^{\dagger}}^{k}a^{k}\right]\\ &=\mathopen{:}\,\mathrel{\left({a^{\dagger}}^{e}\right)^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-e,1,R)}\,\mathclose{:}_{N}\end{aligned}$ (4.5)
 and
-$\begin{aligned} \left({a^{\dagger}}^{L}a{a^{\dagger}}^{R}\right)^{n}&=\left[\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(e,-1,-L)\,a^{k}{a^{\dagger}}^{k}\right]\left({a^{\dagger}}^{e}\right)^{n}\\ &=\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}(a{a^{\dagger}};\,e,-1,-L)\left({a^{\dagger}}^{e}\right)^{n}}\,\mathclose{:}_{A},\end{aligned}$ (51)
-where the $\mathop{{}\it HS}\nolimits_{n,k}$ coefficients can be computed from the summation formula (40) or the recurrence (42).
+$\begin{aligned} \left({a^{\dagger}}^{L}a{a^{\dagger}}^{R}\right)^{n}&=\left[\sum_{k=0}^{n}\mathop{{}\it HS}\nolimits_{n,k}(e,-1,-L)\,a^{k}{a^{\dagger}}^{k}\right]\left({a^{\dagger}}^{e}\right)^{n}\\ &=\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}(a{a^{\dagger}};\,e,-1,-L)\left({a^{\dagger}}^{e}\right)^{n}}\,\mathclose{:}_{A},\end{aligned}$ (4.6)
+where the $\mathop{{}\it HS}\nolimits_{n,k}$ coefficients can be computed from the summation formula (3.17) or the recurrence (3.19).
 Proof.
-Representing $a,{a^{\dagger}}$ by $D_{x},x$, acting on any convenient space of functions of $x$, yields a faithful representation of $\mathcal{W}$, so it suffices to obtain versions of (50),(51) with $a,{a^{\dagger}}$ replaced by $D_{x},x$.
-Replacing $x$ by $xD_{x}$ in the definition (34) of the coefficients $\mathop{{}\it HS}\nolimits_{n,k}$ yields
-$(xD_{x}+r)^{\underline{n},A}=\sum_{k=0}^{n}HS_{n,k}(A,B,r)\,(xD_{x})^{\underline{k},B}.$ (52)
+Representing $a,{a^{\dagger}}$ by $D_{x},x$, acting on any convenient space of functions of $x$, yields a faithful representation of $\mathcal{W}$, so it suffices to obtain versions of (4.5),(4.6) with $a,{a^{\dagger}}$ replaced by $D_{x},x$.
+Replacing $x$ by $xD_{x}$ in the definition (3.11) of the coefficients $\mathop{{}\it HS}\nolimits_{n,k}$ yields
+$(xD_{x}+r)^{\underline{n},A}=\sum_{k=0}^{n}HS_{n,k}(A,B,r)\,(xD_{x})^{\underline{k},B}.$ (4.7)
 This can be rewritten with the aid of Boole’s lemma $(xD_{x})^{\underline{n}}=x^{n}D_{x}^{n}$, which has a long history [35]. The more general forms of the lemma
-$(xD_{x})^{\underline{n},A}$ $=(x^{A})^{n}(x^{1-A}D_{x})^{n},$ (53a)
-$=(xD_{x}x^{A})^{n}(x^{-\alpha})^{n}$ (53b)
+$(xD_{x})^{\underline{n},A}$ $=(x^{A})^{n}(x^{1-A}D_{x})^{n},$ (4.8a)
+$=(xD_{x}x^{A})^{n}(x^{-A})^{n}$ (4.8b)
 (cf. [48]) can be proved by induction, or operationally: the left and right-hand sides act identically on monomials in $x$. The yet more general forms
-$(xD_{x}+r)^{\underline{n},A}$ $=x^{-r}\left[(x^{A})^{n}(x^{1-A}D_{x})^{n}\right]x^{r}=(x^{A})^{n}(x^{1-A-r}D_{x}x^{r})^{n},$ (54a)
-$=x^{-r}\left[(xD_{x}x^{A})^{n}(x^{-A})^{n}\right]x^{r}=(x^{1-r}D_{x}x^{A+r})^{n}(x^{-A})^{n}$ (54b)
+$(xD_{x}+r)^{\underline{n},A}$ $=x^{-r}\left[(x^{A})^{n}(x^{1-A}D_{x})^{n}\right]x^{r}=(x^{A})^{n}(x^{1-A-r}D_{x}x^{r})^{n},$ (4.9a)
+$=x^{-r}\left[(xD_{x}x^{A})^{n}(x^{-A})^{n}\right]x^{r}=(x^{1-r}D_{x}x^{A+r})^{n}(x^{-A})^{n}$ (4.9b)
 follow by the similarity transformation $(\cdot)\mapsto x^{-r}(\cdot)x^{r}$.
-To obtain a version of (50) with $a,{a^{\dagger}}$ replaced by $D_{x},x$, rewrite the left-hand side of (52) with the aid of (54a), and the right-hand side with the aid of (53a); and also set (A,B,r)=(e,1,R)(A,B,r)=(-e,1,R).
-To obtain (51) similarly, use (54b) and (53b), and set $(A,B,r)=\allowbreak(e,-1,1-\nobreak L)$.
+To obtain a version of (4.5) with $a,{a^{\dagger}}$ replaced by $D_{x},x$, rewrite the left-hand side of (4.7) with the aid of (4.9a), and the right-hand side with the aid of (4.8a); and also set (A,B,r)=(e,1,R)(A,B,r)=(-e,1,R).
+To obtain (4.6) similarly, use (4.9b) and (4.8b), and set $(A,B,r)=\allowbreak(e,-1,1-\penalty 10000L)$.
 ∎
 Example 4.3.
-If $L=R=1$, identities (50),(51) of the theorem reduce to
+If $L=R=1$, identities (4.5),(4.6) of the theorem reduce to
 $\left({a^{\dagger}}a{a^{\dagger}}\right)^{n}=\left\{\begin{aligned} &\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\mathop{{}\it HS}\nolimits_{n}({a^{\dagger}}a;\,-1,1,1)}\,\mathclose{:}_{N}=n!\;\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}L_{n}(-{a^{\dagger}}a)}\,\mathclose{:}_{N},\\ &\mathopen{:}\,\mathrel{\mathop{{}\it HS}\nolimits_{n}(a{a^{\dagger}};\,1,-1,-1)a^{n}}\,\mathclose{:}_{A}=(-1)^{n}n!\;\mathopen{:}\,\mathrel{L_{n}(a{a^{\dagger}}){a^{\dagger}}^{n}}\,\mathclose{:}_{A},\end{aligned}\right$.
-(55)
-by (45), so that
+(4.10)
+by (3.22), so that
 $\left({a^{\dagger}}a{a^{\dagger}}\right)^{n}=\left\{\begin{aligned} &{a^{\dagger}}^{n}\cdot\sum_{k=0}^{n}\frac{n!^{2}}{k!^{2}(n-k)!}\,{a^{\dagger}}^{k}a^{k},\\ &\sum_{k=0}^{n}(-1)^{n-k}\frac{n!^{2}}{k!^{2}(n-k)!}\,a^{k}{a^{\dagger}}^{k}\cdot{a^{\dagger}}^{n}.\end{aligned}\right$.
-(56)
-These are known Laguerre-polynomial representations [48].
+(4.11)
+These are known Laguerre-polynomial representations [49], which by the $(s,s^{\prime})=(\pm 1,\mp 1)$ cases of the conversion formula (2.13) can be seen to be equivalent to $({a^{\dagger}}a{a^{\dagger}})^{n}={a^{\dagger}}^{n}a^{n}{a^{\dagger}}^{n}$, the Viskov identity [35].
 Multiplying the identities in Theorem 4.2 by $\lambda^{n}/n!$ and summing over $n$ yields the following identities in $\mathcal{W}[[\lambda]]$, with the two explicit expressions coming from the EGF formula in Proposition 3.4.
-The first subsumes (46b) and (46b).
+The first subsumes (4.1b) and (4.1b).
 Theorem 4.4.
 For all integer $L,R\geqslant 0$ with $e=L+R-1\geqslant 0$, $G={\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ has the ordered representations
-$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,-e,1,R)}\,\mathclose{:}_{N}\\ &=\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl{(}1-e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}\end{aligned}$ (57)
+$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,-e,1,R)}\,\mathclose{:}_{N}\\ &=\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl(1-e\lambda{a^{\dagger}}^{e}\bigr)^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}\end{aligned}$ (4.12)
 and
-$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}(a{a^{\dagger}},\lambda{a^{\dagger}}^{e};\,e,-1,-L)}\,\mathclose{:}_{A}\\ &=\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl{(}1+e\lambda{a^{\dagger}}^{e}\bigr{)}^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A},\end{aligned}$ (58)
+$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}&=\mathopen{:}\,\mathrel{\mathcal{HS}(a{a^{\dagger}},\lambda{a^{\dagger}}^{e};\,e,-1,-L)}\,\mathclose{:}_{A}\\ &=\mathopen{:}\,\mathrel{\exp\left\{-a{a^{\dagger}}\left[\bigl(1+e\lambda{a^{\dagger}}^{e}\bigr)^{-1/e}-1\right]\right\}(1+e\lambda{a^{\dagger}}^{e})^{-L/e}}\,\mathclose{:}_{A},\end{aligned}$ (4.13)
 with the case $e=0$ handled by taking the formal $e\to 0$ limit.
@@
 To convert either to the other, (1) take the adjoint, (2) replace $a,{a^{\dagger}}$ by ${a^{\dagger}},-a$ (an automorphism of $\mathcal{W}$, entailing the interchange of $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{N}$ and $\mathopen{:}\,\mathrel{\dots}\,\mathclose{:}_{A}$), and (3) negate $\lambda$ and interchange $L,R$.
-The $R=0$ case of (57) was previously derived by Lang [14, 15] and Błasiak et al.
+The $R=0$ case of (4.12) was previously derived by Lang [14, 15] and Błasiak et al.
 [16].
@@
 Also, Penson et al.
-[48] obtained what amounted to the $R=1$ case of (57) by exploiting a Dobiński representation of generalized Bell polynomials.
+[49] obtained what amounted to the $R=1$ case of (4.12) by exploiting a Dobiński representation of generalized Bell polynomials.
 More generally, Błasiak et al.
 [16] integrated $q(x)D_{x}+v(x)$, where $q,v\in\mathbb{C}[[x]]$.
-As ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ equals $({a^{\dagger}})^{L+R}a+\nobreak R({a^{\dagger}})^{L+R-1}$, by substituting $q(x)=x^{L+R}$ and $v(x)=Rx^{L+R-1}$ into their formulae one can derive (57) in full generality.
+As ${a^{\dagger}}^{L}a{a^{\dagger}}^{R}$ equals $({a^{\dagger}})^{L+R}a+\penalty 10000R({a^{\dagger}})^{L+R-1}$, by substituting $q(x)=x^{L+R}$ and $v(x)=Rx^{L+R-1}$ into their formulae one can derive (4.12) in full generality.
 Błasiak et al.
@@
 Up to the taking of an adjoint, they derived a formula for ${\rm e}^{\lambda\tilde{M}^{\dagger}}$: they showed that
-$\begin{split}&\exp\left\{\lambda\,\frac{1}{f^{\prime}({a^{\dagger}})}\left[a-\frac{g^{\prime}({a^{\dagger}})}{g({a^{\dagger}})}\right]\right\}\\ &\quad=\mathopen{:}\,\mathrel{\frac{g({a^{\dagger}})}{g\left(\bar{f}(f({a^{\dagger}})+\lambda)\right)}\exp{\Bigl{\{}\left[\bar{f}(f({a^{\dagger}})+\lambda)-{a^{\dagger}}\right]a\Bigr{\}}}}\,\mathclose{:}_{N},\end{split}$ (59)
+$\begin{split}&\exp\left\{\lambda\,\frac{1}{f^{\prime}({a^{\dagger}})}\left[a-\frac{g^{\prime}({a^{\dagger}})}{g({a^{\dagger}})}\right]\right\}\\ &\quad=\mathopen{:}\,\mathrel{\frac{g({a^{\dagger}})}{g\left(\bar{f}(f({a^{\dagger}})+\lambda)\right)}\exp{\Bigl\{\left[\bar{f}(f({a^{\dagger}})+\lambda)-{a^{\dagger}}\right]a\Bigr\}}}\,\mathclose{:}_{N},\end{split}$ (4.14)
 where $g,f\in\mathbb{C}[[D]]$ are any pair that defines a Sheffer sequence.
-Equation (59) was not obtained in the power-series framework of the present paper, as its derivation employed a Fock space representation and projections onto coherent states, but it is a valid identity in $\mathbb{C}[[{a^{\dagger}}]][a][[\lambda]]$.
-By examination, formally substituting $g(D)=D^{-R}$, $f(D)=-D^{-e}/e$ into (59) yields (57).
+Equation (4.14) was not obtained in the power-series framework of the present paper, as its derivation employed a Fock space representation and projections onto coherent states, but it is a valid identity in $\mathbb{C}[[{a^{\dagger}}]][a][[\lambda]]$.
+By examination, formally substituting $g(D)=D^{-R}$, $f(D)=-D^{-e}/e$ into (4.14) yields (4.12).
 But this additional derivation of eq.
-(57) is not rigorous, as these $g,f$ contain negative powers of $D$ and do not belong to $\mathbb{C}[[D]]$.
+(4.12) is not rigorous, as these $g,f$ contain negative powers of $D$ and do not belong to $\mathbb{C}[[D]]$.
 At most, it is an analytic continuation of a rigorous derivation.
@@
 The main result (Theorem 5.3) can now be approached.
-According to Theorem 4.4, the normally ordered ($s=\nobreak-1$) and anti-normally ordered ($s=\nobreak+1$) representations of the exponential operator ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ can be expressed in terms of $\mathcal{HS}(t,z;\allowbreak A,B,r)$, the EGF of the parametric family of exponential Riordan arrays of Hsu–Shiue type, for which there is an explicit expression.
+According to Theorem 4.4, the normally ordered ($s=\penalty 10000-1$) and anti-normally ordered ($s=\penalty 10000+1$) representations of the exponential operator ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ can be expressed in terms of $\mathcal{HS}(t,z;\allowbreak A,B,r)$, the EGF of the parametric family of exponential Riordan arrays of Hsu–Shiue type, for which there is an explicit expression.
 (Recall Proposition 3.4.) For any $s\in\mathbb{C}$, the $s$-ordered representation of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ will now be obtained in terms of a new family of exponential Riordan arrays, or family of Sheffer polynomial sequences, which is of independent interest.
@@
 Firstly, define parametric functions $g_{T},f_{T}$ of $D$, or more accurately parametric Maclaurin series $g_{T},f_{T}\in\mathbb{C}[[D]]$, by
-$g_{T}(D;B,r,r^{\prime};\,s)\mathrel{:=}\left(\frac{1+s}{2}\right)\left[\frac{1+(\frac{1-s}{2})B\,D}{1-(\frac{1+s}{2})B\,D}\right]^{-r/B}\!\!+\left(\frac{1-s}{2}\right)\left[\frac{1-(\frac{1+s}{2})B\,D}{1+(\frac{1-s}{2})B\,D}\right]^{r^{\prime}/B},$ (60)
+$g_{T}(D;B,r,r^{\prime};\,s)\mathrel{:=}\left(\frac{1+s}{2}\right)\left[\frac{1+(\frac{1-s}{2})B\,D}{1-(\frac{1+s}{2})B\,D}\right]^{-r/B}\!\!+\left(\frac{1-s}{2}\right)\left[\frac{1-(\frac{1+s}{2})B\,D}{1+(\frac{1-s}{2})B\,D}\right]^{r^{\prime}/B},$ (5.1)
 $z=f_{T}(D;A,B;\,s)\mathrel{:=}\frac{[1-(\frac{1+s}{2})B\,D]^{-A/B}-1}{A}+\frac{[1+(\frac{1-s}{2})B\,D]^{-A/B}-1}{-A},$
 with the cases $A=0$, $B=0$ handled by taking limits.
-These $g_{T},f_{T}$ have Maclaurin expansions that are series in $\mathbb{C}[[D]]$ of order 0 and $1$ respectively: $g_{T}=1\cdot D^{0}+\nobreak O(D^{1})$ and $f_{T}=1\cdot D^{1}+\nobreak O(D^{2})$.
+These $g_{T},f_{T}$ have Maclaurin expansions that are series in $\mathbb{C}[[D]]$ of order 0 and $1$ respectively: $g_{T}=1\cdot D^{0}+\penalty 10000O(D^{1})$ and $f_{T}=1\cdot D^{1}+\penalty 10000O(D^{2})$.
 Definition 5.1.
@@
 Here, the pair $\bar{g}_{T},\bar{f}_{T}]$ is the inverse in the exponential Riordan group $\mathfrak{R}$ of the parametric element $g_{T},f_{T}]$ defined in eq.
-(60).
+(5.1).
 That is, $D=\bar{f}_{T}(z;s)$ is inverse to $z=f_{T}(D;s)$ and can be obtained by series reversion, and $\bar{g}_{T}=1/g_{T}(D;s)=1/g_{T}(\bar{f}_{T}(z;s))$.
 The notations $\mathbf{S}[g_{T},f_{T}]$ and $\mathbf{R}[\bar{g}_{T},\bar{f}_{T}]$ are equivalent.
-It should be clear from the definition of $f_{T}$ in (60) that for general $s$ and $A,B$, the inverse $D=\bar{f}_{T}(z)$ of $z=f_{T}(D)$ is not an elementary function. But:
+It should be clear from the definition of $f_{T}$ in (5.1) that for general $s$ and $A,B$, the inverse $D=\bar{f}_{T}(z)$ of $z=f_{T}(D)$ is not an elementary function. But:
 Proposition 5.2.
@@
 Proof.
-Set s=1,+1s=-1,+1 in the formulae (60) for $g_{T},f_{T}$, and compare them with the formulae (36) for $d_{\textit{HS}},h_{\textit{HS}}$.
+Set s=1,+1s=-1,+1 in the formulae (5.1) for $g_{T},f_{T}$, and compare them with the formulae (3.13) for $d_{\textit{HS}},h_{\textit{HS}}$.
 The formula for the parametric Hsu–Shiue EGF $\mathcal{HS}(t,z)$ was supplied in Proposition 3.4.
 ∎
-So, exponential Riordan arrays of the newly defined two-point Hsu–Shiue type (and likewise their row polynomials and EGF’s) interpolate as $s$ varies over the $1$-simplex $1\leqslant s\leqslant+1$} between conventional Hsu–Shiue ones, to which they reduce when $s=\pm 1$.
+So, exponential Riordan arrays of the newly defined two-point Hsu–Shiue type (and likewise their row polynomials and EGF’s) interpolate as the parameter $s$ varies over the $1$-simplex $1\leqslant s\leqslant+1$} between a pair of conventional Hsu–Shiue ones, to which they reduce when $s=\pm 1$.
 One could also define three-point interpolation, with a parameter vector that varies over a $2$-simplex, etc.
 5.2 Main result
-The following theorem states how the $s$-ordered representation of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ can be expressed in terms of the two-point Hsu–Shiue EGF $\mathcal{T}(t,z;A,B,r,r^{\prime};s)$, and subsumes Theorem 4.4, which provided only normal ($s=\nobreak-1$) and anti-normal ($s=\nobreak+1$) orderings.
+The following theorem states how the $s$-ordered representation of ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ can be expressed in terms of the two-point Hsu–Shiue EGF $\mathcal{T}(t,z;A,B,r,r^{\prime};s)$, and subsumes Theorem 4.4, which provided only normal ($s=\penalty 10000-1$) and anti-normal ($s=\penalty 10000+1$) orderings.
 Examples are given in Sec.
@@
 For all integer $L,R\geqslant 0$ with $e=L+R-1\geqslant 0$, $G={\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ has for all $s\in\mathbb{C}$ the $s$-ordered representation
-${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ $=\mathopen{:}\,\mathrel{\mathcal{T}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,e,1,-L,R;\,s)}\,\mathclose{:}_{s}$ (61)
-$=\mathopen{:}\,\mathrel{\bar{g}_{T}(\lambda{a^{\dagger}}^{e};\,s)\exp\left[({a^{\dagger}}a)\bar{f}_{T}(\lambda{a^{\dagger}}^{e};\,s)\right]}\,\mathclose{:}_{s},$ (62)
-where $\bar{g}_{T},\bar{f}_{T}]$ is the Riordan-group inverse of $g_{T},f_{T}]$, defined as in eq. (60), with parameters $(A,B,r,r^{\prime};s)=(e,1,-L,R;s)$. This reduces when $s=\pm 1$ to
+${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ $=\mathopen{:}\,\mathrel{\mathcal{T}({a^{\dagger}}a,\lambda{a^{\dagger}}^{e};\,e,1,-L,R;\,s)}\,\mathclose{:}_{s}$ (5.2)
+$=\mathopen{:}\,\mathrel{\bar{g}_{T}(\lambda{a^{\dagger}}^{e};\,s)\exp\left[({a^{\dagger}}a)\bar{f}_{T}(\lambda{a^{\dagger}}^{e};\,s)\right]}\,\mathclose{:}_{s},$ (5.3)
+where $\bar{g}_{T},\bar{f}_{T}]=[\bar{g}_{T}(z;s),\bar{f}_{T}(z;s)]$ is the Riordan-group inverse of $g_{T},f_{T}]$, defined as in eq.
+(5.1), with parameters $(A,B,r,r^{\prime};s)=(e,1,-L,R;s)$.
+This reduces when $s=\pm 1$ to
 $\mathopen{:}\,\mathrel{(1-e\lambda{a^{\dagger}}^{e})^{-R/e}\exp\left\{\left[\bigl(1-e\lambda{a^{\dagger}}^{e}\bigr)^{-1/e}-1\right]{a^{\dagger}}a\right\}}\,\mathclose{:}_{N}$ $\textrm{if }s=-1$
@@
 Proof.
-The idea is this: $G={\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ equals $\mathopen{:}\,\mathrel{F^{(s)}(\alpha,\alpha^{*})}\,\mathclose{:}_{s}$, where $F^{(s)}(\alpha,\alpha^{*})$ is some function that satisfies the diffusion equation (23), a PDE that evolves $F^{(s)}$ toward smaller values of $s$.
+The idea is this: $G={\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}$ equals $\mathopen{:}\,\mathrel{F^{(s)}(\alpha,\alpha^{*})}\,\mathclose{:}_{s}$, where $F^{(s)}(\alpha,\alpha^{*})$ is some function that satisfies the diffusion equation (2.19), a PDE that evolves $F^{(s)}$ toward smaller values of $s$.
 (In this proof, for clarity the commuting arguments of $F^{(s)}$ will be written as $\alpha,\alpha^{*}$ and not as $a,{a^{\dagger}}$.
-The notational convention in quantum optics that dependence on $\alpha,\alpha^{*}$ really means dependence on a single complex variable $\alpha=\allowbreak\alpha_{r}+\nobreak{\rm i}\alpha_{i}$ will briefly be followed.)
-In principle, one could start with $F^{(+1)}=F^{(A)}$ and integrate the diffusion equation backward in $s$ toward $F^{(-1)}=F^{(N)}$, as one does in quantum optics when smoothing the Glauber–Sudarshan $P$-function ($s=\nobreak+1$) to obtain the Husimi $Q$-function ($s=\nobreak-1$) [2, 3].
+The notational convention in quantum optics that dependence on $\alpha,\alpha^{*}$ really means dependence on a single complex variable $\alpha=\allowbreak\alpha_{r}+\penalty 10000{\rm i}\alpha_{i}$ will briefly be followed.)
+In principle, one could start with $F^{(+1)}=F^{(A)}$ and integrate the diffusion equation backward in $s$ toward $F^{(-1)}=F^{(N)}$, as one does in quantum optics when smoothing the Glauber–Sudarshan $P$-function ($s=\penalty 10000+1$) to obtain the Husimi $Q$-function ($s=\penalty 10000-1$) [2, 3].
 The differential operator $\partial^{2}/\partial\alpha\partial\alpha^{*}$ in the PDE is to be interpreted as $\Delta/4$, with $\Delta=\partial^{2}/\partial\alpha_{r}^{2}+\partial^{2}/\partial\alpha_{i}^{2}$ the Laplacian on the complex $\alpha$-plane.
 However, to avoid convergence questions the exponential operator $G$ is viewed in this paper as an element of $\mathcal{W}[[\lambda]]$, and $F^{(s)}$ is accordingly an element of $\mathcal{W}_{c}[[\lambda]]$.
-So $F^{(s)}$ is not a single function on the complex $\alpha$-plane (i.e., a function of $\alpha_{r},\alpha_{i}$ or informally of $\alpha,\alpha^{*}$), and the convolution or smoothing formula (24) cannot easily be used.
+So $F^{(s)}$ is not a single function on the complex $\alpha$-plane (i.e., a function of $\alpha_{r},\alpha_{i}$ or informally of $\alpha,\alpha^{*}$), and the convolution or smoothing formula (2.20) cannot easily be used.
 Instead, $F^{(s)}$ at each value of $s$ is a power series in $\lambda$, with coefficients that are polynomial in $\alpha,\alpha^{*}$, and the series is formal: it is never summed.
@@
 The statement of the theorem incorporates the $s=\pm 1$ expressions for $F^{(s)}$ from Theorem 4.4. To match their functional form, make the Ansatz
-$F^{(s)}(\alpha,\alpha^{*})={V}(\alpha^{*}\alpha,\lambda{\alpha^{*}}^{e};\,s),$ (63)
+$F^{(s)}(\alpha,\alpha^{*})={V}(\alpha^{*}\alpha,\lambda{\alpha^{*}}^{e};\,s),$ (5.4)
 where $\lambda\neq 0$ and the function $V(t,z;s)$ is to be determined.
-Changing the independent variables $\alpha,\alpha^{*}$ (really $\alpha_{r},\alpha_{i}$) to $t=\alpha^{*}\alpha$, $z=\lambda{\alpha^{*}}^{e}$ converts the diffusion equation (23) to an equation of diffusion type on the $(t,z)$-plane,
-$\left[\frac{\partial}{\partial s}+\mathcal{X}\right]V(t,z;\,s)=0,\qquad\mathcal{X}\mathrel{:=}\frac{1}{2}\left(t\frac{\partial^{2}}{\partial t^{2}}+\frac{\partial}{\partial t}+ez\,\frac{\partial^{2}}{\partial t\,\partial z}\right),$ (64)
+Changing the independent variables $\alpha,\alpha^{*}$ (really $\alpha_{r},\alpha_{i}$) to $t=\alpha^{*}\alpha$, $z=\lambda{\alpha^{*}}^{e}$ converts the diffusion equation (2.19) to an equation of diffusion type on the $(t,z)$-plane,
+$\left[\frac{\partial}{\partial s}+\mathcal{X}\right]V(t,z;\,s)=0,\qquad\mathcal{X}\mathrel{:=}\frac{1}{2}\left(t\frac{\partial^{2}}{\partial t^{2}}+\frac{\partial}{\partial t}+ez\,\frac{\partial^{2}}{\partial t\,\partial z}\right),$ (5.5)
 irrespective of $\lambda$; so that formally,
-$V(\cdot,\cdot;\,s)=\exp\left[(s_{0}-s)\mathcal{X}\right]\,V(\cdot,\cdot;\,s_{0}),\qquad s<s_{0}.$ (65)
+$V(\cdot,\cdot;\,s)=\exp\left[(s_{0}-s)\mathcal{X}\right]\,V(\cdot,\cdot;\,s_{0}),\qquad s<s_{0}.$ (5.6)
 There is an obvious $z$-independent solution, namely
-$V(t,z;\,s)\varpropto(s_{0}-s)^{-1}{\rm e}^{-2t/(s_{0}-s)},\qquad s<s_{0},$ (66)
+$V(t,z;\,s)\varpropto(s_{0}-s)^{-1}{\rm e}^{-2t/(s_{0}-s)},\qquad s<s_{0},$ (5.7)
 for any real $s_{0}$.
-Depending on $t=\alpha^{*}\alpha$ but not on $z=\lambda{\alpha^{*}}^{e}$, for $s_{0}=+1$ this is in fact the $s$-ordered quasiprobability density on the complex $\alpha$-plane coming from the density operator of the vacuum state [2, §12.4.4], showing how the $P$-function ($s=\nobreak+1$), which is a $\delta$-function located at $\alpha=0$, broadens into a Gaussian as $s$ is decreased and the $Q$-function ($s=\nobreak-1$) is approached.
-But it is qualitatively different from the expressions $V(t=\alpha^{*}\alpha,\allowbreak z=\lambda{\alpha^{*}}^{e};\pm 1)$ appearing in Theorem 4.4.
+Depending on $t=\alpha^{*}\alpha$ but not on $z=\lambda{\alpha^{*}}^{e}$, for $s_{0}=+1$ this is in fact the $s$-ordered quasiprobability density on the complex $\alpha$-plane coming from the density operator of the vacuum state [2, §12.4.4], showing how the $P$-function ($s=\penalty 10000+1$), which is a $\delta$-function located at $\alpha=0$, broadens into a Gaussian as $s$ is decreased and the $Q$-function ($s=\penalty 10000-1$) is approached.
+But it is qualitatively different from the expressions $V(t=\alpha^{*}\alpha,\allowbreak z=\lambda{\alpha^{*}}^{e};\allowbreak s=\pm 1)$ appearing in Theorem 4.4.
 To incorporate those functions, make the further Riordan/Sheffer Ansatz
-$V(t,z;\,s)=\mathbf{R}[\bar{g}(\cdot;s),\bar{f}(\cdot;s)](t,z)=\mathbf{S}[g(\cdot;s),f(\cdot;s)](t,z),$ (67)
+$V(t,z;\,s)=\mathbf{R}[\bar{g}(\cdot;s),\bar{f}(\cdot;s)](t,z)=\mathbf{S}[g(\cdot;s),f(\cdot;s)](t,z),$ (5.8)
 which says that
-$V(t,z;\,s)=\bar{g}(z;\,s)\,{\rm e}^{t\bar{f}(z;\,s)},$ (68)
+$V(t,z;\,s)=\bar{g}(z;\,s)\,{\rm e}^{t\bar{f}(z;\,s)},$ (5.9)
 where as usual, $\bar{g}(z;s)=1\cdot z^{0}+O(z^{1})$ and $D=\bar{f}(z;s)=1\cdot z^{1}+O(z^{2})$, with $g(D;s)=1\cdot D^{0}+O(D^{1})$ and $z=f(D;s)=1\cdot D^{1}+O(D^{2})$.
-Substituting (68) into (64) and separating the terms proportional to $t^{0}$ and $t^{1}$ gives
-$2\bar{g}_{s}+ez\bar{f}\bar{g}_{z}+(\bar{f}+ez\bar{f}_{z})g=0,$ (69a)
-$2\bar{f}_{s}+ez\bar{f}\bar{f}_{z}+\bar{f}^{2}=0,$ (69b)
+Substituting (5.9) into (5.5) and separating the terms proportional to $t^{0}$ and $t^{1}$ gives
+$2\bar{g}_{s}+ez\bar{f}\bar{g}_{z}+(\bar{f}+ez\bar{f}_{z})g=0,$ (5.10a)
+$2\bar{f}_{s}+ez\bar{f}\bar{f}_{z}+\bar{f}^{2}=0,$ (5.10b)
 where subscripts denote partial derivatives. These are coupled nonlinear PDEs for $\bar{g}(z;s)$, $\bar{f}(z;s)$. The two-point boundary conditions
-$\bar{g}(z;\,s=-1)=(1-ez)^{-R/e},$ $\bar{g}(z;\,s=+1)=(1+ez)^{-L/e},$ (70a)
-$\bar{f}(z;\,s=-1)=(1-ez)^{-1/e}-1,$ $\bar{f}(z;\,s=+1)=1-(1+ez)^{-1/e}$ (70b)
+$\bar{g}(z;\,s=-1)=(1-ez)^{-R/e},$ $\bar{g}(z;\,s=+1)=(1+ez)^{-L/e},$ (5.11a)
+$\bar{f}(z;\,s=-1)=(1-ez)^{-1/e}-1,$ $\bar{f}(z;\,s=+1)=1-(1+ez)^{-1/e}$ (5.11b)
 come from Theorem 4.4.
 The function (or more accurately formal infinite series) $z=f(D;s)$ is the inverse of $D=\bar{f}(z;s)$, computed by series reversion.
-With its aid, the independent variables $z,s$ in (69) can be replaced by $D,s$, and the dependent variables $\bar{g},\bar{f}$ by $g,f$, where $g=1/(\bar{g}\circ f)$, i.e., $\bar{g}=1/(g\circ\bar{f})$.
+With its aid, the independent variables $z,s$ in (5.10) can be replaced by $D,s$, and the dependent variables $\bar{g},\bar{f}$ by $g,f$, where $g=1/(\bar{g}\circ f)$, i.e., $\bar{g}=1/(g\circ\bar{f})$.
 This yields the PDEs
-$-2f_{D}\,g_{s}+D^{2}f_{D}\,g_{D}+(Df_{D}+ef)g=0,$ (71a)
-$-2f_{s}+D^{2}f_{D}+eDf=0,$ (71b)
-which are satisfied by $f(D;s)$ and $g(D;s)$. The two-point boundary conditions on these dual functions follow from (70a),(70b) and are
-$g(D;\,s=-1)=(1+D)^{-R},$ $g(D;\,s=+1)=(1-D)^{-L},$ (72a)
-$f(D;\,s=-1)=\left[1-(1+D)^{-e}\right]/e,$ $f(D;\,s=+1)=\left[(1-D)^{-e}-1\right]/e.$ (72b)
-The “dualized” PDEs (71) are linear and of first order, and can be solved by Lagrange’s method of characteristics.
-The solution of (71b) is
-$z=f(D;\,s)=D^{-e}F(\eta),$ (73)
-a similarity solution in which $\eta\mathrel{:=}s-2/D$ is the similarity variable and the function $F(\eta)$ is arbitrary. To match the boundary conditions (72b), choose
-$F(\eta)=\frac{\left(\frac{-1-\eta}{2}\right)^{-e}-\left(\frac{1-\eta}{2}\right)^{-e}}{e},$ (74)
+$-2f_{D}\,g_{s}+D^{2}f_{D}\,g_{D}+(Df_{D}+ef)g=0,$ (5.12a)
+$-2f_{s}+D^{2}f_{D}+eDf=0,$ (5.12b)
+which are satisfied by $f(D;s)$ and $g(D;s)$. The two-point boundary conditions on these dual functions follow from (5.11a),(5.11b) and are
+$g(D;\,s=-1)=(1+D)^{-R},$ $g(D;\,s=+1)=(1-D)^{-L},$ (5.13a)
+$f(D;\,s=-1)=\left[1-(1+D)^{-e}\right]/e,$ $f(D;\,s=+1)=\left[(1-D)^{-e}-1\right]/e.$ (5.13b)
+The “dualized” PDEs (5.12) are linear and of first order, and can be solved by Lagrange’s method of characteristics.
+The solution of (5.12b) is
+$z=f(D;\,s)=D^{-e}F(\eta),$ (5.14)
+a similarity solution in which $\eta\mathrel{:=}s-2/D$ is the similarity variable and the function $F(\eta)$ is arbitrary. To match the boundary conditions (5.13b), choose
+$F(\eta)=\frac{\left(\frac{-1-\eta}{2}\right)^{-e}-\left(\frac{1-\eta}{2}\right)^{-e}}{e},$ (5.15)
 so that
-$z=f(D;\,s)=\frac{\left[1-(\frac{1+s}{2})D\right]^{-e}-{\left[1+(\frac{1-s}{2})D\right]^{-e}}}{e},$ (75)
-with expansion $f(D;s)=1\cdot D^{1}+O(D^{2})$. With this choice of f(D;s)f(D;\,s), the PDE (71a) is solved in the same way. It has the similarity solution
-$g(D;\,s)=\left[\left(\frac{1+s}{2}\right)+\left(\frac{1-s}{2}\right)\left(\frac{\eta+1}{\eta-1}\right)^{e+1}\right]G(\eta),$ (76)
-where the function $G(\eta)$ is arbitrary. To match the boundary conditions (72a), choose
-$G(\eta)=\left(\frac{\eta-1}{\eta+1}\right)^{L},$ (77)
+$z=f(D;\,s)=\frac{\left[1-(\frac{1+s}{2})D\right]^{-e}-{\left[1+(\frac{1-s}{2})D\right]^{-e}}}{e},$ (5.16)
+with expansion $f(D;s)=1\cdot D^{1}+O(D^{2})$. With this choice of f(D;s)f(D;\,s), the PDE (5.12a) is solved in the same way. It has the similarity solution
+$g(D;\,s)=\left[\left(\frac{1+s}{2}\right)+\left(\frac{1-s}{2}\right)\left(\frac{\eta+1}{\eta-1}\right)^{e+1}\right]G(\eta),$ (5.17)
+where the function $G(\eta)$ is arbitrary. To match the boundary conditions (5.13a), choose
+$G(\eta)=\left(\frac{\eta-1}{\eta+1}\right)^{L},$ (5.18)
 so that
-$g(D;\,s)=\left(\frac{1+s}{2}\right)\left[\frac{1+(\frac{1-s}{2})D}{1-(\frac{1+s}{2})D}\right]^{L}+\left(\frac{1-s}{2}\right)\left[\frac{1-(\frac{1+s}{2})D}{1+(\frac{1-s}{2})D}\right]^{R},$ (78)
-with $g(D;s)=1\cdot D^{0}+O(D^{1})$. Comparing (75),(78) with the definitions of $f_{T},g_{T}$ given in (60) reveals that
-$f(D;\,s)=f_{T}(D;e,1;\,s),\qquad g(D;\,s)=g_{T}(D;1,-L,R;\,s).$ (79)
+$g(D;\,s)=\left(\frac{1+s}{2}\right)\left[\frac{1+(\frac{1-s}{2})D}{1-(\frac{1+s}{2})D}\right]^{L}+\left(\frac{1-s}{2}\right)\left[\frac{1-(\frac{1+s}{2})D}{1+(\frac{1-s}{2})D}\right]^{R},$ (5.19)
+with $g(D;s)=1\cdot D^{0}+O(D^{1})$. Comparing (5.16),(5.19) with the definitions of $f_{T},g_{T}$ given in (5.1) reveals that
+$f(D;\,s)=f_{T}(D;e,1;\,s),\qquad g(D;\,s)=g_{T}(D;1,-L,R;\,s).$ (5.20)
 Therefore
-$V(t,z;\,s)=\mathcal{T}(t,z;\,e,1,-L,R;\,s),$ (80)
+$V(t,z;\,s)=\mathcal{T}(t,z;\,e,1,-L,R;\,s),$ (5.21)
 which is the substance of the theorem. ∎
 Remark.
-Wilcox [49] solves a PDE similar to (64) by normal ordering an exponentiated second-order differential operator analogous to the $\exp\left[(s_{0}-\nobreak s)\mathcal{X}\right]$ in eq.
-(65).
-In his paper the PDE comes first, but in the just-concluded proof of Theorem 5.3 the normal ordering (actually $s$-ordering) came first, and determined the PDE to be solved.
+Wilcox [50] solved a PDE similar to (5.5) by normal ordering an exponentiated second-order differential operator analogous to the $\exp\left[(s_{0}-\penalty 10000s)\mathcal{X}\right]$ in eq.
+(5.6).
+In his paper the PDE came first, but in the just-concluded proof of Theorem 5.3 the normal ordering (actually $s$-ordering) came first, and determined the PDE to be solved.
 Corollary 5.4.
 For all integer $L,R\geqslant 0$ with $e=L+R-1\geqslant 0$,
-$({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}=\mathopen{:}\,\mathrel{T_{n}({a^{\dagger}}a;e,1,-L,R;s)}\,\mathclose{:}_{s},$
+$({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}=\mathopen{:}\,\mathrel{\left({a^{\dagger}}^{e}\right)^{n}T_{n}({a^{\dagger}}a;e,1,-L,R;s)}\,\mathclose{:}_{s},$
 where $T_{n}(t;e,1,-L,R;s)$ is the case $(A,B,r,r^{\prime})=(e,1,-L,R)$ of the nn\!’th parametric Sheffer polynomial $\mathbf{S}[g_{T},f_{T}]_{n}(t)=\mathbf{R}[\bar{g}_{T},\bar{f}_{T}]_{n}(t)$, with $g_{T}(D;s)$ and $z=f_{T}(D;s)$ given in Definition 5.1, i.e., in eq.
-(60).
+(5.1).
 Proof.
-Expand each side of the identity (61) in Theorem 5.3 as a formal power series in $\lambda$, and equate the coefficients of $\lambda^{n}$ on the two sides. ∎
-The following is the Weyl ordering ($s=\nobreak 0$) case of the main result Theorem 5.3, which is the one that can be expressed most simply.
-The functions $\bar{g}(z)$, $D=\bar{f}(z)$ come from $g(D)$, $z=f(D)$, which in turn come from the two-point functions $g_{T}(D;s)$, $z=f_{T}(D;s)$ defined in (60), by setting $s=\nobreak 0$.
+Expand each side of the identity (5.2) in Theorem 5.3 as a formal power series in $\lambda$, and equate the coefficients of $\lambda^{n}$ on the two sides. ∎
+The s=1s=-1 (normal ordering) case of the formula in this corollary is an illustration of Prop. 4.1, the result of Duchamp et al.
+[…]
 Theorem 5.5.
@@
 6 Examples
-The main result (Theorem 5.3) will now be applied, to produce several new $s$-ordered representations of exponential boson operators ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}\in\mathcal{W}[[\lambda]]$, and of $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}\in\nobreak\mathcal{W}$.
-According to the theorem, for any $(L,R)$ one needs the EGF of a certain family of two-point Hsu–Shiue polynomials (of Sheffer type), denoted by $\mathcal{T}(t,z;\allowbreak A,B,r,r^{\prime};s)$.
-The EGF equals $\bar{g}_{T}(z){\rm e}^{t\bar{f}_{T}(z)}$, where $\bar{g}_{T},\bar{f}_{T}]$ is the inverse in the exponential Riordan group $\mathfrak{R}$ of the parametric $s$-dependent element $g_{T},f_{T}]$ defined in eq.
-(60).
-One simply substitutes $(A,B,r,r^{\prime})=(e,1,-L,R)$, where $e\mathrel{:=}\allowbreak L+\nobreak R-\nobreak 1$, and sets $(t,z)=({a^{\dagger}}a,\lambda{a^{\dagger}}^{e})$ (which really means $(t,z)=(\alpha^{*}\alpha,\lambda{\alpha^{*}}^{e})$), to obtain the desired $s$-ordered representation.
+The main result (Theorem 5.3) will now be applied, to produce several new $s$-ordered representations of exponential boson operators ${\rm e}^{\lambda({a^{\dagger}}^{L}a{a^{\dagger}}^{R})}\in\mathcal{W}[[\lambda]]$, and of $({a^{\dagger}}^{L}a{a^{\dagger}}^{R})^{n}\in\penalty 10000\mathcal{W}$.
+[…]
 The only possible stumbling block is the computation of the inverse group element (the function $D=\bar{f}_{T}(z;s)$ is the compositional inverse of $z=f_{T}(D;s)$).
@@
 This includes the operators ${\rm e}^{\lambda({a^{\dagger}}a)}$, ${\rm e}^{\lambda(a{a^{\dagger}})}$, with $(L,R)=(1,0)$, $(0,1)$.
-Substituting $(A,B,r,r^{\prime})=\allowbreak(e,1,-L,R)=\allowbreak(0,1,-1,0)$ into (60) gives
-$g_{T}(D;\,s)=\frac{2}{2-D-sD},\qquad z=f_{T}(D;\,s)=\ln\left(\frac{2+D-sD}{2-D-sD}\right).$ (81)
+Substituting $(A,B,r,r^{\prime})=\allowbreak(e,1,-L,R)=\allowbreak(0,1,-1,0)$ into (5.1) gives
+$g_{T}(D;\,s)=\frac{2}{2-D-sD},\qquad z=f_{T}(D;\,s)=\ln\left(\frac{2+D-sD}{2-D-sD}\right).$ (6.1)
 (In deriving the latter, substituting $A=0$ means taking the $A\to 0$ limit.) The inverse $D=\bar{f}_{T}(z;s)$ and $\bar{g}_{T}(z;s)=1/g_{T}(D=\bar{f}_{T}(z;s);s)$ become
-$\bar{g}_{T}(z;\,s)=\frac{2}{1+{\rm e}^{z}-s(1-{\rm e}^{z})},\qquad D=\bar{f}_{T}(z;\,s)=\frac{2({\rm e}^{z}-1)}{1+{\rm e}^{z}-s(1-{\rm e}^{z})}.$ (82)
+$\bar{g}_{T}(z;\,s)=\frac{2}{1+{\rm e}^{z}-s(1-{\rm e}^{z})},\qquad D=\bar{f}_{T}(z;\,s)=\frac{2({\rm e}^{z}-1)}{1+{\rm e}^{z}-s(1-{\rm e}^{z})}.$ (6.2)
 Substituting into $\bar{g}_{T}(z;s){\rm e}^{t\bar{f}_{T}(z;s)}$ and setting $(t,z)=({a^{\dagger}}a,\lambda{a^{\dagger}}^{e})=\allowbreak({a^{\dagger}}a,\lambda)$ yields
-${\rm e}^{\lambda({a^{\dagger}}a)}=\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}{a^{\dagger}}a\right]}\,\mathclose{:}_{s},$ (83)
-which is precisely the original Cahill–Glauber $s$-ordering result, eq. (2). The case $(L,R)=(0,1)$ is similar.
+${\rm e}^{\lambda({a^{\dagger}}a)}=\mathopen{:}\,\mathrel{\frac{2}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}\exp\left[\frac{2({\rm e}^{\lambda}-1)}{1+{\rm e}^{\lambda}-s(1-{\rm e}^{\lambda})}{a^{\dagger}}a\right]}\,\mathclose{:}_{s},$ (6.3)
+which is precisely the original Cahill–Glauber $s$-ordering result, eq.
+(1.2).
+The functions $g_{T},f_{T}$ given in (6.1) appear in its inverse, eq.
+(1.3).
+The case $(L,R)=(0,1)$ is similar to the case $(L,R)=(1,0)$.
 6.2 The case $e=1$
 This includes the operators ${\rm e}^{\lambda({a^{\dagger}}^{2}a)}$, ${\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}})}$, ${\rm e}^{\lambda(a{a^{\dagger}}^{2})}$, with $(L,R)=(2,0)$, $(1,1)$, $(0,2)$.
-Substituting $(A,B)=(e,1)=(1,1)$ into (60) gives
+Substituting $(A,B)=(e,1)=(1,1)$ into (5.1) gives
 $g_{T}(D;\,s)=\left\{\begin{aligned} &\frac{4+D^{2}-s^{2}D^{2}}{(2-D-sD)^{2}},&\qquad&(L,R)=(2,0);\\ &\frac{4+D^{2}-s^{2}D^{2}}{(2-sD)^{2}-D^{2}},&\qquad&(L,R)=(1,1);\\ &\frac{4+D^{2}-s^{2}D^{2}}{(2+D-sD)^{2}},&\qquad&(L,R)=(0,2);\end{aligned}\right$.
-(84)
+(6.4)
 and
-$z=f_{T}(D;\,s)=\frac{4D}{4-D^{2}}.$ (85)
+$z=f_{T}(D;\,s)=\frac{4D}{4-D^{2}}.$ (6.5)
 Hence,
 $\bar{g}_{T}(z;\,s)=\left\{\begin{aligned} &\frac{1+s}{1+2sz+z^{2}+(z+s)\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(2,0);\\ &\frac{1}{\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(1,1);\\ &\frac{1-s}{1+2sz+z^{2}-(z+s)\sqrt{1+2sz+z^{2}}},&\qquad&(L,R)=(0,2);\end{aligned}\right$.
-(86)
+(6.6)
 and
-$D=\bar{f}_{T}(z;\,s)=\frac{2z}{1+sz+\sqrt{1+2sz+z^{2}}}.$ (87)
+$D=\bar{f}_{T}(z;\,s)=\frac{2z}{1+sz+\sqrt{1+2sz+z^{2}}}.$ (6.7)
 Substituting into $\bar{g}_{T}(z;s){\rm e}^{t\bar{f}_{T}(z;s)}$, with $(t,z)=({a^{\dagger}}a,\lambda{a^{\dagger}}^{e})=({a^{\dagger}}a,\lambda{a^{\dagger}})$, yields
@@
 One has the $s$-ordered representations
-$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{2}a)}&=\mathopen{:}\,\frac{1+s}{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}+(\lambda{a^{\dagger}}+s)\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\\ {\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}})}&=\mathopen{:}\,\frac{1}{\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\\ {\rm e}^{\lambda(a{a^{\dagger}}^{2})}&=\mathopen{:}\,\frac{1-s}{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}-(\lambda{a^{\dagger}}+s)\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\end{aligned}$ (88)
+$\begin{aligned} {\rm e}^{\lambda({a^{\dagger}}^{2}a)}&=\mathopen{:}\,\frac{1+s}{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}+(\lambda{a^{\dagger}}+s)\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\\ {\rm e}^{\lambda({a^{\dagger}}a{a^{\dagger}})}&=\mathopen{:}\,\frac{1}{\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\\ {\rm e}^{\lambda(a{a^{\dagger}}^{2})}&=\mathopen{:}\,\frac{1-s}{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}-(\lambda{a^{\dagger}}+s)\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\times\mathcal{E}\,\mathclose{:}_{s},\end{aligned}$ (6.8)
 where
-$\begin{aligned} \mathcal{E}\mathrel{:=}\exp\left[({a^{\dagger}}a)\,\frac{2\lambda{a^{\dagger}}}{1+s\lambda{a^{\dagger}}+\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\right],\end{aligned}$ (89)
+$\begin{aligned} \mathcal{E}\mathrel{:=}\exp\left[({a^{\dagger}}a)\,\frac{2\lambda{a^{\dagger}}}{1+s\lambda{a^{\dagger}}+\sqrt{1+2s\lambda{a^{\dagger}}+\lambda^{2}{a^{\dagger}}^{2}}}\right],\end{aligned}$ (6.9)
 for all $s\in\mathbb{C}$.
-When $s=\nobreak-1$ and $s=\nobreak+1$, each of these three ordered representations reduces to a normal ordered and an anti-normal ordered one, which are the ones provided by Theorem 4.4.
-(To set $s=\nobreak-1$ in the first or $s=\nobreak+1$ in the third, a limit must be taken.) But, the three Weyl-ordered representations obtained by setting $s=\nobreak 0$ seem to be new to the literature.
-From such identities as those of Theorem 6.1, it is possible to derive explicitly $s$-ordered series expansions of $({a^{\dagger}}^{2}a)^{n}$, $({a^{\dagger}}a{a^{\dagger}})^{n}$, $(a{a^{\dagger}}^{2})^{n}$; at least, when s=1,0,+1s=-1,0,+1.
-For the case of $({a^{\dagger}}a{a^{\dagger}})^{n}$ when $s=\pm 1$, this has already been done: see the Laguerre-polynomial representations in eqs.
-(55),(56) above, each term in which is explicitly normal, resp.
-anti-normal ordered.
-The expansion of $({a^{\dagger}}a{a^{\dagger}})^{n}$ when $s=\nobreak 0$ is of particular interest: it expresses $({a^{\dagger}}a{a^{\dagger}})^{n}$ as the symmetric (Weyl) quantization of a classical function on the complex $\alpha$-plane.
+[…]
+[…]
 Proposition 6.2.
-One has the Weyl-ordered ($s=\nobreak 0$) representation
+[…]
 $({a^{\dagger}}a{a^{\dagger}})^{n}=\sum_{{0\leqslant k\leqslant n}\atop{\text{$n-k$ even}}}(-1)^{(n-k)/2}\,2^{-(n-k)}\binom{n}{(n-k)/2}\,\frac{n!}{k!}\,\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}({a^{\dagger}}a)^{k}}\,\mathclose{:}_{W}.$
 Remark.
-The coefficients in this expansion, unlike those in (55),(56), are rational rather than integer-valued.
+The coefficients in this expansion, unlike those in (4.10),(4.11), are rational rather than integer-valued.
 Proof.
-Consider the $(L,R)=(1,1)$ case of Corollary 5.4 when $s=\nobreak 0$, or more concretely the $(L,R)=(1,1)$ case of Theorem 5.5, or the second identity in Theorem 6.1 when $s=\nobreak 0$.
-They all imply that
-$({a^{\dagger}}a{a^{\dagger}})^{n}=\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\,\mathbf{R}[\bar{g}_{T}(z;\,0),\bar{f}_{T}(z;\,0)]_{n}(t={a^{\dagger}}a)}\,\mathclose{:}_{W},$ (90)
-where $\bar{g}_{T}(z;0)$ comes from the $(L,R)=(1,1)$ case of (86), and $\bar{f}_{T}(z;0)$ comes similarly from (87). That is,
-$({a^{\dagger}}a{a^{\dagger}})^{n}=\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\,\mathbf{R}\left[\frac{1}{\sqrt{1+z^{2}}},\frac{2z}{1+\sqrt{1+z^{2}}}\right]_{n}\!(t={a^{\dagger}}a)}\,\mathclose{:}_{W}.$ (91)
+[…]
+$({a^{\dagger}}a{a^{\dagger}})^{n}=\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\,\mathbf{R}[\bar{g}_{T}(z;\,0),\bar{f}_{T}(z;\,0)]_{n}(t={a^{\dagger}}a)}\,\mathclose{:}_{W},$ (6.10)
+where $\bar{g}_{T}(z;0)$ comes from the $(L,R)=(1,1)$ case of (6.6), and $\bar{f}_{T}(z;0)$ comes similarly from (6.7). That is,
+$({a^{\dagger}}a{a^{\dagger}})^{n}=\mathopen{:}\,\mathrel{{a^{\dagger}}^{n}\,\mathbf{R}\left[\frac{1}{\sqrt{1+z^{2}}},\frac{2z}{1+\sqrt{1+z^{2}}}\right]_{n}\!(t={a^{\dagger}}a)}\,\mathclose{:}_{W}.$ (6.11)
 Up to normalization, the Riordan array $\mathbf{R}[d(z),h(z)]$ appearing here is familiar: versions appear as A108044 and A120616 in the OEIS [51]. The expression
-$\mathbf{R}_{o}\left[\frac{1}{\sqrt{1+4z^{2}}},\frac{2z}{1+\sqrt{1+4z^{2}}}\right]_{n,k}\!=(-1)^{(n-k)/2}\binom{n}{(n-k)/2}$ (92)
-for its matrix elements is known (the binomial coefficient is zero if $n-\nobreak k$ is odd).
-Here, $\mathbf{R}_{o}$ in distinction to $\mathbf{R}$ signifies an ordinary (non-exponential) Riordan array, meaning that in the definition of $\mathbf{R}_{o}[d(z),h(z)]_{n,k}$ the $n!/k!$ factor in the definition (27) used in this paper is absent.
+$\mathbf{R}_{o}\left[\frac{1}{\sqrt{1+4z^{2}}},\frac{2z}{1+\sqrt{1+4z^{2}}}\right]_{n,k}\!=(-1)^{(n-k)/2}\binom{n}{(n-k)/2}$ (6.12)
+for its matrix elements is known (the binomial coefficient is zero if $n-\penalty 10000k$ is odd).
+Here, $\mathbf{R}_{o}$ in distinction to $\mathbf{R}$ signifies an ordinary (non-exponential) Riordan array, meaning that in the definition of $\mathbf{R}_{o}[d(z),h(z)]_{n,k}$ the $n!/k!$ factor in the definition (3.3) used in this paper is absent.
 Altering the normalization and restoring the $n!/k!$ factor yields the claimed formula for $({a^{\dagger}}a{a^{\dagger}})^{n}$.
@@
 As with $e=0,1$, one substitutes $(A,B)=(e,1)=(2,1)$ and $L,R$ into the definitions in eq.
-(60) of the functions $g_{T}(D;s)$ and $z=f_{T}(D;s)$, and computes $\bar{g}_{T}(z;s),\allowbreak\bar{f}_{T}(z;s)]$ by inverting $g_{T},f_{T}]$ in the exponential Riordan group.
+(5.1) of the functions $g_{T}(D;s)$ and $z=f_{T}(D;s)$, and computes $\bar{g}_{T}(z;s),\allowbreak\bar{f}_{T}(z;s)]$ by inverting $g_{T},f_{T}]$ in the exponential Riordan group.
 The $s$-ordered representation of the operator is then $\bar{g}_{T}(\lambda{a^{\dagger}}^{e};s)\times\allowbreak\exp\left[({a^{\dagger}}a)\bar{f}_{T}(\lambda{a^{\dagger}}^{e};s)\right]$.
-But for each of the four operators, the representation is quite complicated. The stumbling block is that according to (60), $z=f_{T}(D;s)$ is given by
-$z=f_{T}(D;s)=\frac{8D(2-sD)}{\left[(2-sD)^{2}-D^{2}\right]^{2}},$ (93)
+But for each of the four operators, the representation is quite complicated. The stumbling block is that according to (5.1), $z=f_{T}(D;s)$ is given by
+$z=f_{T}(D;s)=\frac{8D(2-sD)}{\left[(2-sD)^{2}-D^{2}\right]^{2}},$ (6.13)
 and solving for $D=\bar{f}_{T}(z;s)$ entails solving the quartic
-$(1-s^{2})^{2}z\,D^{4}+8s(1-s^{2})z\,D^{3}-8\left[(1-3s^{2})z-s\right]D^{2}-16(1+2sz)\,D+16z=0.$ (94)
-The solution $D=D(z;s)$ satisfying $D=1\cdot z^{1}+\nobreak O(z^{2})$ can be expressed in radicals by employing the quartic formula, but owing to its complexity it is not given here.
-If $s=\pm 1$, the quartic (94) degenerates to a more easily solved quadratic, which can be viewed as the reason why the normal and anti-normal orderings of each of the four operators, which are supplied by Theorem 4.4, are relatively simple.
-But for general $s$, and even when $s=\nobreak 0$ (the Weyl ordering case), there seems to be no simplification.
-Whether it is possible to express $s$-ordered representations of operators ${\rm e}^{{a^{\dagger}}^{L}a{a^{\dagger}}^{R}}$ with $e=L+\nobreak R-\nobreak 1>2$ in terms of radicals, for general or even specialized $s$, is a question requiring further study.
+$(1-s^{2})^{2}z\,D^{4}+8s(1-s^{2})z\,D^{3}-8\left[(1-3s^{2})z-s\right]D^{2}-16(1+2sz)\,D+16z=0.$ (6.14)
+[…]
+[…]